### v1

# Conceptualization, Methodology, Writing – original draft, Software, Writing – review & editing

Abstract
Real-time prediction of corridor-level crash occurrence risk (COR) remains challenging, as existing near-miss based extreme value models oversimplify collision geometry, exclude vehicle–infrastructure (V–I) interactions, and inadequately capture spatial heterogeneity in vehicle dynamics. This study introduces a geometry-aware two-dimensional time-to-collision (2D-TTC) indicator within a Hierarchical Bayesian spatiotemporal grouped random parameter (HBSGRP) framework using a non-stationary univariate generalized extreme value (UGEV) model to estimate short-term COR in urban corridors. High-resolution trajectories from the Argoverse-2 dataset, covering 28 locations along Miami’s Biscayne Boulevard, were analyzed to extract extreme V–V and V–I near misses. The model incorporates dynamic variables and roadway features as covariates, with partial pooling across locations to address unobserved heterogeneity. Results show that the HBSGRP–UGEV framework outperforms fixed-parameter alternatives, reducing DIC by up to 7.5% for V–V and 3.1% for V–I near-misses. Predictive validation using ROC–AUC confirms strong performance: 0.89 for V–V segments, 0.82 for V–V intersections, 0.79 for V–I segments, and 0.75 for V–I intersections. Model interpretation reveals that relative speed and distance dominate V–V risks at intersections and segments, with deceleration critical in segments, while V–I risks are driven by speed, boundary proximity, and steering/heading adjustments. These findings highlight the value of a statistically rigorous, geometry-sensitive, and spatially adaptive modeling approach for proactive corridor-level safety management, supporting real-time interventions and long-term design strategies aligned with Vision Zero.
keywords:
[…]
1 Introduction
The increasing complexity of urban road networks, rising traffic volumes, and the coexistence of human-driven vehicles (HDVs) and autonomous vehicles (AVs) demand innovative safety assessment approaches. Urban corridors are particularly vulnerable, where a single incident can disrupt mobility, trigger cascading congestion, and impose severe economic losses. In 2023, traffic crashes in the United States claimed 40,901 lives, with a fatality rate of 1.26 per 100 million vehicle miles traveled (VMT) (National Center for Statistics and Analysis, 2025). The total annual cost of crashes is estimated to exceed \$1.85 trillion, including \$460 billion in direct economic losses and \$1.4 trillion in reduced quality of life (Blincoe et al., 2023). Despite extensive research on crash patterns, risk estimation, and mitigation strategies (Lanzaro et al., 2023; Kamel et al., 2023, 2024; Kamel and Sayed, 2024; Singh et al., 2024; Ghoul and Sayed, 2025; Singh et al., 2025; Anis et al., 2025b), accurate prediction of crash risk in urban corridors remains methodologically underdeveloped.
Most existing models focus exclusively on vehicle–vehicle (V–V) near-misses, largely due to the availability of vehicle trajectory data, while systematically excluding vehicle–infrastructure (V–I) interactions such as lane edges, curbs, barriers, and crosswalks. Yet V–I events account for a substantial proportion of severe run-off-road and roadway departure crashes (Islam et al., 2021). This omission leaves corridor-level safety models inherently incomplete. Recent advances in AV sensing and high-definition (HD) mapping, such as the Argoverse-2 dataset (Wilson et al., 2023), now enable the precise integration of roadway boundaries with vehicle trajectories, allowing both V–V and V–I near-misses to be modeled with geometric fidelity.
Crash risk prediction strategies based on historical retrospective crash data (Arun et al., 2021; Lee et al., 2017; Pei et al., 2011) suffer from well-documented limitations (Lord and Mannering, 2010; Lord et al., 2021; Mannering et al., 2020), including the rarity and stochastic nature of crashes, underreporting (Arun et al., 2021; Tarko, 2018), temporal delays, and spatial mismatches with real-time conditions. These challenges motivate a paradigm shift toward proactive, trajectory-based surrogate safety measures (SSMs) (Ali et al., 2023a; Arun et al., 2021; Mahmud et al., 2017). Derived from high-resolution vehicle trajectory data, SSMs reflect underlying near-miss dynamics (Li et al., 2024; Tarko, 2018; Davis et al., 2011). Among SSMs, time-based indicators such as Time-to-Collision (TTC), Post-Encroachment Time (PET), and Modified TTC (MTTC) are widely used (Hayward, 1971; Perkins and Harris, (1967; Minderhoud and Bovy, 2001; Ozbay et al., 2008; Allen et al., 1978; Venthuruthiyil and Chunchu, 2022). However, these formulations often assume constant velocity and linear trajectories, which bias results in boundary-constrained environments. Recent research has extended TTC into higher-fidelity two-dimensional (2D) indicators (Li et al., 2024; Anis et al., 2025b), incorporating latitude and longitude position, heading, and acceleration. Yet, even these advances typically employ simplified circular approximations, which can influence risk estimates in urban contexts where vehicle shape and alignment significantly influence near-miss likelihood and severity. To support reliable short-term crash occurrence risk (COR) estimation, SSM frameworks must enhance geometric fidelity to more accurately represent real-world interactions between vehicles and infrastructure. This study builds on these efforts by introducing a high-fidelity, geometry-aware 2D-TTC indicator that explicitly incorporates vehicle geometry, dynamic states, and roadway boundaries, ensuring realistic and efficient detection of both V–V and V–I near misses.
While SSMs capture frequent near-misses, a robust COR estimation framework requires extrapolating extreme near-misses to rare event outcomes. To date, several studies have proposed probabilistic (Saunier and Sayed, 2008), causal (Davis et al., 2011), and extreme value theory (EVT) (Fu and Sayed, 2021a, b; Zheng et al., 2021b; Ali et al., 2023a; Zheng et al., 2021a; Songchitruksa and Tarko, 2006; Zheng et al., 2014; Tarko, 2012). EVT provides a principled statistical foundation for this task, modeling the tail behavior of the extreme near-miss distribution. Two EVT approaches dominate: the Generalized Extreme Value (GEV) distribution using Block Maxima (BM) sampling and the Generalized Pareto Distribution (GPD) using Peaks Over Threshold (POT) sampling (Coles et al., 2001). The BM approach is particularly suitable for real-time analysis due to its short block structure, lower threshold sensitivity, and compatibility with discrete-time samples (Fu and Sayed, 2022a).
Early EVT applications in traffic safety employed stationary GEV models (Wang et al., 2018; Guo et al., 2019; Orsini et al., 2019; Alozi and Hussein, 2022; Hussain et al., 2022) that assumed constant parameterization across the study area. More recently, non-stationary GEV models have been introduced, allowing covariates to influence location and scale parameters (Tahir and Haque, 2024; Ali et al., 2022; Fu and Sayed, 2022a; Zheng et al., 2019; Nazir et al., 2023; Kar et al., 2024; Ali et al., 2023b; Alozi and Hussein, 2022; Zheng et al., 2014; Kar et al., 2023) and thereby capture heterogeneity across traffic states (Lanzaro et al., 2023; Kamel et al., 2023, 2024; Kamel and Sayed, 2024; Singh et al., 2024; Ghoul and Sayed, 2025; Singh et al., 2025; Anis et al., 2025b). Bayesian hierarchical extensions have further addressed data sparsity and unobserved heterogeneity (Zheng and Sayed, 2019; Fu and Sayed, 2022b). However, most applications remain geographically narrow and functionally limited to either intersections (Fu and Sayed, 2023, 2021c, 2022b) or segments (Kumar and Mudgal, 2024; Kamel et al., 2023; Kamel and Sayed, 2024; Anis et al., 2025b; Kamel et al., 2024). Although some studies have incorporated vehicle dynamics (Kumar and Mudgal, 2024; Fu and Sayed, 2022b; Anis et al., 2025b) and examined heterogeneity across traffic flow directions or roadway geometries (Kamel et al., 2023, 2024), these frameworks generally omit explicit integration of roadway geometry and directional heterogeneity at the corridor level. Such integration requires a modeling structure capable of jointly representing vehicle dynamics (which vary temporally) and roadway features (which vary spatially across segments and intersections). Therefore, a critical gap remains for a comprehensive Bayesian hierarchical GEV framework capable of jointly modeling vehicle dynamics, roadway geometry, flow directionality, and multilevel heterogeneity. Such a framework is essential for accurate corridor-level COR estimation that reflects the complex spatiotemporal dynamics of urban traffic environments. This paper makes the following unique contributions:
  • Develops a corridor-level COR framework that simultaneously incorporates both V–V and V–I interactions, addressing a critical omission in prior studies.
  • Introduces a high-fidelity, geometry-aware 2D-TTC indicator that realistically captures extreme near-miss events by accounting for vehicle geometry, dynamic states, and roadway boundaries.
  • Advances a Hierarchical Bayesian Spatiotemporal Grouped Random Parameter (HBSGRP) Univariate GEV model that integrates roadway geometry, traffic flow directionality, and multilevel heterogeneity, enabling corridor-wide yet location-sensitive crash risk estimation
Together, these contributions establish the first infrastructure-aware, statistically rigorous framework for a proactive COR framework in urban corridors, leveraging Arogovers-2 AV trajectory and HD map data to support both operational interventions and long-term safety planning.
The remainder of this paper introduces the proposed 2D-TTC indicators, the HBSGRP–UGEV modeling framework, and the data preparation used in the analysis. This is followed by the presentation of model estimation, validation, and comparative results. The paper concludes with key findings, limitations, and directions for future research.
2 Near-miss detection framework
This section presents the modeling framework for computing a geometry-aware, high-fidelity 2D-TTC to support COR estimation. The framework integrates a continuous-time kinematic bicycle model, vehicle geometry representation, and a spatial near-miss detection algorithm, allowing capture of V-V and V-I interactions. These components collectively enable the accurate forecasting of vehicle trajectories and the identification of potential near misses. The resulting architecture forms the computational backbone of the SSM analysis, ensuring a realistic representation of near-miss events in mixed-traffic environments.
2.1 Vehicle dynamics model
The vehicle dynamics module provides the predictive foundation of the framework. This adopts the kinematic bicycle model to represent vehicle footprint, as it is suitable for short-term near-miss analysis, where steering and acceleration dominate, and lateral slip can be neglected. The motion is represented by a four-dimensional (4D) dynamic state vector. At any time $t$, the state of vehicle $i$ is given by:
$\mathbf{x}_{i}(t)=\begin{bmatrix}x_{i}(t)\\[2.84526pt] y_{i}(t)\\[2.84526pt] \theta_{i}(t)\\[2.84526pt] v_{i}(t)\end{bmatrix}$ (1)
Here, $x_{i}(t),y_{i}(t)$ denote the vehicle’s global center coordinates, $\theta_{i}(t)$ its heading angle, and $v_{i}(t)$ its speed. Over the prediction window, each vehicle is assumed to maintain constant acceleration $a_{i}$ and steering angle $\delta_{i}$, simplifying short-horizon trajectory integration. The motion then follows the standard kinematic bicycle equations:
$\dot{x}_{i}(t)$ $=v_{i}(t)\cos\theta_{i}(t)$ (2)
$\dot{y}_{i}(t)$ $=v_{i}(t)\sin\theta_{i}(t)$ (3)
$\dot{\theta}_{i}(t)$ $=\frac{v_{i}(t)}{L_{i}}\tan\delta_{i}$ (4)
$\dot{v}_{i}(t)$ $=a_{i}$ (5)
Here, $L_{i}$ is the vehicle wheelbase, influencing the turning radius and path curvature of vehicle $i$. The steering angle $\delta_{i}$ and acceleration $a_{i}$ are treated as fixed inputs over the lookahead period. Equations (2-5) describe the rates of change in position, orientation, and velocity, determining the vehicle’s trajectory. To model interactions, individual states are concatenated into an eight-dimensional (8D) joint state vector, defined as:
$X(t)=\begin{bmatrix}x_{A}(t)&y_{A}(t)&\theta_{A}(t)&v_{A}(t)&x_{B}(t)&y_{B}(t)&\theta_{B}(t)&v_{B}(t)\end{bmatrix}^{\top}$ (6)
The vector field that governs the evolution of this combined state is F{F} defined as:
$\dot{X}(t)=F\bigl{(}X(t)\bigr{)}=\begin{bmatrix}v_{A}\cos\theta_{A}&v_{A}\sin\theta_{A}&\frac{v_{A}}{L_{A}}\tan\delta_{A}&a_{A}&v_{B}\cos\theta_{B}&v_{B}\sin\theta_{B}&\frac{v_{B}}{L_{B}}\tan\delta_{B}&a_{B}\end{bmatrix}^{\top}$ (7)
The state-space vector field in Eqn.7 gives the instantaneous rate of change of the joint state $\dot{X}(t)$. Its forward evolution is computed using the fourth-order Runge–Kutta (RK4) method, which balances accuracy and efficiency for nonlinear vehicle dynamics. The time horizon is discretized into $N$ steps of size $\Delta t$ ($t_{n}=n\Delta t$, $n=0,1,\dots,N$), and the RK4 update for state ${X}(t_{n+1})$ from ${X}(t_{n})$ is:
$X_{n+1}=X_{n}+\frac{\Delta t}{6}\left(k_{1}+2k_{2}+2k_{3}+k_{4}\right)$ (8)
where the intermediate slope evaluations ${k}_{1},\dots,{k}_{4}\in\mathbb{R}^{8}$ are defined as:
$k_{1}$ $=F(X_{n})$
$k_{2}$ $=F\left(X_{n}+\frac{\Delta t}{2}k_{1}\right)$
$k_{3}$ $=F\left(X_{n}+\frac{\Delta t}{2}k_{2}\right)$
$k_{4}$ $=F\left(X_{n}+\Delta t\cdot k_{3}\right)$ (9)
This integration is repeated for each timestep $n\in\{0,\dots,N-1\}$, producing a sequence of predicted states $\{{X}_{n}\}$ over the horizon. These trajectories form the basis for geometric collision checks between vehicles and roadway boundaries.
2.2 Vehicle geometry representation
While the dynamic model predicts vehicle positions and headings, accurate near-miss detection requires a realistic representation of the vehicle footprint. Simplifying vehicles as points or circles often misrepresents interactions, especially during lane changes, swerving, or angled approaches. To address this, each vehicle is modeled as a rigid rectangle defined by its actual length $L_{i}^{\text{veh}}$ and width $W_{i}^{\text{veh}}$ (vehicle $i\in\{A,B\}$). The vehicle is centered at its center of gravity, with four corners specified in the body-fixed coordinate system as:
${b}_{i}^{(j)}=\begin{bmatrix}\xi_{i}^{(j)}\\ \eta_{i}^{(j)}\end{bmatrix},\quad j=1,2,3,4$ (10)
where $\xi_{i}^{(j)}\in\left\{+\frac{L_{i}^{\text{veh}}}{2},\ -\frac{L_{i}^{\text{veh}}}{2}\right\}$ and $\eta_{i}^{(j)}\in\left\{+\frac{W_{i}^{\text{veh}}}{2},\ -\frac{W_{i}^{\text{veh}}}{2}\right\}$. This results in four corners: front-left, front-right, rear-left, and rear-right, expressed relative to the vehicle’s center in the local frame.
To compute the position of each corner in the global coordinate frame at time $t$, the body-fixed corner coordinates are first rotated by the vehicle’s heading angle $\theta_{i}(t)$, and then translated by the vehicle’s center position $(x_{i}(t),y_{i}(t))$. This transformation is written as:
${p}_{i}^{(j)}(t)=\begin{bmatrix}x_{i}(t)\\ y_{i}(t)\end{bmatrix}+R(\theta_{i}(t))\cdot{b}_{i}^{(j)}$ (11)
where $R(\theta_{i}(t))$ is the 2D rotation matrix defined as:
$R(\theta_{i}(t))=\begin{bmatrix}\cos\theta_{i}(t)&-\sin\theta_{i}(t)\\ \sin\theta_{i}(t)&\cos\theta_{i}(t)\end{bmatrix}$ (12)
This transformation rotates the local corner coordinates by the heading angle and translates them into the global frame. Repeating it for all corners at each timestep produces the vehicle’s time-varying geometric footprint, which serves as input for near-miss detection. These footprints enable precise evaluation of proximity-based events, both V-V and V-I, ensuring accurate 2D-TTC estimation while reducing false positives from geometric simplifications.
2.3 V-V near-miss detection
Vehicles are modeled as rigid rectangles, and their interactions are determined by predicted states and orientations. At each timestep of the RK4 integration, the global positions of all four corners are computed using the kinematic and rotational transformations. A collision is defined to occur when a corner of one vehicle lies within a proximity threshold of a corner of another. Let ${p}_{A}^{(j)}(t)$ and ${p}_{B}^{(k)}(t)$ denote the global coordinates of corners $j$ and $k$ of vehicles A and B at time $t$. A V–V near-miss is flagged if any pair of these corners satisfies the proximity condition.
${p}_{A}^{(j)}(t)={p}_{B}^{(k)}(t)$ (13)
An exact match between corners of two vehicles is highly unlikely due to floating-point arithmetic and continuous motion. Instead, a practical criterion is applied: a collision is flagged if the difference in both the x and y coordinates of any corner pair is less than or equal to a small threshold $\epsilon$:
$|x_{A}^{(j)}-x_{B}^{(k)}|$ $\leq\varepsilon$ (14)
$|y_{A}^{(j)}-y_{B}^{(k)}|$ $\leq\varepsilon$ (15)
The rotation–translation transformation from each vehicle’s center and heading angle gives the global corner coordinates, for vehicles $A$ and $B$:
$x_{A}^{(j)}$ $=x_{A}+\xi_{A}^{(j)}\cos\theta_{A}-\eta_{A}^{(j)}\sin\theta_{A}$ (16)
$y_{A}^{(j)}$ $=y_{A}+\xi_{A}^{(j)}\sin\theta_{A}+\eta_{A}^{(j)}\cos\theta_{A}$ (17)
$x_{B}^{(k)}$ $=x_{B}+\xi_{B}^{(k)}\cos\theta_{B}-\eta_{B}^{(k)}\sin\theta_{B}$ (18)
$y_{B}^{(k)}$ $=y_{B}+\xi_{B}^{(k)}\sin\theta_{B}+\eta_{B}^{(k)}\cos\theta_{B}$ (19)
At each timestep, all 16 corner-pair combinations are evaluated. A V–V near-miss is flagged if any pair satisfies the proximity condition, and the corresponding collision time $t_{c}$ is recorded for 2D-TTC computation. This discrete-time approach enables accurate detection of spatial overlap while avoiding the oversimplifications of circular vehicle approximations.
2.4 V-I near-miss detection
In addition to V-V interactions, the framework also detects near-misses between vehicles and roadway infrastructure such as curbs, medians, or barriers. Capturing these V–I events is critical for identifying edge encroachments or departures that often precede run-off-road crashes.
Road boundaries are represented as static polylines of $M$ discrete points $(X_{\ell},Y_{\ell})$ consistent with HD map formats in datasets such as Argoverse-2. At each timestep, the global positions of the four vehicle corners ${p}_{i}^{(j)}(t)$ are obtained from the predicted state. The distance between a vehicle corner $j$ and boundary point $\ell$, the distance at time $t$ is computed as:
$d_{j,\ell}(t)=\sqrt{(x^{(j)}(t)-X_{\ell})^{2}+(y^{(j)}(t)-Y_{\ell})^{2}}$ (20)
A near-miss is flagged if $d_{j,\ell}(t)\leq\varepsilon$ for any corner and boundary point, where is a proximity threshold accounting for numerical tolerances. When this occurs, the corresponding time $t_{c}$ is recorded as the instant of boundary contact. These events are then passed to the 2D-TTC framework to support extreme risk estimation.
2.5 Spatiotemporal COR
The near-miss detection framework integrates vehicle dynamics, geometric transformations, and spatial proximity checks to identify both V–V and V–I events. Forward trajectories are simulated using RK4 integration, vehicle corners are transformed into global coordinates, and geometric overlaps are evaluated to detect potential near-misses across the prediction horizon. The procedure is summarized in Algorithm 1, and a conceptual illustration is provided in Fig. 2.5. These steps yield a time series of 2D-TTC values for each interaction block.
Input: Initial state $\mathbf{X}(0)$, vehicle parameters $l_{i}$, $L_{i}^{\text{veh}},W_{i}^{\text{veh}}$, control inputs $a_{i}$, $\delta_{i}$, and boundary points $\{(X_{\ell},Y_{\ell})\}$
Output: Earliest collision time $t_{c}$
1
2for each timestep $n=0$ to N1N-1 do
3   Simulate $\mathbf{X}_{n}\rightarrow\mathbf{X}_{n+1}$ using RK4 integration;
4   Compute global corner positions $\mathbf{p}_{i}^{(j)}$ for each vehicle;
5
6 for each corner pair $(j,k)\in\{1,2,3,4\}^{2}$ do
7    if $|x_{A}^{(j)}-x_{B}^{(k)}|\leq\varepsilon$ or $|y_{A}^{(j)}-y_{B}^{(k)}|\leq\varepsilon$ then
8         Record V-V near-miss, set $t_{c}$;
9
10
11
12 for each corner $j$ and boundary point $\ell$ do
13      Compute $d_{j,\ell}(t)=\sqrt{(x^{(j)}-X_{\ell})^{2}+(y^{(j)}-Y_{\ell})^{2}}$;
14    if $d_{j,\ell}(t)\leq\varepsilon$ then
15         Record V-I near-miss, set $t_{c}$;
16
17
18
return the earliest collision time $t_{c}$ as 2D-TTC
Algorithm 1 Near-miss detection based on state integration and geometric overlap
Fig. 1.: Computation of 2D-TTC for near-miss scenarios
While this geometric framework enables precise localization of near-misses, it does not capture the probabilistic structure or temporal evolution of COR. To address this, the extracted blockwise minima of 2D-TTC are modeled using the EVT model. For each block, the model estimates the probability that 2D-TTC falls below a critical safety threshold, which defines the COR metric as a short-term indicator of crash risk. Block duration is chosen to balance temporal sensitivity with statistical stability across diverse contexts. This framework, therefore, combines deterministic near-miss detection with probabilistic EVT modeling to enable localized, data-driven assessment of crash-prone conditions. The complete methodological pipeline, from trajectory simulation to dynamic risk estimation, is summarized in Fig. 2.5.
Fig. 2.: COR time series generation
3 Extreme value theory (EVT)
Extreme Value Theory (EVT) provides the statistical foundation for estimating COR by extrapolating rare, high-severity events from frequent near-misses. In this study, extreme events are represented by block maxima of 2D-TTC derived from high-resolution vehicle trajectories (Algorithm 1). EVT allows risk estimation without requiring large volumes of historical crash data (Zheng and Sayed, 2020; Fu and Sayed, 2022a). Two EVT approaches are commonly applied: the Block Maxima (BM) method, which fits maxima or minima from fixed-duration blocks using the Generalized Extreme Value (GEV) distribution (Coles et al., 2001), and the Peaks Over Threshold (POT) method, which models threshold exceedances using the Generalized Pareto Distribution (GPD). While POT can be data-efficient, it is highly sensitive to threshold choice and less compatible with discrete, sensor-derived datasets. By contrast, BM aligns naturally with real-time AV data, avoids subjective thresholds, and supports scalable, time-series analysis (Fu and Sayed, 2022a; Songchitruksa and Tarko, 2006). Accordingly, this study adopts the BM–GEV framework (see Fig. 2.5).
The GEV has been widely applied in traffic safety research. Early studies modeled near-missets with univariate and multivariate GEV formulations (Orsini et al., 2020; Zhang et al., 2019; Zheng et al., 2014). Recent advances incorporate nonstationary structures, linking GEV parameters to covariates such as vehicle dynamics (Anis et al., 2025b; Fu and Sayed, 2021c, 2022b). Bayesian hierarchical models have further addressed sparse data and unobserved heterogeneity across vehicles, sites, and traffic contexts. The increasing availability of high-resolution AV trajectory and HD map data has highlighted the need for flexible GEV formulations that jointly capture V–V and V–I interactions.
Despite this progress, most UGEV models still neglect the joint influence of roadway geometry and vehicle dynamics, two critical but distinct sources of heterogeneity. Roadway geometry (e.g., lane count, curvature, median type) is spatially structured and relatively stable within segments or intersections, while vehicle dynamics (e.g., speed, acceleration) vary temporally and spatially, driving short-term behavioral volatility. Capturing this multilevel heterogeneity is crucial for estimating corridor- and network-level crash risk. To address this gap, this study develops a Hierarchical Bayesian Spatiotemporal Grouped Random Parameter (HBSGRP) UGEV model. Unlike standard random-parameter models, which assign variation independently to each unit, grouped random parameters capture shared variation across clusters (e.g., segment types, corridor directions) while allowing for within-group variability. This structure provides both flexibility and parsimony, making it well-suited for high-resolution, corridor-scale crash risk modeling (Washington et al., 2020; Cai et al., 2018; Islam et al., 2023).
3.1 Univariate generalized extreme value (UGEV)
As established in the previous section, the BM-based UGEV model is employed to characterize the tail behavior of 2D-TTC, which represents the most safety-critical near-miss events in both V–V and V–I interactions. To implement this framework, continuous vehicle trajectories are partitioned into fixed-duration windows (e.g., 11 s). Within each window, the minimum 2D-TTC is extracted and negated, consistent with the upper-tail modeling convention of UGEV (Coles et al., 2001). The resulting sequence of BM serves as the input for parameter estimation. Throughout this study, we focus on maxima, denoted as $M_{n}$, under the assumption that each block consists of $n$ independent and identically distributed (i.i.d.) random variables, $X_{1},X_{2},\ldots,X_{n}$. To align with conventional UGEV formulations, we consider the negative values of the original variables derived from Algorithm 1, thereby defining the block maximum as $M_{n}=\max\{-X_{1},-X_{2},\ldots,-X_{n}\}$. Under suitable regularity conditions, if there exist sequences of constants $a_{n}\in\mathbb{R}$ and $b_{n}>0$, the normalized block maxima $M_{n}^{*}=(M_{n}-a_{n})/b_{n}$ converge in distribution to a non-degenerate limit $G(x)$ as $n\to\infty$, then: $\Pr\left(\frac{M_{n}-a_{n}}{b_{n}}\leq X\right)\xrightarrow[n\to\infty]{}f(x)$, where $G(x)$ is a non-degenerate distribution function. According to EVT, the limiting distribution $G(x)$ must belong to one of three families: Gumbel, Fréchet, or Weibull, each of which can be captured by the parametric form of the GEV distribution (Coles et al., 2001). Let the UGEV be defined as:
$G(x;\mu_{x},\sigma_{x},\xi_{x})=\exp\left\{-\left[1+\xi_{x}\left(\frac{x-\mu_{x}}{\sigma_{x}}\right)\right]^{-1/\xi_{x}}\right\},\quad\text{where }1+\xi_{x}\left(\frac{x-\mu_{x}}{\sigma_{x}}\right)>0$ (21)
Here, $\mu_{x}\in\mathbb{R}$ is the location parameter, $\sigma_{x}>0$ the scale, and $\xi_{x}\in\mathbb{R}$ the shape. The value of $\xi_{x}$ determines tail heaviness: $\xi_{x}>0$ corresponds to the heavy-tailed Fréchet, $\xi_{x}<0$ to the bounded Weibull, and $\xi_{x}=0$ to the light-tailed Gumbel case.
3.1.1 Hierarchical Bayesian Spatiotemporal Grouped Random Parameter (HBSGRP) model
To capture structured heterogeneity in extreme outcomes across the corridor, we employ a Hierarchical Bayesian Spatiotemporal Grouped Random Parameter (HBSGRP) model. Unlike conventional random-parameter models that assign variation to individual units, the HBSGRP formulation introduces variation at the group level. This structure allows systematic differences across predefined observational groups while maintaining parsimony and interpretability. In this specification, covariates are partitioned into two sets: fixed effects $X$, which apply uniformly across all observations, and group-varying effects $Z$, whose coefficients differ across groups. For a dataset divided into $K$ groups, with observation $i$ belonging to group $k(i)$, both the intercept and selected slopes are allowed to vary by group. This enables the capture of localized relationships while preserving overall model coherence. The model is expressed within a three-layer hierarchical Bayesian framework. The data layer specifies the likelihood of observed block maxima of 2D-TTC from V–V and V–I interactions, assuming a UGEV distribution. Process layer links covariates to UGEV parameters through fixed and group-level random effects, with optional latent terms (e.g., Gaussian processes) to capture residual spatial–temporal dependence. Prior layer defines prior distributions for all parameters and hyperparameters, supporting regularization and uncertainty quantification.
Bayesian inference proceeds by updating prior beliefs with observed data to obtain the posterior distribution. Formally, the joint posterior of parameters $\Theta$, given data $X$, is:
$q(\Theta\mid X)\propto q_{\text{data}}(X\mid\Theta)\cdot q_{\text{process}}(\Theta\mid\Psi)\cdot q_{\text{prior}}(\Psi)$ (22)
In this formulation, $q(\Theta\mid X)$ denotes the posterior distribution, integrating both the data likelihood and prior knowledge. The term $q_{\text{data}}(X\mid\Theta)$ corresponds to the likelihood function defined by the EVT assumptions. The component $q_{\text{process}}(\Theta\mid\Psi)$ captures the hierarchical dependency structure, including the group-varying parameters. Lastly, $q_{\text{prior}}(\Psi)$ encompasses the prior distributions of hyperparameters that control the model’s flexibility and penalize over-complexity.
The likelihood function for the data layer is specified as each extreme value sample $X_{k,i}$, observed for event $i$ within group $K$, is modeled as:
$q_{\text{data}}(X_{k,i}\mid\Theta)=\prod_{k=1}^{K}\prod_{i=1}^{N_{k}}\frac{1}{\sigma_{k,i}}\exp\left\{-\left[1+\xi_{k,i}\left(\frac{X_{k,i}-\mu_{k,i}}{\sigma_{k,i}}\right)\right]^{-1/\xi_{k,i}}\right\}\left[1+\xi_{k,i}\left(\frac{X_{k,i}-\mu_{k,i}}{\sigma_{k,i}}\right)\right]^{-1-1/\xi_{k,i}}$ (23)
The location ($\mu_{k,i}$), scale ($\sigma_{k,i}$), and shape ($\xi_{k,i}$) parameters are expressed as functions of group-level fixed-effect covariates $X_{k,j}$ (e.g., road geometry, speed limit) and sample-specific random-effect covariates $W_{k,i,j}$ (e.g., speed, acceleration, time-to-collision) as follows:
$\mu_{k,i}=\beta_{\mu,0}+\sum_{j=1}^{J}\beta_{\mu,j}X_{k,j}+\sum_{j=1}^{J^{\prime}}\gamma_{\mu,j,k}W_{k,i,j}$ (24)
$\vartheta_{k,i}=\beta_{\sigma,0}+\sum_{j=1}^{J}\beta_{\sigma,j}X_{k,j}+\sum_{j=1}^{J^{\prime}}\gamma_{\sigma,j,k}W_{k,i,j},\quad\sigma_{k,i}=\exp(\vartheta_{k,i})$ (25)
$\xi_{k,i}=\beta_{\xi,0}+\sum_{j=1}^{J}\beta_{\xi,j}X_{k,j}$ (26)
In the above equations $X_{k,j}$ defined as group-level fixed-effect covariate $j$, $W_{k,i,j}$is sample-specific random-effect covariate $j$. $\beta_{\mu,0},\beta_{\sigma,0},\beta_{\xi,0}$ as fixed intercepts for respective parameters. $\beta_{\mu,j},\beta_{\sigma,j},\beta_{\xi,j}$: fixed-effect coefficients. $\gamma_{\mu,j,k},\gamma_{\sigma,j,k}$: random-effect coefficients, varying across groups $k$. Therefore, the likelihood for the process layer, modeling within-group variability, is defined as:
$q_{\text{process}}(\Theta\mid\Psi)=\prod_{k=1}^{K}\prod_{i=1}^{N_{k}}\frac{1}{\sqrt{2\pi\tau_{\mu}^{2}}}\exp\left\{-\frac{1}{2\tau_{\mu}^{2}}(\mu_{k,i}-\bar{\mu}_{k})^{2}\right\}\\ \times\frac{1}{\sqrt{2\pi\tau_{\vartheta}^{2}}}\exp\left\{-\frac{1}{2\tau_{\vartheta}^{2}}(\vartheta_{k,i}-\bar{\vartheta}_{k})^{2}\right\}\times\frac{1}{\sqrt{2\pi\tau_{\xi}^{2}}}\exp\left\{-\frac{1}{2\tau_{\xi}^{2}}(\xi_{k,i}-\bar{\xi}_{k})^{2}\right\}$ (27)
where $\bar{\mu}_{k}$, $\bar{\vartheta}_{k}$, and $\bar{\xi}_{k}$ are group-specific average parameters.
Prior distributions are assigned to all model parameters. The priors for fixed and random effect coefficients are as follows:
$q_{\text{prior}}(\Psi)=q_{\beta_{\mu,0}}(\beta_{\mu,0})\cdot\prod_{j=1}^{J}q_{\beta_{\mu,j}}(\beta_{\mu,j})\cdot\prod_{k=1}^{K}\prod_{j=1}^{J^{\prime}}q_{\gamma_{\mu,j,k}}(\gamma_{\mu,j,k})\\ \times q_{\beta_{\sigma,0}}(\beta_{\sigma,0})\cdot\prod_{j=1}^{J}q_{\beta_{\sigma,j}}(\beta_{\sigma,j})\cdot\prod_{k=1}^{K}\prod_{j=1}^{J^{\prime}}q_{\gamma_{\sigma,j,k}}(\gamma_{\sigma,j,k})\ \times q_{\beta_{\xi,0}}(\beta_{\xi,0})\cdot\prod_{j=1}^{J}q_{\beta_{\xi,j}}(\beta_{\xi,j})$ (28)
Hyperpriors for the variance parameters $\tau_{\mu}^{2},\tau_{\vartheta}^{2},\tau_{\xi}^{2}$ are modeled using inverse gamma distributions:
$\tau_{\mu}^{2}\sim\text{IG}(\alpha_{\mu},\beta_{\mu}),\quad\tau_{\vartheta}^{2}\sim\text{IG}(\alpha_{\vartheta},\beta_{\vartheta}),\quad\tau_{\xi}^{2}\sim\text{IG}(\alpha_{\xi},\beta_{\xi})$ (29)
where $\text{IG}(\alpha,\beta)$ denotes an inverse gamma prior distribution ensuring positivity constraints for variance parameters.
3.2 COR estimation
Building on the HBSGRP–UGEV estimates, the framework compute COR by evaluating the probability that 2D-TTC for both V–V and V–I interactions falls below a critical threshold within each block. Unlike severity-stratified approaches, this framework estimates the total number of latent extreme near-miss events regardless of outcome severity, enabling a holistic assessment of potential COR under mixed traffic conditions.
Let $i$ denote block $i$ within group $k$, where groups correspond to spatial or functional corridor classifications (e.g., intersections or directional segments). The minimum 2D-TTC in each block is modeled with the HBSGRP-UGEV distribution, parameterized by $\mu_{k,i}$, $\sigma_{k,i}$, and $\xi_{k,i}$. Let $\omega$ be the critical 2D-TTC threshold below which a near-miss is considered crash-prone. The probability of a crash-level event in block $i$ of group $k$ is:
$P_{k,i}^{\text{crash}}=\Pr(2D-TTC_{k,i}\leq\omega)=1-\exp\left\{-\left[1-\xi_{k,i}\left(\frac{\omega-\mu_{k,i}}{\sigma_{k,i}}\right)\right]^{-1/\xi_{k,i}}\right\},\quad\text{where }1-\xi_{k,i}\left(\frac{\omega-\mu_{k,i}}{\sigma_{k,i}}\right)>0$ (30)
Assuming a homogeneous Poisson process for exceedances, the expected COR for group $k$ is obtained by aggregating exceedance probabilities across blocks and normalizing by exposure duration $T$. Here $y_{k,i}$ is the number of near-misses (V-V or V-I) in block $i$, and $N_{k}$ the number of blocks in group $k$. The expected COR for group $k$ is:
$COR_{k}=\frac{1}{T}\sum_{i=1}^{N_{k}}y_{k,i}\cdot P_{k,i}^{\text{crash}}=\frac{1}{T}\sum_{i=1}^{N_{k}}y_{k,i}\cdot\left[1-\exp\left\{-\left[1-\xi_{k,i}\left(\frac{\omega-\mu_{k,i}}{\sigma_{k,i}}\right)\right]^{-1/\xi_{k,i}}\right\}\right]$ (31)
The total expected crash frequency across the study area is then:
$CF_{\text{total}}=\sum_{k=1}^{K}COR_{k}=\frac{1}{T}\sum_{k=1}^{K}\sum_{i=1}^{N_{k}}y_{k,i}\cdot\left[1-\exp\left\{-\left[1+\xi_{k,i}\left(\frac{\omega-\mu_{k,i}}{\sigma_{k,i}}\right)\right]^{-1/\xi_{k,i}}\right\}\right]$ (32)
This formulation integrates exceedance probabilities from the HBSGRP–UGEV model with observed V–V and V–I near-miss counts, maintaining a direct link between real-world interaction intensity and modeled crash propensity. By separating probability from exposure, the framework yields interpretable, proactive crash frequency estimates aligned with EVT-based safety assessment.
4 Data description
The effectiveness of real-time COR estimation depends critically on the fidelity and coverage of trajectory data. Traditional sensing technologies, such as roadside cameras, loop detectors, and fixed LiDAR, have supported safety analysis by capturing speed, spacing, and acceleration patterns (Islam et al., 2021; Yuan et al., 2019; Li et al., 2020). However, these systems suffer from narrow fields of view, occlusion, overlapping detections, and high maintenance costs, which limit scalability and data quality (St-Aubin et al., 2013). Recent advances in AV sensing offer a promising alternative. AV datasets offer high-frequency, multimodal trajectories with unprecedented spatial precision, allowing for detailed analysis of kinematics, spacing, and maneuvers. This richness allows for realistic near-miss detection in complex traffic environments. Despite this potential, AV data remain underutilized in generalized frameworks for real-time COR estimation. For example, studies using the Waymo Open Dataset have largely focused on localized near-miss events without incorporating roadway geometry or V-I interactions, limiting broader applicability (Anis et al., 2025b).
To address these challenges, this study utilizes the Argoverse-2 dataset (Wilson et al., 2023), which was collected in 2021 across six U.S. cities (Austin, Detroit, Miami, Palo Alto, Pittsburgh, and Washington, D.C.) using SAE Level 4 AVs. Argoverse-2 comprises more than 250k driving scenarios spanning 113 roadway segments, with each 11-second scenario sampled at 10 Hz. Compared with Waymo Open (Ettinger et al., 2021), Lyft Level-5 (Houston et al., 2021), and nuScenes (Caesar et al., 2020), Argoverse-2 is uniquely curated to emphasize safety-critical interactions such as intersection near-misses, lane changes, and pedestrian crossings—conditions most indicative of crash risk. A defining feature of Argoverse-2 is its scene-adaptive sampling, which prioritizes high-interaction episodes rather than routine free-flow traffic. Scenarios are annotated with agent IDs, AV driving mode, and map-relative lane-following behaviors. The sensor suite includes seven ring cameras and two front-facing stereo cameras (20 Hz) supported by dual 64-beam VLP-32C LiDAR sensors, producing  107k 3D points per frame. HD maps contain lane centerlines, crosswalks, medians, and drivable areas registered to local coordinates.
Table 1: Open-source AV datasets
Dataset Sensora Tasksb Hzc HD Map Scope / Notes
Argoverse-2 (Wilson et al. 2023) C, L D, T, S, M 10 Yes 250k scenarios across 113 segments; 11s each; 10 classes; 6 cities
nuScenes (Caesar et al. 2020) C, L, R D, T, S, M 2 No 40k scenarios across  1k segments; 23 classes; 2 cities
Waymo (Ettinger et al. 2021) C, L D, T, M 10 Yes 390k scenarios over  1950 segments; 12.6M trajectories; 6 cities
Lyft Level 5 (Kesten et al. 2019) C, L, R M 10 Yes 170k AV instances (25s) in one city; 10 classes
KITTI (Geiger et al. 2013) L D, T, M 10 No 15k scenes; classic full-stack AV benchmark
DAIR-V2X (Yu et al. 2022) L D Yes 71254 image frames for V2X scenes
OPV2V (Xu et al. 2022) C, L D, T, M 10 No 33k V2V simulation samples; 18k V2X frames (CARLA)
DeepAccident (Wang et al. 2024) C, L, R D, T, S, M 10 Yes 285k simulation samples incl. 57k V2X scenarios
Shift (Sun et al., 2022) C, L S, M 10 No 2.5M frames; 4850 segments; 8 cities
  • a
    Sensor: Camera (C), LiDAR (L), Radar (R).
  • b
    Tasks: Detection (D), Tracking (T), Segmentation (S), Motion Forecasting (M).
  • c
    Hz: Sampling frequency; “–” not reported.
Argoverse-2 was selected after benchmarking open-source AV datasets (Table 1). Waymo Open offers a greater total mileage but lower near-miss density and no AV-behavior labels. Lyft Level-5 is confined to a single city and lacks scene-level filtering. nuScenes samples at 2 Hz, limiting the capture of rapid interactions. In contrast, Argoverse-2 balances spatial coverage, sensor fidelity, near-miss-rich scenarios, and behavioral annotations, making it well-suited to this study’s goals. This study used motion-forecasting and HD-map modules of Argoverse-2 to extract vehicle trajectories, compute near-miss indicators, and encode roadway geometry.
4.1 Data Preprocessing
To operationalize the proposed framework, both the motion forecasting and HD map modules of the Argoverse-2 dataset were processed through a structured pipeline. The preprocessing consisted of four key stages: (i) extraction of multi-agent trajectories, (ii) transformation of coordinates from local map space to global geographic space, (iii) integration with HD map features to contextualize V-I interactions, and (iv) selection of a high-risk urban corridor for detailed analysis.
4.1.1 Motion forecasting dataset
Each motion forecasting scenario follows a hierarchical structure with three levels: scenario metadata, agent tracks, and frame-level states. Scenarios are 11s duration and are sampled at 10 Hz, producing 110 time steps per vehicle for fine-grained interaction modeling. At the scenario level, metadata include a unique scenario ID, the recording city, and the focal agent, typically the ego AV. Each scenario then contains multiple agent tracks, where each track corresponds to a road user and specifies the object type (e.g., vehicle, pedestrian, cyclist), scoring category (focal, scored, or unscored), and a time-ordered sequence of states. Frame-level states capture position, velocity, and heading at each timestep, forming the basis for computing 2D-TTC.
To implement the framework, we developed a Python-based batch pipeline using the Argoverse-2 SDK. The pipeline parsed motion forecasting scenarios across all cities and extracted tracks for all agents—not just the focal vehicle—to retain critical multi-agent interactions. For each track, frame-level kinematic attributes were retrieved, including $(x,y)$ position, heading, velocity components, and timestamps.
Because positions were reported in local, city-specific frames, all coordinates were transformed to globally referenced WGS84. This involved (i) projecting each city’s local origin into its UTM zone (e.g., Zone 17 for Miami, Pittsburgh, and Detroit), and (ii) offsetting local $(x,y)$ positions and inverse-projecting them to latitude–longitude using pyproj. This geospatial normalization provided consistent referencing across cities and ensured compatibility with HD map overlays. Batch scripts automatically traversed scenario folders, performed attribute extraction and coordinate conversion, and exported unified CSV datasets containing both original city-frame and converted lat/lon values. These datasets were then used for spatial filtering, interaction detection, and overlay validation.
The processed files formed the basis for V–V and V–I near-miss detection. For each valid scenario, pairwise distances and relative approach rates were computed for all vehicle pairs, while V-I distances were derived from buffered polygons of lane edges and curbs in the HD map. Scenarios with no substantive interaction (e.g., straight-lane cruising without lane changes or turns) were excluded, ensuring the dataset emphasized behaviorally informative, high-risk conditions consistent with proactive safety modeling. Non-vehicle agents (e.g., pedestrians, cyclists) were also excluded at this stage, though the pipeline is extensible to incorporate vulnerable road users in future work.
4.1.2 HD map
The Argoverse-2 HD map dataset provides vectorized representations of roadway geometry and drivable space, enabling infrastructure-aware risk estimation. These maps are distributed as JSON files in local city coordinates and contain lane centerlines, left/right boundaries, crosswalk polygons, and drivable area definitions.
Lane-level geometry was extracted using the official Argoverse-2 API. The get_all_lane_segments() method returns centerline polylines along with left/right boundaries that define the navigable road space. Additional features, such as crosswalk polygons and drivable area masks, were also retrieved to support validation and visualization, though only lane boundaries and centerlines were directly used in near-miss detection.
Because all coordinates are expressed in city-specific Cartesian frames, global georeferencing was required to align HD maps with trajectory data. For each city, the WGS84 geodetic origin was projected into its respective UTM zone (e.g., Zone 17 for Miami and Pittsburgh, Zone 14 for Austin, Zone 10 for Palo Alto, Zone 18 for Washington, D.C.) using the pyproj library. Each local $(x,y)$ coordinate was then offset by this origin and inverse-projected into latitude–longitude, ensuring spatial consistency across cities. Once transformed, all HD map components: lane boundaries, centerlines, and crosswalks—were indexed using their global coordinates and converted into shapefiles or GeoPandas geometries. This geospatial transformation enabled efficient spatial querying, collision proximity analysis, and visual overlay with satellite imagery and trajectory data. In particular, the lane boundaries were instrumental in detecting V–I near-misses by allowing for the computation of lateral vehicle encroachments beyond the legal driving space. From the processed dataset, a high-density urban corridor in Miami was present to showcase the transformation (see Fig. 4.1.2). This illustrates the transformation of AV data into real-world coordinates, its alignment with HD maps and satellite imagery, and the resulting traffic patterns that highlight interaction hotspots across intersections and midblocks.
Fig. 3.: AV data transformed to real-world coordinates and aligned: (a) trajectories with lane boundaries and crosswalks, (b) validation with satellite imagery, and (c–f) directional traffic patterns highlighting interaction hotspots.
4.1.3 Trajectory preprocessing and derivation of vehicle dynamics
Before computing 2D-TTC, a series of preprocessing steps was applied to ensure the precision, continuity, and physical plausibility of vehicle motion data from the Argoverse-2 dataset. These steps included the extraction of raw trajectory data, trajectory smoothing and filtering, and the derivation of dynamic variables such as speed, acceleration, and steering angle. First, all multi-agent trajectories were transformed from local city-specific frames to global WGS84 coordinates, enabling consistent alignment with HD map geometries. Raw trajectories were then examined for noise caused by sensor jitter, occlusion, or missed detections—issues more common in human-driven vehicles than in AV tracks. To suppress such noise while preserving curvature and lane-change dynamics, cubic B-spline smoothing was applied to the position coordinates (Eilers and Marx, 1996; Choi et al., 2024). This method effectively removed local fluctuations while retaining key geometric features (see Fig.4.1.3). Stationary or irrelevant agents, such as parked vehicles and bicycles, were filtered out to reduce computational load and focus only on interaction-relevant entities.
Fig. 4.: Effect of B-spline smoothing on a sample trajectory, showing removal of noise while preserving geometric fidelity
Dynamic control inputs were then computed. Instantaneous speed $v_{t}$ was derived as the Euclidean norm of velocity components:
$v_{t}=\sqrt{v_{x,t}^{2}+v_{y,t}^{2}}$ (33)
Residual noise in $v_{t}$ was suppressed with a Savitzky–Golay filter (Chen et al., 2004) using a second-order polynomial and a 110-frame window (11s). Longitudinal acceleration ($a_{t}$) was obtained by first-order numerical differentiation:
$a_{t}=\frac{v_{t+1}-v_{t}}{\Delta t}$ (34)
where $\Delta t=0.01$s reflects the 10 Hz sampling rate. Heading $\theta_{t}$ was computed from directional vectors of each track and was unwrapped to maintain continuity across discontinuities at $\pm\pi$. The yaw rate $\dot{\delta}_{t}$ was derived from the kinematic bicycle model:
$\dot{\theta}_{t}=\frac{\theta_{t+1}-\theta_{t}}{\Delta t}$ (35)
$\delta_{t}=\tan^{-1}\left(\frac{L\cdot\dot{\theta}_{t}}{v_{t}}\right)$ (36)
The final dataset included per-frame values of speed, acceleration, heading, yaw rate, and steering angle for each track, providing stable and interpretable control inputs. These preprocessed trajectories formed the basis for subsequent near-miss detection using Algorithm 1 and the HBSGRP-UGEV model.
4.2 Study area
Under realistic and heterogeneous urban traffic conditions, a representative high-volume corridor was selected along Biscayne Boulevard in Miami, Florida. This arterial extends from Downtown Miami through the Upper East Side toward the northern city limits, serving as a critical connector to major activity centers, including the Miami International Airport region. It is characterized by multimodal interactions, geometric complexity, frequent turning maneuvers, and diverse traffic control features. The corridor was chosen for its rich coverage in the Argoverse-2 motion forecasting dataset, HD map, and its suitability for extracting sufficient samples for extreme near-miss scenarios.
As shown in Fig. 4.2, the study area was extracted using ArcGIS Pro based on scenario density, HD map fidelity, and heterogeneity of vehicle maneuvers. The extracted network includes a continuous corridor with multiple intersections, midblock segments, and bidirectional traffic operations. For analytical consistency, the corridor was subdivided into nine directional segments (S1–S9) and ten intersections (I1–I10). Segment boundaries were defined midblock to maintain flow continuity, while intersections were delineated to capture near-miss-prone zones such as turning paths, crosswalks, and lane merges. Each subregion was mapped with lane boundaries, crosswalks, and vehicle trajectories to preserve geometric fidelity and directional alignment. Insets in Fig. 4.2 illustrate the spatial configuration of the segmented units, which serve as the basis for grouped modeling and localized risk estimation.
Fig. 5.: Segmented study corridor along Biscayne Boulevard, Miami, with lane boundaries and crosswalks. Insets show direction-specific segments (S1–S9) and intersections (I1–I10)
Table 2 provides a detailed profile of the 28 study subregions, consisting of 10 intersections and 18 directional segments. Intersections generally show fewer scenarios and unique vehicles (e.g., I2 with 47 scenarios and 220 vehicles, I4 with 96 scenarios and 568 vehicles) compared to adjacent midblock segments, reflecting shorter dwell times but higher near-miss intensity. In contrast, segments accumulate a significantly higher number of scenarios and vehicles due to sustained travel. For example, S7 (NE 19-17 St) records 308 northbound and 332 southbound scenarios involving more than 4000 vehicles, while S8 (NE 17–15 St) captures nearly 4400 vehicles in 608 scenarios.
Table 2: Profile of study locations along Biscayne Boulevard
ID Location Direction Scenarios Unique vehicles Lane no Lane width (m) Driveway/Minor intersection Median type
Segment
S1 (NE 36–33 St) NB / SB 131 / 111 1706 / 813 3 / 3 3.5 3 Raised
S2 (NE 33–29 St) NB / SB 151 / 155 1532 / 1148 3 / 3 3.64 3 Raised
S3 (NE 29–26 St) NB / SB 191 / 186 1721 / 1295 3 / 3 3.36 4 Flush paved
S4 (NE 26–22 St) NB / SB 176 / 199 1542 / 1581 3 / 3 3.51 4 Flush paved
S5 (NE 22–21 St) NB / SB 127 / 123 705 / 643 3 / 2 3.3 No Flush paved
S6 (NE 21–19 St) NB / SB 235 / 245 1588 / 1861 3 / 3 3.62 1 Flush paved
S7 (NE 19–17 St) NB / SB 308 / 332 1704 / 2410 2 / 2 3.12 1 Flush paved
S8 (NE 17–15 St) NB / SB 294 / 314 1955 / 2406 3 / 3 3.57 2 Flush paved
S9 (NE 15–13 St) NB / SB 232 / 216 1565 / 1217 3 / 2 3.43 1 Flush paved
Intersection
I1 NE 36 St 97 743
I2 NE 33 St 47 220
I3 NE 29 St 132 699
I4 NE 26 St 96 568
I5 NE 22 St 106 640
I6 NE 21 St 134 647
I7 NE 19 St 214 1332
I8 NE 17 St 205 1328
I9 NE 15 St 252 1332
I10 NE 13 St 232 1139
Lane configurations vary between 2 and 3 lanes per direction, with average widths ranging from 3.1m (S7) to 3.6m (S6), indicating subtle but important differences in cross-sectional design. Driveway and minor intersection density also differs across the corridor: segments such as S3 and S4 contain four access points, while others like S5 and S6 have few or none, creating heterogeneous exposure to side-entry near-misses. Median type alternates between raised (e.g., S1–S2) and flush paved (e.g., S3–S9), influencing turning maneuvers and midblock encroachments. Overall, the table highlights that intersections serve as near-miss-prone nodes with concentrated but lower-volume interactions, whereas segments provide richer samples of continuous traffic flow, characterized by greater geometric variability.
Building on the spatial segmentation and detailed subregion profiles outlined above, the next step involves quantifying dynamic safety conditions using the proposed 2D-TTC framework (Algorithm 1). For each block, extreme near-miss events were identified by forward simulating vehicle trajectories under the geometric and traffic conditions described in Table 2. Both V–V and V–I interactions were modeled to reflect the heterogeneous near-miss environments of intersections and segments. Trajectories were projected over a 3s horizon, and spatial overlaps with surrounding traffic or roadway boundaries were detected as potential near-miss events. Representative outcomes from this process are shown in Fig. 4.2, illustrating the spatial dynamics of predicted V–V and V–I near-misses within the Miami study corridor. These event visualizations provide the empirical basis for block-level 2D-TTC extraction across the 10 intersections (I1–I10), enabling localized evaluation of crash-prone conditions.
Table 3: Summary statistics for V–V and V–I near-misses at Intersection
Metric V–V V–I
Avg Min Max Std Avg Min Max Std
Continuous variables
TTC 1.617 0.11 2.99 0.863 1.304 0.10 2.99 0.889
Relative speed 5.257 0.00 23.724 4.446 8.753 0.00 22.276 4.035
Relative acceleration -0.535 -13.706 10.343 2.227 0.161 -5.925 13.199 0.702
Relative distance 11.673 2.419 46.857 6.844 1.520 0.60 3.80 1.110
Jerk -5.550 -13.609 9.315 2.652 -7.900 -10.800 2.800 14.700
Heading difference 1.077 0.00 6.251 1.410 -0.045 -3.138 3.141 1.647
Steering difference 0.409 0.00 3.142 0.666 -0.006 -1.571 1.571 0.130
Volume 13.462 2.00 31.00 5.969 10.461 1.00 31.00 5.392
Vehicle length 4.78 3.18 6.92 0.49 5.1 3.95 12.65 1.30
Vehicle width 2.22 1.76 3.16 0.30 2.04 1.36 3.08 0.23
Categorical variable: Near-miss location (Proportion %)
Through 72.29 0.00
Left Turn 25.99 46.08
Right Turn 8.90 53.92
(a) V-V
(b) V-I
Fig. 6.: Representative intersection-level near-miss detection using the 2D-TTC framework: (a) V–V and (b) V–I
Descriptive statistics of intersection-level V–V and V–I near-misses are summarized in Table 3. The continuous indicators include 2D-TTC, relative speed, acceleration, distance, jerk, and angular deviations (heading and steering), along with vehicle dimensions and exposure (volume). On average, 2D-TTC values for both near-miss types range between 1.3 and 1.6 s, with extreme cases falling below 0.1 s, reflecting highly critical near-miss conditions. V–I near-misses are associated with higher relative speeds and tighter distance gaps, indicating more abrupt encroachment against roadway boundaries. Conversely, V–V near-misses exhibit larger heading and steering deviations, along with higher jerk variability, suggesting unstable maneuvers such as sudden turns or evasive swerves. The categorical distribution of near-misses further illustrates key behavioral differences. Nearly three-quarters (72%) of V–V events occurred during through movements, with fewer during left (26%) or right (9%) turns. In contrast, V–I near-misses were almost entirely turning-related, with 46% occurring in left turns and 54% in right turns, and none during through movements. This pattern highlights that infrastructure-related near-misses typically occur during lateral maneuvers at intersections, where vehicles cross lane boundaries or encroach on curbs and medians.
Beyond intersections, the analysis was extended to directional roadway segments (S1–S9) using the same 2D-TTC framework (See Fig. 4.2).Table 4 summarizes descriptive statistics of both V–V and V–I near misses at the segment level. For V–V near-misses, average TTC values were longer (1.83 s) than those observed at intersections, suggesting greater temporal headway between vehicles in segment settings. However, the extreme minima (0.11 s) and high maximum relative speeds (up to 33.9 m/s) highlight the presence of aggressive overtaking and high-speed convergence. Relative distances also exhibited a much broader range (up to 141.5 m), indicating that near-misses often arise in more dispersed longitudinal interactions rather than tightly constrained intersection maneuvers.
Table 4: Summary statistics for V–V and V–I near-misses at segments
Metric V–V V–I
Avg Min Max Std Avg Min Max Std
Continuous variables
TTC 1.829 0.11 2.99 0.805 1.272 0.10 2.99 0.912
Relative speed 4.513 0.00 33.909 4.278 10.262 0.000 28.067 4.212
Relative acceleration -0.871 -23.014 18.336 3.139 0.164 -14.067 23.343 1.400
Relative distance 15.361 1.931 141.548 12.773 1.720 0.32 6.56 1.02
Jerk -7.121 -17.502 8.423 1.552 -5.70 -12.72 3.556 4.77
Heading difference 0.368 0.000 6.256 0.990 -0.032 -3.102 3.138 1.586
Steering difference 0.577 0.000 3.142 0.756 -0.001 -1.571 1.571 0.167
Volume 16.801 1.000 65.000 11.009 12.816 1.000 65.000 7.906
Vehicle Length 4.74 3.0 7.85 1.11 4.79 2.95 8.65 0.87
Vehicle width 2.12 1.36 3.08 0.23 2.22 1.76 4 0.18
Lane number 2.73 2 4 0.578 2.51 2.0 4.0 0.52
Lane width 3.28 2.99 3.96 0.14 3.27 2.99 3.96 0.141
Driveway intensity 0.009 0.000 0.015 0.005 0.009 0.000 0.015 0.005
Categorical variable: (Proportion %)
Through 58.55 0.00
Collision location Left Turn 9.52 46.08
Right Turn 31.93 53.92
Median type Divided 23.60 13.98
Undivided 76.4 86.02
(a) V-V
(b) V-I
Fig. 7.: Representative segment-level near-miss detection using the 2D-TTC framework: (a) V–V and (b) V–I
For V–I near-misses, the mean TTC was shorter (1.27s), with relative speeds peaking at 28.1 m/s, underscoring the abrupt nature of vehicle–boundary encounters. Compared with intersections, segment-based V–I near-misses involved longer approach distances but also showed greater instability in jerk and heading, reflecting frequent lateral drifts toward lane edges or curbs. These events often occurred near driveways and minor access points, where vehicles deviated from their intended path, creating encroachment risks. The categorical breakdown reinforces these patterns. For V–V events, nearly 59% of near-misses were associated with through movements, while about 32% and 10% were linked to right and left turns, respectively. In contrast, V–I near-misses were dominated by turning maneuvers, with more than half (53.9%) linked to right turns and 46% to left turns, and no near-misses detected during through movements. This distribution highlights the heightened role of lateral maneuvers in producing infrastructure-related risks. Furthermore, median type also influenced outcomes: 76% of V–V near-misses and 86% of V–I near-misses occurred along undivided sections, underscoring the vulnerability of open medians and flush-paved designs in segment environments.
Taken together, these segment-level findings reveal that while intersections are hotspots for tightly constrained V-V near-misses, segments carry distinctive risks tied to high-speed encounters and infrastructure encroachments. The observed distributions underscore the importance of treating intersections and segments as functionally distinct safety domains when modeling near-miss dynamics and designing targeted countermeasures.
(a) Intersection
(b) Segment: NB
(c) Segment: SB
Fig. 8.: BM near-miss frequencies by location type and direction
After identifying and compiling the BM near-miss events for each observation unit (segments and intersections), the frequencies of extreme V–V and V–I events were systematically aggregated and visualized in Fig. 4.2. This plot illustrates the spatial distribution of extreme near-miss frequencies across 28 locations, comprising 10 intersections and 18 segments, organized into intersection zones, primary-direction segments, and opposite-direction segments. This spatial aggregation confirms the integrity of the BM dataset and provides the foundation for modeling. The study extends prior approaches by incorporating both V–V and V–I near misses into the HBSGRP–UGEV framework, explicitly integrating roadway geometry and infrastructure features alongside vehicle dynamics. Prior research has emphasized the role of vehicle kinematics in crash risk (Desai et al., 2021; Jun et al., 2007; Tak et al., 2015); here, additional geometric covariates, such as lane configuration and boundary proximity, are introduced. For example, high steering angles, often indicating abrupt lateral maneuvers—were included as predictors, given their link to boundary incursions and sudden evasive actions that may elevate near-miss risk (Gilbert et al., 2021). Fig. 4.2 highlights several notable spatial patterns. Opposite-direction segments (S82, S72, S62) exhibited the highest V–I frequencies, suggesting greater lateral variability and exposure to curb or median near-misses. Primary-direction segments (S11, S21, S81) showed elevated V–V near-miss frequencies, consistent with midblock environments dominated by longitudinal interactions such as close-following, lane changes, or overtaking. Among intersections, I7, I8, and I9 emerged as critical hotspots, with elevated frequencies of both near-miss types. This dual concentration reflects the operational complexity of intersections, where near-missing maneuvers, pedestrian flows, and signal control interact to amplify crash potential.
The distribution and magnitude of near-miss counts across all locations validate the effectiveness of the BM approach in capturing rare but safety-critical events. Specifically, nearly all study sites meet or exceed the recommended minimum threshold of 30 BM events for reliable GEV-based inference (Zheng et al., 2014), with the exception of intersection I2. This ensures that the dataset is statistically robust for subsequent HBSGRP–UGEV analysis. The spatial heterogeneity revealed in Fig. 4.2 further underscores the need to incorporate direction-specific geometric and behavioral covariates within the modeling framework. These results provide compelling justification for developing localized, high-resolution safety models that integrate both roadway configuration and observed vehicle dynamics, thereby enabling precise and context-sensitive real-time risk assessments in complex urban environments.
5 Modeling estimation result
This study develops a series of Bayesian hierarchical UGEV models to estimate COR from extreme 2D-TTC near-miss events observed along Biscayne Boulevard in Miami. Extreme near-misses are identified using the BM approach, wherein the minimum 2D-TTC within each temporal block is treated as the most critical interaction. The theoretical foundations and implementation of this sampling methodology are detailed in prior research (2025b). Each V–V and V–I interaction is treated as an independent block. Because extreme events are sparse at individual sites, direct location-specific estimation can yield unstable posteriors with wide uncertainty intervals. To address this, a grouping strategy was adopted, pooling samples across functionally similar sites (intersections, segments) while still preserving local heterogeneity. This clustering approach improves estimation stability and aligns with prior network-level applications of UGEV models in traffic safety (Zheng et al., 2014; Fu and Sayed, 2021c).
The primary analytical framework is the HBSGRP–UGEV model, while a simpler fixed-effect hierarchical model (HBSFP) was estimated as a benchmark. Accordingly, two models were developed for each near-miss type: (1) HBSFP and (2) HBSGRP. The grouped random parameter specification allows selected coefficients to vary across spatial clusters, thereby capturing latent differences linked to roadway geometry and interaction dynamics. Model performance was evaluated using the DIC (Spiegelhalter et al., 2002; El-Basyouny and Sayed, 2012b, a), where lower values indicate stronger predictive performance (Zheng et al., 2019).
Estimation was conducted using MultiBUGS, a parallelized extension of OpenBUGS tailored for high-dimensional Bayesian inference. Its support for block updating and parallel execution makes it well-suited for capturing spatial dependence, random effects, and extreme value behavior within complex data structures. Parameters were estimated by Markov Chain Monte Carlo (MCMC) simulation, two parallel MCMC chains of 50000 iterations were run, with the first 20000 discarded as burn-in and the remaining 30000 retained for posterior inference. Convergence was verified both visually (trace plots) and quantitatively (Brooks–Gelman–Rubin diagnostic), with all monitored parameters achieving BGR values below 1.1, indicating satisfactory convergence (Gelman and Rubin, 1992; Brooks and Gelman, 1998; El-Basyouny and Sayed, 2009). This estimation procedure yielded robust posterior summaries of means, standard deviations, and 95% credible intervals, even under correlation structures induced by spatial clustering and random effects.
The study incorporated a comprehensive set of covariates into the location ($\mu$) and scale ($\sigma$) parameters of the non-stationary UGEV distribution to reflect contextual and behavioral variability in COR. For both V–V and V–I interactions, dynamic covariates included relative speed, distance, acceleration, deceleration, headway difference, steering angle difference, jerk and traffic volume, factors capturing key dimensions of vehicle dynamics and near-miss severity. Geometric and infrastructure related covariates were also included: lane number, lane width, driveway intensity, median type, and lane position for segment models; and a categorical indicator of turning movement for intersection models.
All covariates were specified in both $\mu$ and $\sigma$ components, enabling the models to capture variation in both the central tendency and dispersion of extreme 2D-TTC. In contrast, the shape parameter ($\xi$) was held constant across all locations. This choice is deliberate: $\xi$ governs tail heaviness in the UGEV distribution but is notoriously difficult to estimate from sparse BM samples. Allowing $\xi$ to vary with covariates often induces instability and overfitting (Coles et al., 2001; Cooley et al., 2006). Fixing $\xi$ improves convergence and yields more interpretable posterior distributions.
5.1 Posterior results and model performance
Tables 5 through 8 report the posterior summaries for both HBSFP and HBSGRP models, estimated separately for V–V and V–I near-misses at intersections and segments. Across all cases, the HBSGRP specification consistently outperformed the HBSFP baseline, confirming its ability to capture unobserved heterogeneity and locatione-level variability. The DIC values for the HBSGRP models were 508.9 (V–V) and 3179.6 (V–I) at intersections, and 4189.0 (V–V) and 6550.0 (V–I) at segments. These correspond to relative reductions in DIC of 3.41% and 0.54% at intersections, and 7.49% and 3.11% at segments, compared with their HBSFP counterparts. Such reductions exceed the widely accepted threshold of five DIC units, indicating statistically meaningful improvements in fit (El-Basyouny and Sayed, 2012b, a). To further assess model robustness, WAIC and LOOIC were also computed. Both measures produced rankings consistent with DIC, reinforcing the superior performance of the HBSGRP structure. By allowing selected coefficients to vary across spatial groups, the HBSGRP models capture latent heterogeneity in both the central tendency and dispersion of near-miss severity, which fixed-effect structures fail to represent. Collectively, these results demonstrate that grouped random parameter models provide a more flexible and accurate framework for modeling COR dynamics in complex urban traffic environments.
5.1.1 Intersection-level
The intersection-level HBSGRP–UGEV model was designed to capture the heterogeneity of extreme 2D-TTC distributions arising from V-V interactions at signalized locations. In the model specification stage, several covariates were tested as random effects; however, most roadway geometry variables exhibited negligible variation across sites and did not improve posterior fit. To maintain parsimony and interpretability, random effects were therefore restricted to the intercept and two dynamic covariates (relative speed and relative distance), while other vehicle dynamics and geometric features were retained as fixed effects, following recommendations from (Anis et al., 2025a). This choice reflects theoretical expectations: vehicle dynamics are highly variable across intersections depending on signal phasing, approach alignment, and driver behavior, whereas roadway structure tends to exert a more stable influence. The resulting specification captures both site-specific dynamics and broader geometric constraints, group-specific random effects illustrated in Fig. 5.1.1 with posterior summaries reported in Table 5.
(a) $\beta_{\mu,0[1,10]}$
(b) $\gamma_{\mu,1[1,10]}$
(c) $\gamma_{\mu,2[1,10]}$
(d) $\beta_{\sigma,0[1,10]}$
(e) $\gamma_{\sigma,1[1,10]}$
(f) $\gamma_{\sigma,2[1,10]}$
(g) $\beta_{\xi,0[1,10]}$
Fig. 9.: Posterior estimates for V-V near-misses at intersections
Table 5: Posterior estimates for V-V near misses at intersections
Model Parameters Hyperparameter Covariate HBSFP HBSGRP
Mean SDa 95% CRIb Mean SDa 95% CRIb
Location Parameter ($\mu_{k,i}$)
$\beta_{\mu,0}$ Fixed Intercept -1.870 0.039 $1.917,-1.811]^{\dagger} - - -$
$\beta_{\mu,0,k}$ Random Intercept - - - -1.893 0.186 $2.309,-1.572]^{\dagger}$
$\gamma_{\mu,1,k}$ Random Relative speed - - - 0.745 0.181 $0.393,1.134]^{\dagger}$
$\gamma_{\mu,2,k}$ Random Relative distance - - - -0.491 0.157 $0.819,-0.280]^{\dagger}$
$\beta_{\mu,1}$ Fixed Relative speed 0.639 0.050 $0.246,0.739]^{\dagger} - - -$
$\beta_{\mu,2}$ Fixed Relative acceleration 1.064 0.074 $0.126,2.018]^{\dagger} -$0.155 0.032 $0.221,-0.093]^{\dagger}$
$\beta_{\mu,3}$ Fixed Relative deceleration -0.013 0.037 [-0.088, 0.059] 0.140 0.063 $0.004,0.238]^{\dagger}$
$\beta_{\mu,4}$ Fixed Relative distance -0.442 0.036 $0.505,-0.418]^{\dagger} - - -$
$\beta_{\mu,5}$ Fixed Jerk -0.012 0.006 $021,-0.002]^{\dagger} -0.0$27 0.015 $0.046,-0.006]^{\dagger}$
$\beta_{\mu,6}$ Fixed Heading difference 0.024 0.026 [-0.029, 0.072] 0.028 0.039 [-0.052, 0.097]
$\beta_{\mu,7}$ Fixed Steering difference -0.004 0.049 [-0.096, 0.092] 0.059 0.026 $0.003,0.107]^{\dagger}$
$\beta_{\mu,8}$ Fixed Volume -0.003 0.005 [-0.012, 0.006] -0.018 0.035 [-0.085, 0.051]
$\beta_{\mu,9}$ Fixed Turn movement (Left =1, else=0) 0.086 0.026 [-0.102, 0.242] 0.073 0.011 $0.042,0.205]^{\dagger}$
Scale Parameter ($\log\sigma_{k,i}$)
$\beta_{\sigma,0}$ Fixed Intercept -0.571 0.058 $0.680,-0.455]^{\dagger} - - -$
$\beta_{\sigma,0,k}$ Random Intercept - - - -0.599 0.311 $1.240,-0.056]^{\dagger}$
$\gamma_{\sigma,1,k}$ Random Relative speed - - - -0.265 0.254 $0.810,-0.043]^{\dagger}$
$\gamma_{\sigma,2,k}$ Random Relative distance - - - -0.653 0.297 $1.121,-0.033]^{\dagger}$
$\beta_{\sigma,1}$ Fixed Relative speed 0.012 0.084 [-0.139, 0.183] - - -
$\beta_{\sigma,2}$ Fixed Relative acceleration -1.686 0.865 $3.066,-0.627]^{\dagger} -$0.226 0.066 $0.360,-0.108]^{\dagger}$
$\beta_{\sigma,3}$ Fixed Relative deceleration -0.218 0.051 $0.321,-0.120]^{\dagger} -$0.253 0.053 $0.356,-0.151]^{\dagger}$
$\beta_{\sigma,4}$ Fixed Relative distance -0.062 0.008 $0.078,-0.046]^{\dagger} - - -$
$\beta_{\sigma,5}$ Fixed Jerk 0.019 0.009 $005,0.030]^{\dagger} 0.0$12 0.003 $0.007,0.018]^{\dagger}$
$\beta_{\sigma,6}$ Fixed Heading difference -0.063 0.071 [-0.212, 0.060] -0.055 0.014 $0.133,-0.026]^{\dagger}$
$\beta_{\sigma,7}$ Fixed Steering difference 0.063 0.065 [-0.061, 0.191] -0.053 0.040 $0.117,-0.011]^{\dagger}$
$\beta_{\sigma,8}$ Fixed Volume -0.001 0.007 [-0.015, 0.013] -0.004 0.044 [-0.090, 0.084]
$\beta_{\sigma,9}$ Fixed Turn movement(Left =1, else=0) 0.042 0.120 [-0.183, 0.285] 0.217 0.123 $0.025,0.471]^{\dagger}$
Shape Parameter ($\xi_{k,i}$)
$\beta_{\xi,0}$ Fixed Intercept -0.305 0.051 $0.413,-0.208]^{\dagger} - - -$
$\beta_{\xi,0,k}$ Random Intercept - - - -0.684 0.253 $1.309,-0.314]^{\dagger}$
Model Fit
DIC 526.9 508.9
WAIC 539.4 521.5
LOOIC 540.1 522.3
  • HBSFP
    Hierarchical Bayesian Spatial Fixed Parameter
  • HBSGRP
    Hierarchical Bayesian Spatial Grouped Random Parameter
  • a
    Standard deviation
  • b
    95% Bayesian credible interval
    Covariate not included in the model
  • $\dagger$}
    Indicates statistical significance at the 95% level (interval excludes 0)
The fixed-parameter benchmark (HBSFP) provides a baseline against which the grouped random specification can be compared. Consistent reductions in DIC, WAIC, and LOOIC confirm the superior fit of the HBSGRP model, underscoring the explanatory value of allowing relative speed and spacing effects to vary across intersection clusters. The direction and magnitude of these coefficients provide valuable insights into the mechanisms underlying COR. The random-parameter specification highlights the spatially varying influence of relative speed and relative distance on both the location and scale components of the HBSGRP distribution. For $\mu$, relative speed shows a positive effect across intersections, indicating that larger speed differentials between interacting vehicles lead to sharply reduced 2D-TTC values and more severe near-miss conditions. In contrast, relative distance is negatively associated with $\mu$, meaning that shorter headways systematically increase near-miss severity by limiting the buffer available for evasive action. For $\sigma$, both covariates are associated with reduced variability in near-miss severity. Relative speed has a negative coefficient, suggesting that high-speed interactions produce consistently critical outcomes, dangerous but less variable, because drivers have little opportunity to recover through braking or swerving. Similarly, relative distance exerts a negative effect on $\sigma$, reflecting that short headways yield uniformly critical near-miss with little dispersion in outcomes. Together, these results confirm the physical intuition that speed amplifies severity while distance compresses variability, producing deterministic near miss outcomes under constrained conditions. Fig. 5.1.1 further illustrate this spatial heterogeneity across intersection, while most locations show consistent trends, the I2 site exhibits unusually wide CI for the $\mu$ effects and CI exceed zero for the $\sigma$, across intercept, relative speed, and relative distance. This instability is directly attributable to I2’s small BM sample size, which weakens posterior precision and introduces greater uncertainty in estimated group-level effects. Despite this, the overall pattern across intersections remains robust: relative speed consistently intensifies near-miss severity, while shorter distances exacerbate risk and constrain outcome variability.
Beyond the random effects, several covariates were modeled as fixed parameters due to their relatively stable contributions across intersections. Their interpretation reveals distinct roles of longitudinal dynamics, lateral control, maneuver type, and exposure effects in shaping near-miss severity and variability. Table 5 indicate that relative acceleration shows a significant negative effect on both the location and scale parameters. This suggests that acceleration is generally associated with smoother and less variable near-miss, consistent with situations where vehicles accelerate out of a green phase or clear an intersection without abrupt interaction. By contrast, relative deceleration exerts a positive effect on severity ($\mu$) but a negative effect on variability ($\sigma$). This indicates that abrupt braking sharply intensifies near-misses but also produces more uniform outcomes, reflecting the deterministic nature of emergency deceleration in high-risk encounters. Jerk, which captures the rate of change in acceleration, has the opposite pattern: a negative effect on $\mu$ but a positive effect on $\sigma$. Although jerk slightly reduces average near-miss severity (perhaps due to hesitation or driver corrections), it simultaneously increases unpredictability in outcomes, highlighting the unstable nature of rapid throttle/brake oscillations.
Lateral vehicle dynamics also contribute meaningfully to near-miss risk estimation. Steering difference emerges as a significant contributor in the HBSGRP specification, showing a positive association with $\mu$ and a negative association with $\sigma$. This suggests that abrupt steering maneuvers intensify near-miss severity but lead to more consistent outcomes within constrained intersection geometries. Heading difference also shows a significant positively association with $\mu$ and a negatively with $\sigma$. Indicating that even small misalignments in trajectory lead to predictable patterns of near-miss outcomes, consistent with the rigid geometry of signalized approaches.
The categorical indicator for left-turn movements is significant and has a positive effect on both $\mu$ and $\sigma$. This aligns with the well-documented complexity of left-turn phases, which involve interactions with opposing through traffic and diverse driver decision strategies. The result suggests that left turns not only elevate near-miss severity but also introduce greater variability in outcomes, reflecting the wide range of intersection designs and driver behaviors under turning conditions. Traffic volume is statistically insignificant in both $\mu$ and $\sigma$. Nonetheless, it was retained in the model to ensure conceptual completeness, since it represents exposure. The lack of effect implies that short-duration 2D-TTC near-miss events are governed more by instantaneous vehicle interactions than by aggregate traffic demand.
Finally, the estimated shape parameter ($\xi$) is negative and significant across all intersections (Fig. 8(g)), indicating a bounded upper tail of the UGEV distribution. This supports the Weibull domain as the appropriate tail form for modeling extreme near-misses in signalized urban settings. Importantly, group-level variability in $\xi$ suggests that intersection design and control strategies influence not only the frequency but also the extremal nature of near-miss outcomes.
The HBSGRP model for V–I near-misses specifies the intercept as a random parameter in both the location and scale components, capturing baseline heterogeneity in near-miss severity and variability across intersections. Relative speed (location) and relative distance (scale) were also treated as random effects (Fig. 5.1.1), though their 95% CI overlapped zero. They were retained on theoretical grounds, given their inclusion improve overall metrics. All other covariates were modeled as fixed effects, representing more stable behavioral and geometric influences. Fit statistics (Table 6) show modest but consistent improvements in DIC, WAIC, and LOOIC over the HBSFP benchmark, suggesting that the grouped random parameter structure adds nuance but yields smaller performance gains compared to the V–V case.
Table 6: Posterior estimates for V-I near misses at intersections
Model Parameters Hyperparameter Covariate HBSFP HBSGRP
Mean SDa 95% CRIb Mean SDa 95% CRIb
Location Parameter ($\mu_{k,i}$)
$\beta_{\mu,0}$ Fixed Intercept -1.392 0.039 $1.461,-1.299]^{\dagger} - - -$
$\beta_{\mu,0,k}$ Random Intercept - - - -1.401 0.092 $1.598,-1.236]^{\dagger}$
$\gamma_{\mu,1,k}$ Random Relative speed - - - 0.025 0.020 [-0.012, 0.065]
$\beta_{\mu,1}$ Fixed Relative speed -0.012 0.022 [-0.054, 0.031] - - -
$\beta_{\mu,2}$ Fixed Relative acceleration -0.020 0.023 [-0.067, 0.022] -0.020 0.024 [-0.068, 0.029]
$\beta_{\mu,3}$ Fixed Relative deceleration -0.066 0.013 $075,-0.055]^{\dagger} -0.0$26 0.011 $0.071,-0.009]^{\dagger}$
$\beta_{\mu,4}$ Fixed Relative distance -0.357 0.035 $0.427,-0.290]^{\dagger} -$0.359 0.026 $0.405,-0.306]^{\dagger}$
$\beta_{\mu,5}$ Fixed Jerk 0.270 0.355 [-0.457, 0.556] -0.021 0.005 [-0.026, 0.01]
$\beta_{\mu,6}$ Fixed Heading difference 0.040 0.022 $005,0.084]^{\dagger} 0.0$62 0.026 $0.010,0.127]^{\dagger}$
$\beta_{\mu,7}$ Fixed Steering difference 0.009 0.018 [-0.027, 0.045] 0.038 0.022 $0.014,0.074]^{\dagger}$
$\beta_{\mu,8}$ Fixed Volume -0.017 0.021 [-0.061, 0.023] -0.023 0.023 [-0.069, 0.027]
$\beta_{\mu,9}$ Fixed Turn movement (Left =1, else=0) -0.129 0.080 $0.265,-0.061]^{\dagger} -$0.579 0.310 $1.231,-0.220]^{\dagger}$
Scale Parameter ($\log\sigma_{k,i}$)
$\beta_{\sigma,0}$ Fixed Intercept -0.100 0.038 $0.167,-0.009]^{\dagger} - - -$
$\beta_{\sigma,0,k}$ Random Intercept - - - -0.197 0.098 $0.374,-0.001]^{\dagger}$
$\gamma_{\sigma,1,k}$ Random Relative distance - - - 0.042 0.034 [-0.028, 0.107]
$\beta_{\sigma,1}$ Fixed Relative speed 0.036 0.019 $013,0.063]^{\dagger} 0.0$40 0.019 $0.001,0.076]^{\dagger}$
$\beta_{\sigma,2}$ Fixed Relative acceleration 0.012 0.015 [-0.015, 0.042] 0.046 0.011 $0.009,0.084]^{\dagger}$
$\beta_{\sigma,3}$ Fixed Relative deceleration -0.147 0.056 $0.285,-0.027]^{\dagger}$g -0.177 0.056 $0.289,-0.047]^{\dagger}$
$\beta_{\sigma,4}$ Fixed Relative distance 0.209 0.038 $0.140,0.282]^{\dagger} - - -$
$\beta_{\sigma,5}$ Fixed Jerk 0.611 0.133 $0.227,0.886]^{\dagger}$ 0.109 0.033 $0.016,0.156]^{\dagger}$
$\beta_{\sigma,6}$ Fixed Heading difference -0.023 0.018 $057,-0.004]^{\dagger} -0.0$90 0.021 $0.130,-0.049]^{\dagger}$
$\beta_{\sigma,7}$ Fixed Steering difference -0.012 0.018 [-0.045, 0.023] -0.024 0.020 $0.062,-0.017]^{\dagger}$
$\beta_{\sigma,8}$ Fixed Volume 0.004 0.016 [-0.028, 0.034] 0.029 0.018 $0.003,0.064]^{\dagger}$
$\beta_{\sigma,9}$ Fixed Turn movement (Left =1, else=0) 0.100 0.053 $0.019,0.197]^{\dagger}$ 0.264 0.049 $0.163,0.349]^{\dagger}$
Shape Parameter ($\xi_{k,i}$)
$\beta_{\xi,0}$ Fixed Intercept -0.677 0.042 $0.749,-0.594]^{\dagger} - - -$
$\beta_{\xi,0,k}$ Random Intercept - - - -0.680 0.096 $0.876,-0.501]^{\dagger}$
Model Fit
DIC 3197 3179.6
WAIC 3202 3194.4
LOOIC 3212.87 3209.8
  • HBSFP
    Hierarchical Bayesian Spatial Fixed Parameter
  • HBSGRP
    Hierarchical Bayesian Spatial Grouped Random Parameter
  • a
    Standard deviation
  • b
    95% Bayesian credible interval
    Covariate not included in the model
  • $\dagger$}
    Indicates statistical significance at the 95% level (interval excludes 0)
(a) $\beta_{\mu,0[1,10]}$
(b) $\beta_{\sigma,0[1,10]}$
(c) $\beta_{\xi,0[1,10]}$
Fig. 10.: Posterior estimates for V-I near-misses at intersections
In the location parameter, several interpretable effects emerge. Relative distance is strongly negative and significant, confirming that reduced clearance to the roadway edge sharply lowers 2D-TTC, elevating severity. Relative deceleration also carries a significant negative sign, which contrasts with its positive effect in V–V near-miss. This suggests that braking near infrastructure reflects proactive yielding or evasive action, thereby limiting severity rather than escalating it. Heading and steering differences are positive and significant, indicating that lateral deviations (e.g., turning toward or away from curbs) increase near-miss severity. Left-turn movement, however, shows a significant negative association, suggesting that left-turn trajectories may provide greater clearance to boundaries, reducing average severity compared to right turns. This finding requires careful interpretation, as it may partly reflect geometric factors such as wider turning radii or reduced pedestrian/boundary overlap in left-turn phasing. Relative speed, though included as a random effect, is not significant in the pooled analysis, highlighting that V–I risks are less sensitive to speed differentials than V–V encounters.
The scale parameter results add further insight into variability. The negative random intercept suggests baseline differences across intersections, with some sites producing more predictable outcomes than others. Relative speed exhibits a positive and significant association with scale, implying that faster approaches to infrastructure yield greater variability in severity—contrasting with the V–V case, where speed reduced dispersion. This difference reflects the inherently diverse driver responses to boundary proximity: some brake aggressively, while others take wider trajectories or maintain speed. Relative acceleration is also positive and significant, reinforcing the idea that heterogeneous driver behavior during acceleration near boundaries widens outcome variability. In contrast, relative deceleration has a significant negative effect on scale, meaning that strong braking compresses the severity distribution into consistently critical outcomes. Jerk is similarly positive, confirming that abrupt longitudinal control amplifies variability. Lateral dynamics again exhibit distinct roles: heading and steering differences reduce variability, suggesting that once a driver commits to a lateral maneuver, outcomes become more predictable. Finally, left-turn movement increases variability despite lowering average severity, consistent with heterogeneous geometric layouts and driver strategies in left-turn contexts.
The shape parameter remains negative and significant across intersections, supporting the Weibull domain for V–I near misses and confirming that these events have a bounded upper tail. This implies a physical or behavioral limit to how severe boundary near-misses can become, constrained by roadway geometry and driver response capacity. Spatial variation in the shape parameter underscores the role of intersection design and infrastructure placement in shaping the extremal properties of near-miss distributions. Overall, the intersection-level results suggest that V–I near-misses are driven more by clearance, braking behavior, and lateral maneuvers than by speed differentials. Unlike V–V near-misses, where high speeds produce deterministic severity, V–I cnear-misses are marked by greater variability, reflecting the complex interplay between infrastructure geometry, driver avoidance strategies, and intersection control.
5.2 Segment-level
The HBSGRP model for segment-level V–V near misses introduces random parameters within a hierarchical framework (Table 7). Allowing the intercept and three time-varying vehicle-based measures, relative speed, relative deceleration, and relative distance, to vary across segments (See Fig. 5.2). This specification acknowledges that longitudinal dynamics in uninterrupted midblock contexts are shaped not only by vehicle behavior but also by geometric features such as lane configuration, median presence, and driveway density. All remaining covariates, including relative acceleration, jerk, heading and steering differences, traffic volume, lane attributes, and access density, were modeled as fixed effects due to their comparatively stable influence across segments. The HBSGRP specification produces substantial improvements in model fit over the HBSFP baseline, confirming the value of partial pooling in capturing heterogeneity across the nine directional segments analyzed.
For the location parameter ($\mu$), random effects highlight systematic heterogeneity across the nine directional segments. Relative speed exerts a significant positive effect, confirming that higher closing speeds consistently compress available 2D-TTC and elevate near-miss severity. However, the posterior distributions in Fig. 5.2 reveal that the magnitude of this effect varies substantially across segments. Segments with uninterrupted, high-speed flows (e.g., S11) show stronger effects, while others with denser access points (e.g., S6, S7 and S13) exhibit wider credible intervals. Relative deceleration also carries a negative sign, though with more overlap in its credible intervals across segments, suggesting that emergency braking may sometimes mitigate severity in smoother-flow environments, while in higher-speed segments it intensifies near-misses. Notably, in some lower-sample segments, credible intervals for the location intercept approach or cross zero, reflecting weaker estimation stability under sparse BM events. Relative distance displays a consistently negative effect across segments, meaning that shorter headways reliably reduce 2D-TTC, yet its posterior variation underscores differences in spacing behavior along divided versus undivided facilities.
Refer to caption Refer to caption
(a) $\beta_{\mu,0[1,18]}$ (b) $\gamma_{\mu,1[1,18]}$
Refer to caption Refer to caption
(c) $\gamma_{\mu,2[1,18]}$ (d) $\gamma_{\mu,3[1,18]}$
Refer to caption Refer to caption
(e) $\beta_{\sigma,0[1,18]}$ (f) $\gamma_{\sigma,1[1,18]}$
Refer to caption Refer to caption
(g) $\gamma_{\sigma,2[1,18]}$ (h) $\gamma_{\sigma,3[1,18]}$
(i) $\beta_{\xi,0[1,18]}$
Fig. 11.: Posterior estimates for V-V near-misses at segments
Table 7: Posterior estimates V-V near misses at segments
Model Parameters Hyperparameter Covariate HBSFP HBSGRP
Mean SDa 95% CRIb Mean SDa 95% CRIb
Location Parameter ($\mu_{k,i}$)
$\beta_{\mu,0}$ Fixed Intercept -2.117 0.032 $2.181,-2.025]^{\dagger} - - -$
$\beta_{\mu,0,k}$ Random Intercept - - - -1.047 0.171 $1.369,-0.772]^{\dagger}$
$\gamma_{\mu,1,k}$ Random Relative speed - - - 0.406 0.068 $0.286,0.543]^{\dagger}$
$\gamma_{\mu,2,k}$ Random Relative deceleration - - - -0.082 0.016 $0.197,-0.057]^{\dagger}$
$\gamma_{\mu,3,k}$ Random Relative distance - - - -0.600 0.082 $0.779,-0.454]^{\dagger}$
$\beta_{\mu,1}$ Fixed Relative speed 0.446 0.037 $0.376,0.519]^{\dagger} - - -$
$\beta_{\mu,2}$ Fixed Relative acceleration -0.068 0.017 $154,-0.034]^{\dagger} -0.0$79 0.037 $0.163,-0.032]^{\dagger}$
$\beta_{\mu,3}$ Fixed Relative deceleration -0.074 0.018 $0.108,-0.040]^{\dagger} - - -$
$\beta_{\mu,4}$ Fixed Relative distance -0.669 0.063 $0.801,-0.552]^{\dagger} - - -$
$\beta_{\mu,5}$ Fixed Jerk -0.004 0.001 $0.007,-0.002]^{\dagger} -$0.156 0.043 $0.234,-0.101]^{\dagger}$
$\beta_{\mu,6}$ Fixed Heading difference 0.043 0.016 $011,0.073]^{\dagger} 0.0$44 0.008 $0.029,0.059]^{\dagger}$
$\beta_{\mu,7}$ Fixed Steering difference 0.042 0.012 $018,0.066]^{\dagger} 0.0$26 0.010 $0.007,0.046]^{\dagger}$
$\beta_{\mu,8}$ Fixed Volume 0.035 0.013 $009,0.060]^{\dagger} 0.0$19 0.008 $0.002,0.038]^{\dagger}$
$\beta_{\mu,9}$ Fixed Lane no -0.006 0.011 [-0.027, 0.015] 0.009 0.011 [-0.011, 0.032]
$\beta_{\mu,10}$ Fixed Lane width 0.003 0.001 $001,0.007]^{\dagger} 0.0$17 0.009 $0.002,0.031]^{\dagger}$
$\beta_{\mu,11}$ Fixed Driveway density -0.001 0.015 [-0.029, 0.027] 0.489 0.114 $0.256,0.578]^{\dagger}$
$\beta_{\mu,12}$ Fixed Median (Undivided=1, else=0) -0.004 0.033 [-0.070, 0.062] -0.412 0.155 $0.711,-0.070]^{\dagger}$
$\beta_{\mu,13}$ Fixed Vehicle position (Left lane=1, else=0) 0.029 0.038 [-0.048, 0.103] 0.055 0.006 $0.012,0.105]^{\dagger}$
Scale Parameter ($\log\sigma_{k,i}$)
$\beta_{\mu,0}$ Fixed Intercept -0.622 0.045 $0.711,-0.533]^{\dagger} - - -$
$\beta_{\mu,0,k}$ Random Intercept - - - -0.253 0.047 $0.952,-0.156]^{\dagger}$
$\gamma_{\sigma,1,k}$ Random Relative speed - - - -0.055 0.026 $0.107,-0.003]^{\dagger}$
$\gamma_{\sigma,2,k}$ Random Relative deceleration - - - -0.119 0.018 $0.154,-0.084]^{\dagger}$
$\gamma_{\sigma,3,k}$ Random Relative distance - - - -0.304 0.117 $0.520,-0.084]^{\dagger}$
$\beta_{\sigma,1}$ Fixed Relative speed -0.116 0.050 $0.217,-0.023]^{\dagger} - - -$
$\beta_{\sigma,2}$ Fixed Relative acceleration -0.611 0.097 $817,-0.439]^{\dagger} -0.0$24 0.173 [-0.277, 0.298]
$\beta_{\sigma,3}$ Fixed Relative deceleration -0.060 0.019 $0.097,-0.024]^{\dagger} - - -$
$\beta_{\sigma,4}$ Fixed Relative distance 0.003 0.064 [-0.121, 0.130] - - -
$\beta_{\sigma,5}$ Fixed Jerk 0.006 0.000 $0.004,0.008]^{\dagger}$ 0.357 0.147 $0.061,0.611]^{\dagger}$
$\beta_{\sigma,6}$ Fixed Heading difference 0.065 0.017 $2,0.099]^{\dagger} 0.033 0.0$11 $0.017,0.068]^{\dagger}$
$\beta_{\sigma,7}$ Fixed Steering difference 0.011 0.016 [-0.020, 0.043] -0.001 0.019 [-0.031, 0.029]
$\beta_{\sigma,8}$ Fixed Volume -0.011 0.016 [-0.043, 0.020] -0.012 0.015 [-0.030, 0.029]
$\beta_{\sigma,9}$ Fixed Lane no -0.031 0.014 $049,-0.005]^{\dagger} -0.0$28 0.017 $0.061,-0.002]^{\dagger}$
$\beta_{\sigma,10}$ Fixed Lane width -0.006 0.003 $012,-0.000]^{\dagger} -0.0$15 0.009 $0.044,-0.001]^{\dagger}$
$\beta_{\sigma,11}$ Fixed Driveway density 0.040 0.018 $0.004,0.075]^{\dagger}$ 0.284 0.215 $0.019,0.648]^{\dagger}$
$\beta_{\sigma,12}$ Fixed Median (Undivided=1, else=0) 0.076 0.042 [-0.006, 0.158] -0.205 0.095 $0.369,-0.058]^{\dagger}$
$\beta_{\sigma,13}$ Fixed Vehicle position (Left lane=1, else=0) -0.049 0.052 [-0.156, 0.051] -0.109 0.052 $0.209,-0.006]^{\dagger}$
Shape Parameter ($\xi_{k,i}$)
$\beta_{\xi,0}$ Fixed Intercept -0.116 0.030 $0.178,-0.059]^{\dagger} - - -$
$\beta_{\xi,0,k}$ Random Intercept - - - -0.185 0.144 $0.436,-0.163]^{\dagger}$
Model Fit
DIC 4528 4189
WAIC 4594 4276
LOOIC 4608.25 4278
  • HBSFP
    Hierarchical Bayesian Spatial Fixed Parameter
  • HBSGRP
    Hierarchical Bayesian Spatial Grouped Random Parameter
  • a
    Standard deviation
  • b
    95% Bayesian credible interval
    Covariate not included in the model
  • $\dagger$}
    Indicates statistical significance at the 95% level (interval excludes 0)
Among the fixed effects, relative acceleration shows a significant negative coefficient in both models, implying that acceleration phases are systematically linked to higher near-miss severity on segments. This contrasts with intersection contexts, where acceleration often reflected flow resumption. Here, aggressive acceleration during overtaking or merging maneuvers reduces spacing and increases risk. Jerk is also negative, reinforcing the role of abrupt longitudinal control in producing critical outcomes. Lateral control indicators—heading and steering differences—show positive signs, suggesting that lateral adjustments modestly increase 2D-TTC where lane-changing flexibility exists. Driveway density exhibits a strong positive effect on $\mu$, which may appear counterintuitive, but is likely explained by anticipatory driver caution in access-rich environments. Conversely, undivided medians sharply reduce $\mu$, indicating that exposure to opposing flows creates consistently severe interactions.
For the scale parameter ($\sigma$), the random effects for relative speed, deceleration, and distance are all negative, showing that extreme dynamics compress outcome variability, leading to uniformly critical near misses. The plots confirm this: high-speed and short-spacing conditions cluster tightly around lower 2D-TTC values, with narrow posteriors in segments such as S12 and S15. In contrast, segments like S10 display wider CI for random scale effects, again reflecting limited BM sample sizes. Fixed effects add further nuance: jerk and heading difference increase scale, suggesting that volatile acceleration or lateral corrections introduce heterogeneity in outcomes, while lane number and lane width reduce scale, indicating that greater maneuvering space fosters more predictable near-misses. Driveway density increases dispersion, consistent with the variability introduced by frequent merging and crossing maneuvers. Median type again exerts a strong negative influence, reinforcing that the absence of raised medians compresses severity outcomes into consistently critical ranges.
Finally, the shape parameter ($\xi$) remains negative and significant across segment groups, confirming a bounded upper tail consistent with the Weibull domain. This indicates physical and behavioral limits to maximum 2D-TTC values during critical segment near-misses, constrained by speed environment, geometry, and driver reaction times. Posterior variation in the shape intercept suggests that undivided segments experience steeper extremal limits, while divided facilities allow marginally greater spread in outcomes. Overall, the segment-level results highlight the primacy of longitudinal dynamics in shaping corridor risk, with acceleration, braking, and headway dominating near-miss severity, while roadway design elements such as medians, lane width, and access density govern both severity and variability. The superior fit of the HBSGRP model underscores the importance of capturing these segment-specific dynamics, as fixed-parameter structures fail to account for the spatially contingent influence of speed and spacing across midblock contexts.
The HBSGRP specification for segment-level V–I near-misses provides a detailed view of how longitudinal dynamics and roadway design influence the severity and variability of extreme 2D-TTC outcomes (Table 8, Fig. 5.2). In this model, the intercept and relative distance are treated as group-level random effects in both the location and scale components of the HBSGRP model, allowing for spatial heterogeneity in how spacing between interacting vehicles shapes COR across segments. All other predictors—including relative speed, acceleration, deceleration, jerk, heading difference, steering angle difference, traffic volume, lane attributes, driveway density, median type, and left-lane presence, are modeled as fixed effects, reflecting their expected stability across segment environments.
(a) $\beta_{\mu,0[1,18]}$
(b) $\gamma_{\mu,1[1,18]}$
(c) $\beta_{\sigma,0[1,18]}$
(d) $\gamma_{\sigma,1[1,18]}$
(e) $\beta_{\xi,0[1,18]}$
Fig. 12.: Posterior estimates V-I near misses at segments
Table 8: Posterior estimates V-I near misses at segments
Model Parameters Hyperparameter Covariate HBSFP HBSGRP
Mean SDa 95% CRIb Mean SDa 95% CRIb
Location Parameter ($\mu_{k,i}$)
$\beta_{\mu,0}$ Fixed Intercept -1.287 0.037 $1.352,-1.22]^{\dagger} - - -$
$\beta_{\mu,0,k}$ Random Intercept - - - -1.263 0.122 $1.494,-1.037]^{\dagger}$
$\gamma_{\mu,1,k}$ Random Relative distance - - - -0.317 0.148 $0.632,-0.104]^{\dagger}$
$\beta_{\mu,1}$ Fixed Relative speed -0.027 0.018 $061,-0.008]^{\dagger} 0.0$09 0.001 $0.006,0.027]^{\dagger}$
$\beta_{\mu,2}$ Fixed Relative acceleration 8.36E-04 0.019 [-0.045, 0.036] 0.025 0.013 $0.002,0.057]^{\dagger}$
$\beta_{\mu,3}$ Fixed Relative deceleration -0.039 0.008 $098,-0.004]^{\dagger} -0.0$32 0.008 $0.087,-0.014]^{\dagger}$
$\beta_{\mu,4}$ Fixed Relative distance -0.257 0.025 $0.304,-0.207]^{\dagger} - - -$
$\beta_{\mu,5}$ Fixed Jerk -0.124 0.069 [-0.224, 0.289 ] -0.074 0.039 [-0.124, 0.202]
$\beta_{\mu,6}$ Fixed Heading difference 0.025 0.025 [-0.025, 0.069] 0.027 0.014 $0.001,0.049]^{\dagger}$
$\beta_{\mu,7}$ Fixed Steering difference 0.033 0.017 $004,0.065]^{\dagger} 0.0$41 0.016 $0.003,0.066]^{\dagger}$
$\beta_{\mu,8}$ Fixed Volume -0.002 0.007 [-0.016,0.010] 0.043 0.021 $0.015,0.084]^{\dagger}$
$\beta_{\mu,9}$ Fixed Lane no -0.001 0.012 [-0.019, 0.016] 0.042 0.021 $0.004,0.078]^{\dagger}$
$\beta_{\mu,10}$ Fixed Lane width -0.027 0.018 $061,-0.008]^{\dagger} 0.0$13 0.009 $0.000,0.032]^{\dagger}$
$\beta_{\mu,11}$ Fixed Driveway density -0.007 0.024 [-0.057, 0.042] -0.237 0.105 $0.403,-0.020]^{\dagger}$
$\beta_{\mu,12}$ Fixed Median (Undivided=1, else=0) -0.644 0.102 $773,-0.513]^{\dagger} -0.0$83 0.059 $0.196,-0.019]^{\dagger}$
$\beta_{\mu,13}$ Fixed Vehicle position (Left lane=1, else=0) 0.057 0.033 $028,0.091]^{\dagger} -0.0$64 0.014 $0.082,-0.038]^{\dagger}$
Scale Parameter ($\log\sigma_{k,i}$)
$\beta_{\mu,0}$ Fixed Intercept -0.495 0.156 $0.703,-0.303]^{\dagger} - - -$
$\beta_{\sigma,0,k}$ Random Intercept - - - -0.076 0.104 [-0.265, 0.083]
$\gamma_{\sigma,1,k}$ Random Relative distance - - - 0.135 0.103 $0.011,0.318]^{\dagger}$
$\beta_{\sigma,1}$ Fixed Relative speed 0.039 0.016 $012,0.061]^{\dagger} 0.0$13 0.001 $0.004,0.038]^{\dagger}$
$\beta_{\sigma,2}$ Fixed Relative acceleration 0.002 0.017 [-0.030, 0.042] -0.026 0.014 $0.064,-0.001]^{\dagger}$
$\beta_{\sigma,3}$ Fixed Relative deceleration -0.036 0.013 $075,-0.014]^{\dagger} -0.0$48 0.021 $0.089,-0.022]^{\dagger}$
$\beta_{\sigma,4}$ Fixed Relative distance 0.167 0.018 $0.136,0.199]^{\dagger} - - -$
$\beta_{\sigma,5}$ Fixed Jerk -0.108 0.054 [0.452, -0.119] -0.078 0.119 [-0.312, 0.078]
$\beta_{\sigma,6}$ Fixed Heading difference -0.019 0.020 [-0.053, 0.020] -0.029 0.017 $0.049,-0.001]^{\dagger}$
$\beta_{\sigma,7}$ Fixed Realative steering angle -0.033 0.012 $065,-0.001]^{\dagger} -0.0$32 0.024 $0.059,-0.003]^{\dagger}$
$\beta_{\sigma,8}$ Fixed Volume -0.007 0.018 [-0.039, 0.027] 0.001 0.016 [-0.026, 0.030]
$\beta_{\sigma,9}$ Fixed Lane no -0.028 0.017 $058,-0.006]^{\dagger} -0.0$21 0.012 $0.046,-0.002]^{\dagger}$
$\beta_{\sigma,10}$ Fixed Lane width -0.010 0.020 [-0.044, 0.030] 0.018 0.016 $0.010,0.045]^{\dagger}$
$\beta_{\sigma,11}$ Fixed Driveway density 0.016 0.019 [-0.021, 0.055] -0.125 0.028 [-0.239, 0.018]
$\beta_{\sigma,12}$ Fixed Median (Undivided=1, else=0) 0.663 0.166 $0.440,0.880]^{\dagger}$ 0.174 0.090 $0.263,0.553]^{\dagger}$
$\beta_{\sigma,13}$ Fixed Vehicle position (Left lane=1, else=0) -0.046 0.027 $080,-0.012]^{\dagger} 0.0$55 0.010 $0.039,0.071]^{\dagger}$
Shape Parameter ($\xi_{k,i}$)
$\beta_{\xi,0}$ Fixed Intercept -0.845 0.028 $0.897,-0.788]^{\dagger}$g - - -
$\beta_{\xi,0,k}$ Random Intercept - - - -0.887 0.118 $1.123,-0.659]^{\dagger}$
Model Fit
DIC 6760 6550
WAIC 6775 6595
LOOIC 6791.2 6601
  • HBSFP
    Hierarchical Bayesian Spatial Fixed Parameter
  • HBSGRP
    Hierarchical Bayesian Spatial Grouped Random Parameter
  • a
    Standard deviation
  • b
    95% Bayesian credible interval
    Covariate not included in the model
  • $\dagger$}
    Indicates statistical significance at the 95% level (interval excludes 0)
For the location parameter, shorter relative distances strongly reduce 2D-TTC values, confirming that closer spacing consistently produces more critical near-miss outcomes. The random effect for distance further indicates that influence varies by segment group, suggesting sensitivity to local geometric or operational conditions. Relative speed displays a notable divergence: while the fixed-parameter benchmark shows a negative coefficient, the grouped model produces a small but positive estimate, implying that in certain contexts, higher speeds are associated with marginally longer 2D-TTC values. This reversal highlights that segment-level heterogeneity in speed environments, such as free-flow arterials versus constrained multilane facilities, can fundamentally alter the role of speed in shaping severity. Relative acceleration emerges as a significant positive effect in the grouped specification, suggesting that acceleration phases along segments may often represent clearance or recovery maneuvers rather than heightened near-miss risk. In contrast, relative deceleration maintains a negative coefficient across both models, reflecting the reality that braking during close-following conditions is tied to critical reductions in 2D-TTC, likely triggered by sudden evasive actions.
Several geometric and operational features are also important for location. Heading and steering differences exhibit significant positive coefficients, indicating that lateral adjustments yield small gains in clearance and mitigate near-miss severity. Volume, lane number, and lane width also emerge as positive contributors in the HBSGRP specification, suggesting that higher-capacity, wider facilities afford drivers greater maneuvering space, leading to slightly larger 2D-TTC values. The effect of driveway density, however, is particularly striking: while insignificant in the fixed-parameter model, the grouped specification produces a large negative estimate with a wide CI. This instability suggests that driveway effects are highly localized—segments with dense access points may concentrate more severe near-miss outcomes, but the data are sparse and sensitive to segment-specific characteristics. Median type is also significant, with undivided facilities consistently producing lower 2D-TTC values, reflecting heightened exposure to opposing traffic flows. The left-lane variable shows a sign reversal, with the fixed model suggesting higher severity in interior lanes and the grouped model indicating greater severity in outer lanes. This again illustrates heterogeneity across segment groups, where the positional context of a lane relative to roadside boundaries and merging activity shifts the locus of risk.
The scale parameter provides complementary insights into the dispersion of near-miss severity. The random effect for relative distance is positive, showing that tighter spacing increases not only severity but also variability in outcomes, as different drivers adopt divergent strategies under minimal headway conditions. Relative speed also has a positive association with scale, indicating that higher speeds amplify heterogeneity in outcomes across V–I interactions, unlike in V–V segment near-misses, where speed reduces dispersion. Acceleration, by contrast, has a negative coefficient in the grouped model, suggesting that acceleration phases yield more predictable outcomes, consistent with smoother-flow recovery dynamics. Relative deceleration also reduces variability, implying that braking in near-crash contexts converges toward uniformly critical outcomes.
Geometric attributes reveal nuanced effects on variability. Lane width exerts a positive effect, suggesting that while wider lanes modestly reduce average severity, they also introduce greater variability in outcomes by enabling diverse driver responses. The undivided median type strongly increases scale, underscoring that undivided facilities not only heighten the average risk but also broaden the range of severity outcomes, likely due to variation in driver avoidance maneuvers. The left-lane effect is positive in scale, indicating that severity outcomes in these lanes are more heterogeneous, perhaps reflecting the diversity of driver strategies in navigating outer versus inner lanes. Driveway density, although statistically unstable, shows a wide and imprecise effect that suggests variability is strongly contingent on site-specific access management patterns.
The shape parameter remains negative and significant across both specifications, confirming that the upper tail of the 2D-TTC distribution is bounded. This supports the Weibull characterization of extremes, implying that even in segment-level V–I interactions, physical and behavioral limits constrain how severe the most critical near-miss events can be. Variation in shape across segment groups highlights the role of local geometry, access density, and operational features in shaping the extremal properties of near-miss distributions.
Together, the segment-level analyses of V–V and V–I near misses underscore that roadway geometry and vehicle dynamics interact to shape both the severity and variability of extreme near-misses. For V–V events, relative speed and distance dominate as key random effects, with speed consistently amplifying severity and distance compressing variability, reflecting deterministic outcomes under close-following conditions. For V–I events, however, the role of speed is more heterogeneous—sometimes protective, sometimes risk-enhancing—depending on the segment context. Access-related variables, particularly driveway density and median type, emerge as critical for V–I outcomes, underscoring the importance of access management and median design in shaping roadside safety. In both near-miss types, the negative and significant shape parameter indicates bounded extremes, reinforcing the physical and behavioral limits to crash risk escalation.
6 Model validation
The dataset for the study corridor provides precise geolocation for each extreme near-miss event and readily available annual corridor-level observed crashes. The dataset for the study corridor provides precise geolocation for each extreme near-miss event from Eqn. 32. To address this limitation, validation was conducted as an event-level case–control approach, designed to assess whether the model assigns higher predicted probabilities to more severe near-miss events. Posterior exceedance probabilities were computed from Eqn. 30 for severity thresholds ranging from –0.1 to –0.9 seconds, and these probabilities were then used to evaluate classification performance using the area under the receiver operating characteristic curve (ROC–AUC). This metric is particularly well-suited to rare-event prediction because it is independent of a specific threshold.
(a) V–V Segment
(b) V–I Segment
(c) V–V Intersection
(d) V–I Intersection
Fig. 13.: ROC curves for four HBSGRP–UGEV models across severity thresholds from $\omega=-0.1$s to $\omega=-0.9$s
ROC curves were generated for the four HBSGRP–UGEV variants, covering both V-V, V–I interactions at segment and intersection levels. The results (Figure 6) reveal clear distinctions across near-miss types and roadway contexts. The segment-level V-V model consistently achieved the highest accuracy, with mean AUC values of about 0.89 and very little variation across thresholds. This indicates stable and robust predictive performance, reflecting the ability of grouped random parameters to capture high-speed longitudinal near-miss dynamics along uninterrupted midblock segments. The V-I model at the segment level achieved slightly lower accuracy, averaging around 0.79, and exhibited moderate sensitivity to the choice of severity threshold, a finding consistent with the greater variability in driver responses when vehicles approach roadside boundaries. At intersections, the V-V model demonstrated moderately strong predictive accuracy, with an average AUC of nearly 0.82 and relatively low volatility across thresholds. This indicates that grouped speed and spacing dynamics remain reliable even under more complex signalized and turning conditions. By contrast, the intersection-level V-I model performed weakest, averaging an AUC of approximately 0.75 and showing large fluctuations across thresholds, which reflects the complexity of boundary-related near-misses in intersections where turning trajectories, curb proximity, traffic signals, and pedestrian activity all interact.
Taken together, these findings confirm that the HBSGRP–UGEV framework provides strong and reliable performance for V-V interactions, particularly in segment environments, while predictions for V-I interactions are more sensitive to threshold selection and local variability. This suggests that segment-level V-V estimates are the most operationally reliable for proactive safety monitoring, whereas intersection-level V-I near-misses may require further refinement. Incorporating more detailed information on curb geometry, pedestrian exposure, and high-resolution vehicle–infrastructure proximity measures could improve stability and predictive accuracy in future applications.
7 Summary and conclusions
This study addresses the persistent challenge of accurately estimating short-term crash occurrence risk (COR) in complex, mixed urban traffic environments, where conventional surrogate safety models often overlook detailed vehicle geometry, roadway context, and site-specific behavioral variability. Existing approaches are limited by oversimplified near-miss representations, a predominant focus on vehicle–vehicle (V–V) interactions, and insufficient integration of spatial heterogeneity in statistical modeling. To overcome these gaps, we propose a high-fidelity, geometry-aware, and statistically robust framework for real-time COR estimation that introduces a novel two-dimensional time-to-collision (2D-TTC) indicator. This metric accounts for vehicle dimensions, dynamic states, and precise geometric positioning, enabling accurate detection of extreme near-miss events from both V–V and vehicle–infrastructure (V–I) interactions.
Extreme events are translated into probabilistic risk estimates through a non-stationary Univariate Generalized Extreme Value (UGEV) framework implemented within a Hierarchical Bayesian Spatial Grouped Random Parameter (HBSGRP) structure. A corridor-based grouping strategy was employed to address event sparsity, pooling near-miss samples across intersections and directional segments while preserving local heterogeneity. Execution of the framework involved extracting high-resolution vehicle trajectories from the Argoverse-2 motion forecasting dataset, integrating them with HD maps and supplementary roadway information, and identifying near-miss events. The model incorporated both vehicle dynamics covariate (relative speed, acceleration, deceleration, jerk, heading difference, steering difference, and traffic volume) and geometric or operational covariates (lane number, lane width, driveway density, median type, turning movement, and vehicle lane position). Grouped random parameters were specified for temporally varying dynamics, relative speed, relative deceleration, and relative distance, to capture their context-dependent effects. For benchmarking, a baseline hierarchical Bayesian fixed-parameter (HBSFP) model was also estimated.
Results demonstrate that HBSGRP consistently outperformed the HBSFP across both intersections and segments, achieving lower DIC, WAIC, and LOOIC values and providing greater estimation stability. Relative speed emerged as the dominant predictor of near-miss severity in intersections, while relative distance was more influential for segment-level V–V events. Deceleration played a critical role in segment-level V–V near-miss, underscoring the importance of longitudinal control. Fixed covariates such as heading difference, steering difference, lane width, driveway density, and median type also showed significant effects, demonstrating the ability of the model to capture both dynamic behaviors and roadway design influences.
Validation results further confirmed model credibility. ROC–AUC analysis showed strong predictive performance, with V–V segment models achieving the highest accuracy (AUC $\approx$ 0.89), followed by intersection V–V ($\approx$ 0.82). V–I models performed reasonably well (segment $\approx$ 0.79; intersection $\approx$ 0.75) but displayed greater sensitivity to severity threshold choice. These findings highlight that while V–V interactions are modeled with high stability, V–I near-misses especially at intersections, remain more complex and variable.
Despite its contributions, the study has several limitations. First, potential bias may arise from the dependence on AV onboard sensors, which may affect traffic state accuracy and near-miss detection. Second, the 2D-TTC formulation, while tractable, simplifies nonlinear vehicle dynamics, which could understate risks in extreme evasive maneuvers. Third, the current framework does not explicitly model near-miss severity levels. Future research should extend the framework to multivariate UGEV models incorporating complementary surrogate measures such as post-collision velocity change ($\Delta V$) and explicitly address vulnerable road user (VRU) near-misses. Broader validation across multiple cities, transferability calibration, and comparison with peaks-over-threshold (POT) GPD methods will further strengthen robustness and generalizability. As autonomous vehicle deployment increases, access to larger and richer trajectory datasets will enhance the statistical power of extreme value analysis and enable more nuanced safety assessments. In practice, the proposed HBSGRP–UGEV framework offers a scalable, data-driven tool for proactive safety monitoring at the corridor level. By enabling early identification of high-risk conditions and supporting real-time traffic management, it provides a pathway toward Vision Zero objectives, reducing fatalities and serious injuries while promoting safe, efficient, and equitable mobility in urban environments.
Acknowledgement:
This research is funded by Federal Highway Administration (FHWA) Exploratory Advanced Research 693JJ323C000010. The results do not reflect FHWA’s opinions.
Disclaimer:
The results presented in this document do not necessarily reflect those from the Federal Highway Administration.
Data availability:
Data and code will be made available on reasonable request.
[…]

### v2

@@
 Real-time prediction of corridor-level crash occurrence risk (COR) remains challenging, as existing near-miss based extreme value models oversimplify collision geometry, exclude vehicle–infrastructure (V–I) interactions, and inadequately capture spatial heterogeneity in vehicle dynamics.
-This study introduces a geometry-aware two-dimensional time-to-collision (2D-TTC) indicator within a Hierarchical Bayesian spatiotemporal grouped random parameter (HBSGRP) framework using a non-stationary univariate generalized extreme value (UGEV) model to estimate short-term COR in urban corridors.
-High-resolution trajectories from the Argoverse-2 dataset, covering 28 locations along Miami’s Biscayne Boulevard, were analyzed to extract extreme V–V and V–I near misses.
-The model incorporates dynamic variables and roadway features as covariates, with partial pooling across locations to address unobserved heterogeneity.
-Results show that the HBSGRP–UGEV framework outperforms fixed-parameter alternatives, reducing DIC by up to 7.5% for V–V and 3.1% for V–I near-misses.
+This study introduces a geometry-aware two-dimensional time-to-collision (2D-TTC) indicator within a Hierarchical Bayesian spatiotemporal grouped random parameter (HBSGRP) framework using a nonstationary univariate generalized extreme value (UGEV) model to estimate short-term COR in urban corridors.
+High-resolution trajectories from the Argoverse-2 dataset, covering 28 locations along Miami’s Biscayne Boulevard, were analyzed to extract extreme near misses (V–V and V–I).
+The model incorporates dynamic variables and road features as covariates, with partial pooling across locations to address unobserved heterogeneity.
+The results show that the HBSGRP–UGEV framework outperforms fixed-parameter alternatives, reducing the DIC by up to 7.5% for V–V and 3.1% for V–I near-misses.
 Predictive validation using ROC–AUC confirms strong performance: 0.89 for V–V segments, 0.82 for V–V intersections, 0.79 for V–I segments, and 0.75 for V–I intersections.
@@
 1 Introduction
-The increasing complexity of urban road networks, rising traffic volumes, and the coexistence of human-driven vehicles (HDVs) and autonomous vehicles (AVs) demand innovative safety assessment approaches.
+The increasing complexity of urban road networks, rising traffic volumes, and the coexistence of human-driven vehicles (HDVs) and autonomous vehicles (AVs) require innovative safety assessment approaches.
 Urban corridors are particularly vulnerable, where a single incident can disrupt mobility, trigger cascading congestion, and impose severe economic losses.
-In 2023, traffic crashes in the United States claimed 40,901 lives, with a fatality rate of 1.26 per 100 million vehicle miles traveled (VMT) (National Center for Statistics and Analysis, 2025).
+In 2023, traffic crashes in the United States claimed 40901 lives, with a fatality rate of 1.26 per 100 million vehicle miles traveled (VMT) (National Center for Statistics and Analysis, 2025).
 The total annual cost of crashes is estimated to exceed \$1.85 trillion, including \$460 billion in direct economic losses and \$1.4 trillion in reduced quality of life (Blincoe et al., 2023).
@@
 Most existing models focus exclusively on vehicle–vehicle (V–V) near-misses, largely due to the availability of vehicle trajectory data, while systematically excluding vehicle–infrastructure (V–I) interactions such as lane edges, curbs, barriers, and crosswalks.
-Yet V–I events account for a substantial proportion of severe run-off-road and roadway departure crashes (Islam et al., 2021).
-This omission leaves corridor-level safety models inherently incomplete.
-Recent advances in AV sensing and high-definition (HD) mapping, such as the Argoverse-2 dataset (Wilson et al., 2023), now enable the precise integration of roadway boundaries with vehicle trajectories, allowing both V–V and V–I near-misses to be modeled with geometric fidelity.
+However, V–I events account for a substantial proportion of severe run-off-road and roadway departure crashes (Islam et al., 2021).
+This omission makes corridor-level safety models inherently incomplete.
+Recent advances in AV sensing and high-definition (HD) mapping, such as the Argoverse-2 dataset (Wilson et al., 2023), now allow the precise integration of roadway boundaries with vehicle trajectories, allowing both V–V and V–I near-misses to be modeled with geometric fidelity.
 Crash risk prediction strategies based on historical retrospective crash data (Arun et al., 2021; Lee et al., 2017; Pei et al., 2011) suffer from well-documented limitations (Lord and Mannering, 2010; Lord et al., 2021; Mannering et al., 2020), including the rarity and stochastic nature of crashes, underreporting (Arun et al., 2021; Tarko, 2018), temporal delays, and spatial mismatches with real-time conditions.
-These challenges motivate a paradigm shift toward proactive, trajectory-based surrogate safety measures (SSMs) (Ali et al., 2023a; Arun et al., 2021; Mahmud et al., 2017).
-Derived from high-resolution vehicle trajectory data, SSMs reflect underlying near-miss dynamics (Li et al., 2024; Tarko, 2018; Davis et al., 2011).
+These challenges motivate a paradigm shift towards proactive trajectory-based surrogate safety measures (SSMs) (Ali et al., 2023a; Arun et al., 2021; Mahmud et al., 2017).
+Using high-resolution vehicle trajectory data, SSMs reflect the underlying near-miss dynamics (Li et al., 2024; Tarko, 2018; Davis et al., 2011).
 Among SSMs, time-based indicators such as Time-to-Collision (TTC), Post-Encroachment Time (PET), and Modified TTC (MTTC) are widely used (Hayward, 1971; Perkins and Harris, (1967; Minderhoud and Bovy, 2001; Ozbay et al., 2008; Allen et al., 1978; Venthuruthiyil and Chunchu, 2022).
-However, these formulations often assume constant velocity and linear trajectories, which bias results in boundary-constrained environments.
-Recent research has extended TTC into higher-fidelity two-dimensional (2D) indicators (Li et al., 2024; Anis et al., 2025b), incorporating latitude and longitude position, heading, and acceleration.
-Yet, even these advances typically employ simplified circular approximations, which can influence risk estimates in urban contexts where vehicle shape and alignment significantly influence near-miss likelihood and severity.
-To support reliable short-term crash occurrence risk (COR) estimation, SSM frameworks must enhance geometric fidelity to more accurately represent real-world interactions between vehicles and infrastructure.
+However, these formulations often assume constant velocity and linear trajectories, leading to biased results in boundary-constrained environments.
+Recent research has extended TTC to two-dimensional (2D) indicators of higher-fidelity (Li et al., 2024; Anis et al., 2025b), incorporating latitude and longitude position, heading, and acceleration.
+However, even these advances typically employ simplified circular approximations, which can influence risk estimates in urban contexts where vehicle shape and alignment significantly influence the likelihood and severity of near misses.
+To support a reliable short-term crash occurrence risk (COR) estimate, SSM frameworks must enhance geometric fidelity to more accurately represent real-world interactions between vehicles and infrastructure.
 This study builds on these efforts by introducing a high-fidelity, geometry-aware 2D-TTC indicator that explicitly incorporates vehicle geometry, dynamic states, and roadway boundaries, ensuring realistic and efficient detection of both V–V and V–I near misses.
@@
 The BM approach is particularly suitable for real-time analysis due to its short block structure, lower threshold sensitivity, and compatibility with discrete-time samples (Fu and Sayed, 2022a).
-Early EVT applications in traffic safety employed stationary GEV models (Wang et al., 2018; Guo et al., 2019; Orsini et al., 2019; Alozi and Hussein, 2022; Hussain et al., 2022) that assumed constant parameterization across the study area.
+Early applications of EVT in traffic safety employed stationary GEV models (Wang et al., 2018; Guo et al., 2019; Orsini et al., 2019; Alozi and Hussein, 2022; Hussain et al., 2022) that assumed constant parameterization across the study area.
 More recently, non-stationary GEV models have been introduced, allowing covariates to influence location and scale parameters (Tahir and Haque, 2024; Ali et al., 2022; Fu and Sayed, 2022a; Zheng et al., 2019; Nazir et al., 2023; Kar et al., 2024; Ali et al., 2023b; Alozi and Hussein, 2022; Zheng et al., 2014; Kar et al., 2023) and thereby capture heterogeneity across traffic states (Lanzaro et al., 2023; Kamel et al., 2023, 2024; Kamel and Sayed, 2024; Singh et al., 2024; Ghoul and Sayed, 2025; Singh et al., 2025; Anis et al., 2025b).
-Bayesian hierarchical extensions have further addressed data sparsity and unobserved heterogeneity (Zheng and Sayed, 2019; Fu and Sayed, 2022b).
+Bayesian hierarchical extensions have also addressed data sparsity and unobserved heterogeneity (Zheng and Sayed, 2019; Fu and Sayed, 2022b).
 However, most applications remain geographically narrow and functionally limited to either intersections (Fu and Sayed, 2023, 2021c, 2022b) or segments (Kumar and Mudgal, 2024; Kamel et al., 2023; Kamel and Sayed, 2024; Anis et al., 2025b; Kamel et al., 2024).
@@
 Therefore, a critical gap remains for a comprehensive Bayesian hierarchical GEV framework capable of jointly modeling vehicle dynamics, roadway geometry, flow directionality, and multilevel heterogeneity.
-Such a framework is essential for accurate corridor-level COR estimation that reflects the complex spatiotemporal dynamics of urban traffic environments.
+This framework is crucial for accurately estimating the COR at the corridor level, which reflects the complex spatiotemporal dynamics of urban traffic environments.
 This paper makes the following unique contributions:
-  • Develops a corridor-level COR framework that simultaneously incorporates both V–V and V–I interactions, addressing a critical omission in prior studies.
-  • Introduces a high-fidelity, geometry-aware 2D-TTC indicator that realistically captures extreme near-miss events by accounting for vehicle geometry, dynamic states, and roadway boundaries.
+  • Develops a corridor-level COR framework that simultaneously incorporates both V–V and V–I interactions, addressing a critical omission in previous studies.
+  • Introduces a geometry-aware high-fidelity 2D-TTC indicator that realistically captures extreme near-miss events by accounting for vehicle geometry, dynamic states, and roadway boundaries.
   • Advances a hierarchical Bayesian spatiotemporal grouped random parameter (HBSGRP) Univariate GEV model that integrates roadway geometry, traffic flow directionality, and multilevel heterogeneity, enabling corridor-wide yet location-sensitive crash risk estimation
@@
 The framework integrates a continuous-time kinematic bicycle model, vehicle geometry representation, and a spatial near-miss detection algorithm, allowing capture of V-V and V-I interactions.
-These components collectively enable the accurate forecasting of vehicle trajectories and the identification of potential near misses.
+Together, these components enable the accurate forecasting of vehicle trajectories and identification of potential near misses.
 The resulting architecture forms the computational backbone of the SSM analysis, ensuring a realistic representation of near-miss events in mixed-traffic environments.
@@
 2.2 Vehicle geometry representation
-While the dynamic model predicts vehicle positions and headings, accurate near-miss detection requires a realistic representation of the vehicle footprint.
+Although the dynamic model predicts vehicle positions and headings, accurate near-miss detection requires a realistic representation of the vehicle footprint.
 Simplifying vehicles as points or circles often misrepresents interactions, especially during lane changes, swerving, or angled approaches.
@@
 At each timestep of the RK4 integration, the global positions of all four corners are computed using the kinematic and rotational transformations.
-A collision is defined to occur when a corner of one vehicle lies within a proximity threshold of a corner of another.
+A collision is defined when a vehicle’s corner lies within a proximity threshold of a corner of another vehicle.
 Let ${p}_{A}^{(j)}(t)$ and ${p}_{B}^{(k)}(t)$ denote the global coordinates of corners $j$ and $k$ of vehicles A and B at time $t$.
@@
 2.4 V-I near-miss detection
-In addition to V-V interactions, the framework also detects near-misses between vehicles and roadway infrastructure such as curbs, medians, or barriers.
+In addition to V-V interactions, the framework also detects near misses between vehicles and roadway infrastructure, including curbs, medians, or barriers.
 Capturing these V–I events is critical for identifying edge encroachments or departures that often precede run-off-road crashes.
@@
 Fig. 1.: Computation of 2D-TTC for near-miss scenarios
-While this geometric framework enables precise localization of near-misses, it does not capture the probabilistic structure or temporal evolution of COR.
+While this geometric framework enables precise localization of near misses, it does not capture the probabilistic structure or temporal evolution of COR.
 To address this, the extracted blockwise minima of 2D-TTC are modeled using the EVT model.
@@
 Two EVT approaches are commonly applied: the Block Maxima (BM) method, which fits maxima or minima from fixed-duration blocks using the Generalized Extreme Value (GEV) distribution (Coles et al., 2001), and the Peaks Over Threshold (POT) method, which models threshold exceedances using the Generalized Pareto Distribution (GPD).
-While POT can be data-efficient, it is highly sensitive to threshold choice and less compatible with discrete, sensor-derived datasets.
+Although POT can be data-efficient, it is highly sensitive to threshold choice and less compatible with discrete sensor-derived datasets.
 By contrast, BM aligns naturally with real-time AV data, avoids subjective thresholds, and supports scalable, time-series analysis (Fu and Sayed, 2022a; Songchitruksa and Tarko, 2006).
-Accordingly, this study adopts the BM–GEV framework (see Fig.
+Consequently, this study adopts the BM–GEV framework (see Fig.
 2.5).
@@
 Recent advances incorporate nonstationary structures, linking GEV parameters to covariates such as vehicle dynamics (Anis et al., 2025b; Fu and Sayed, 2021c, 2022b).
-Bayesian hierarchical models have further addressed sparse data and unobserved heterogeneity across vehicles, sites, and traffic contexts.
+Bayesian hierarchical models have also addressed sparse data and unobserved heterogeneity across vehicles, sites, and traffic contexts.
 The increasing availability of high-resolution AV trajectory and HD map data has highlighted the need for flexible GEV formulations that jointly capture V–V and V–I interactions.
@@
 Roadway geometry (e.g., lane count, curvature, median type) is spatially structured and relatively stable within segments or intersections, while vehicle dynamics (e.g., speed, acceleration) vary temporally and spatially, driving short-term behavioral volatility.
-Capturing this multilevel heterogeneity is crucial for estimating corridor- and network-level crash risk.
+Capturing this multilevel heterogeneity is crucial for estimating corridor and network level crash risk.
 To address this gap, this study develops a hierarchical Bayesian Spatiotemporal Grouped Random Parameter (HBSGRP) UGEV model.
@@
 3.1 Univariate generalized extreme value (UGEV)
-As established in the previous section, the BM-based UGEV model is employed to characterize the tail behavior of 2D-TTC, which represents the most safety-critical near-miss events in both V–V and V–I interactions.
+As established in the previous section, the BM-based UGEV model is employed to characterize the tail behavior of 2D-TTC, which represents the most critical near-miss events in both V–V and V–I interactions.
 To implement this framework, continuous vehicle trajectories are partitioned into fixed-duration windows (e.g., 11 s).
@@
 This structure allows systematic differences across predefined observational groups while maintaining parsimony and interpretability.
-In this specification, covariates are partitioned into two sets: fixed effects $X$, which apply uniformly across all observations, and group-varying effects $Z$, whose coefficients differ across groups.
-For a dataset divided into $K$ groups, with observation $i$ belonging to group $k(i)$, both the intercept and selected slopes are allowed to vary by group.
+In this specification, covariates are divided into two sets: fixed effects $X$, which apply uniformly across all observations, and group-varying effects $Z$, whose coefficients differ across groups.
+For a dataset divided into $K$ groups, with observations $i$ belonging to group $k(i)$, both the intercept and selected slopes are allowed to vary by group.
 This enables the capture of localized relationships while preserving overall model coherence.
@@
 The data layer specifies the likelihood of observed block maxima of 2D-TTC from V–V and V–I interactions, assuming a UGEV distribution.
-Process layer links covariates to UGEV parameters through fixed and group-level random effects, with optional latent terms (e.g., Gaussian processes) to capture residual spatial–temporal dependence.
-Prior layer defines prior distributions for all parameters and hyperparameters, supporting regularization and uncertainty quantification.
+The process layer links covariates to UGEV parameters through fixed and group-level random effects, with optional latent terms (e.g., Gaussian processes) to capture residual spatial–temporal dependence.
+The prior layer defines prior distributions for all parameters and hyperparameters, supporting regularization and uncertainty quantification.
 Bayesian inference proceeds by updating prior beliefs with observed data to obtain the posterior distribution. Formally, the joint posterior of parameters $\Theta$, given data $X$, is:
@@
 $\xi_{k,i}=\beta_{\xi,0}+\sum_{j=1}^{J}\beta_{\xi,j}X_{k,j}$ (26)
-In the above equations $X_{k,j}$ defined as group-level fixed-effect covariate $j$, $W_{k,i,j}$is sample-specific random-effect covariate $j$.
+In the above equations, $X_{k,j}$ defined as a group-level fixed-effect covariate $j$, $W_{k,i,j}$is a sample-specific random-effect covariate $j$.
 $\beta_{\mu,0},\beta_{\sigma,0},\beta_{\xi,0}$ as fixed intercepts for respective parameters.
@@
 3.2 COR estimation
-Building on the HBSGRP–UGEV estimates, the framework compute COR by evaluating the probability that 2D-TTC for both V–V and V–I interactions falls below a critical threshold within each block.
-Unlike severity-stratified approaches, this framework estimates the total number of latent extreme near-miss events regardless of outcome severity, enabling a holistic assessment of potential COR under mixed traffic conditions.
-Let $i$ denote block $i$ within group $k$, where groups correspond to spatial or functional corridor classifications (e.g., intersections or directional segments).
+Building on the HBSGRP–UGEV estimates, the framework computes COR by evaluating the probability that 2D-TTC for both V–V and V–I interactions falls below a critical threshold within each block.
+Unlike severity-stratified approaches, this framework estimates the total number of latent extreme near-miss events regardless of outcome severity, allowing for a holistic assessment of potential COR under mixed traffic conditions.
+Let $k,i$ denote the block $i$ within the group $k$, where groups correspond to spatial or functional corridor classifications (e.g., intersections or directional segments).
 The minimum 2D-TTC in each block is modeled with the HBSGRP-UGEV distribution, parameterized by $\mu_{k,i}$, $\sigma_{k,i}$, and $\xi_{k,i}$.
@@
 $P_{k,i}^{\text{crash}}=\Pr(2D-TTC_{k,i}\leq\omega)=1-\exp\left\{-\left[1-\xi_{k,i}\left(\frac{\omega-\mu_{k,i}}{\sigma_{k,i}}\right)\right]^{-1/\xi_{k,i}}\right\},\quad\text{where }1-\xi_{k,i}\left(\frac{\omega-\mu_{k,i}}{\sigma_{k,i}}\right)>0$ (30)
-Assuming a homogeneous Poisson process for exceedances, the expected COR for group $k$ is obtained by aggregating exceedance probabilities across blocks and normalizing by exposure duration $T$.
+Assuming a homogeneous Poisson process for exceedances, the expected COR for the group $k$ is obtained by aggregating exceedance probabilities across blocks and normalizing by exposure duration $T$.
 Here $y_{k,i}$ is the number of near-misses (V-V or V-I) in block $i$, and $N_{k}$ the number of blocks in group $k$.
@@
 4 Data description
-The effectiveness of real-time COR estimation depends critically on the fidelity and coverage of trajectory data.
+The effectiveness of real-time COR estimation is critically dependent on the fidelity and coverage of trajectory data.
 Traditional sensing technologies, such as roadside cameras, loop detectors, and fixed LiDAR, have supported safety analysis by capturing speed, spacing, and acceleration patterns (Islam et al., 2021; Yuan et al., 2019; Li et al., 2020).
@@
 Recent advances in AV sensing offer a promising alternative.
-AV datasets offer high-frequency, multimodal trajectories with unprecedented spatial precision, allowing for detailed analysis of kinematics, spacing, and maneuvers.
+AV datasets provide high-frequency, multimodal trajectories with unprecedented spatial precision, allowing for detailed analysis of kinematics, spacing, and maneuvers.
 This richness allows for realistic near-miss detection in complex traffic environments.
 Despite this potential, AV data remain underutilized in generalized frameworks for real-time COR estimation.
-For example, studies using the Waymo Open Dataset have largely focused on localized near-miss events without incorporating roadway geometry or V-I interactions, limiting broader applicability (Anis et al., 2025b).
-To address these challenges, this study utilizes the Argoverse-2 dataset (Wilson et al., 2023), which was collected in 2021 across six U.S.
-cities (Austin, Detroit, Miami, Palo Alto, Pittsburgh, and Washington, D.C.) using SAE Level 4 AVs.
-Argoverse-2 comprises more than 250k driving scenarios spanning 113 roadway segments, with each 11-second scenario sampled at 10 Hz.
-Compared with Waymo Open (Ettinger et al., 2021), Lyft Level-5 (Houston et al., 2021), and nuScenes (Caesar et al., 2020), Argoverse-2 is uniquely curated to emphasize safety-critical interactions such as intersection near-misses, lane changes, and pedestrian crossings—conditions most indicative of crash risk.
+For example, studies using the Waymo Open Dataset have focused mainly on localized near-miss events without incorporating roadway geometry or V-I interactions, limiting broader applicability (Anis et al., 2025b).
+To address these challenges, this study uses the Argoverse-2 dataset (Wilson et al., 2023), which was collected in 2021 across six U.S.
+cities (Austin, Detroit, Miami, Palo Alto, Pittsburgh, and D.C.) using SAE Level 4 AVs.
+Argoverse-2 comprises more than 250k driving scenarios that span 113 roadway segments, with each 11-second scenario sampled at 10 Hz.
+Compared to Waymo Open (Ettinger et al., 2021), Lyft Level-5 (Houston et al., 2021), and nuScenes (Caesar et al., 2020), Argoverse-2 is uniquely curated to emphasize safety-critical interactions, such as intersection near-misses, lane changes, and pedestrian crossings, conditions most indicative of crash risk.
 A defining feature of Argoverse-2 is its scene-adaptive sampling, which prioritizes high-interaction episodes rather than routine free-flow traffic.
@@
 HD maps contain lane centerlines, crosswalks, medians, and drivable areas registered to local coordinates.
-Table 1: Open-source AV datasets
-Dataset Sensora Tasksb Hzc HD Map Scope / Notes
-Argoverse-2 (Wilson et al. 2023) C, L D, T, S, M 10 Yes 250k scenarios across 113 segments; 11s each; 10 classes; 6 cities
-nuScenes (Caesar et al. 2020) C, L, R D, T, S, M 2 No 40k scenarios across  1k segments; 23 classes; 2 cities
-Waymo (Ettinger et al. 2021) C, L D, T, M 10 Yes 390k scenarios over  1950 segments; 12.6M trajectories; 6 cities
-Lyft Level 5 (Kesten et al. 2019) C, L, R M 10 Yes 170k AV instances (25s) in one city; 10 classes
-KITTI (Geiger et al. 2013) L D, T, M 10 No 15k scenes; classic full-stack AV benchmark
-DAIR-V2X (Yu et al. 2022) L D Yes 71254 image frames for V2X scenes
-OPV2V (Xu et al. 2022) C, L D, T, M 10 No 33k V2V simulation samples; 18k V2X frames (CARLA)
-DeepAccident (Wang et al. 2024) C, L, R D, T, S, M 10 Yes 285k simulation samples incl. 57k V2X scenarios
-Shift (Sun et al., 2022) C, L S, M 10 No 2.5M frames; 4850 segments; 8 cities
+Table 1: Representative open-source AV datasets with applications in safety and crash-risk analysis
+Dataset Trajectory coveragea Safety relevanceb Duration / Frequency HD map Notes
+Argoverse-2 250k scenarios; 113 segments; 6 cities Near-miss analysis, surrogate safety measures 11s per scenario, 10 Hz Yes Rich lane-level geometry; 10 object classes (Wilson et al. 2023)
+nuScenes 40k scenes; $\sim$1000 segments; 2 cities Multi-agent interactions, trajectory risk indicators 20s per scene, 2 Hz No 23 object classes; widely used urban dataset (Caesar et al. 2020)
+Waymo 390k scenarios; $\sim$1950 segments; 6 cities Motion forecasting, crash-prone interaction detection Up to 20s, 10 Hz Yes 12.6M trajectories across diverse geometries (Ettinger et al. 2021)
+Lyft Level 5 170k instances; 1 city Vehicle interactions; trajectory-based safety studies 25s per instance, 10 Hz Yes Limited geographic diversity; 10 object classes (Kesten et al. 2019)
+KITTI 15k scenes; 1 city Early surrogate measures, safety benchmarks 1–2 min sequences, 10 Hz No Classic full-stack dataset, still widely cited (Geiger et al. 2013)
+DAIR-V2X 71k frames Infrastructure-based safety monitoring Yes First large-scale real-world V2X dataset (Yu et al. 2022)
+OPV2V 33k samples; CARLA simulation Cooperative perception, proactive safety design 10 Hz No 18k V2X frames, sim environment (Xu et al. 2022)
+DeepAccident 285k samples; sim dataset Rare crash/near-crash events, proactive safety 10 Hz Yes Includes 57k V2X scenarios (Wang et al. 2024)
+Shift 2.5M frames; 4850 segments; 8 cities Long-term safety and mobility analysis 10 Hz No Corridor-scale benchmarking dataset (Sun et al., 2022)
   • a
-    Sensor: Camera (C), LiDAR (L), Radar (R).
+    Trajectory coverage: Number of driving scenarios and geographic spread.
   • b
-    Tasks: Detection (D), Tracking (T), Segmentation (S), Motion Forecasting (M).
-  • c
-    Hz: Sampling frequency; “–” not reported.
+    Safety relevance: Potential application for crash risk, near-miss, or surrogate safety studies.
 Argoverse-2 was selected after benchmarking open-source AV datasets (Table 1).
@@
 Relative speed 5.257 0.00 23.724 4.446 8.753 0.00 22.276 4.035
-Relative acceleration -0.535 -13.706 10.343 2.227 0.161 -5.925 13.199 0.702
+Relative acc/dec -0.535 -13.706 10.343 2.227 0.161 -5.925 13.199 0.702
 Relative distance 11.673 2.419 46.857 6.844 1.520 0.60 3.80 1.110
@@
 Relative speed 4.513 0.00 33.909 4.278 10.262 0.000 28.067 4.212
-Relative acceleration -0.871 -23.014 18.336 3.139 0.164 -14.067 23.343 1.400
+Relative acc/dec -0.871 -23.014 18.336 3.139 0.164 -14.067 23.343 1.400
 Relative distance 15.361 1.931 141.548 12.773 1.720 0.32 6.56 1.02

### v3

@@
 Abstract
-Real-time prediction of corridor-level crash occurrence risk (COR) remains challenging, as existing near-miss based extreme value models oversimplify collision geometry, exclude vehicle–infrastructure (V–I) interactions, and inadequately capture spatial heterogeneity in vehicle dynamics.
-This study introduces a geometry-aware two-dimensional time-to-collision (2D-TTC) indicator within a Hierarchical Bayesian spatiotemporal grouped random parameter (HBSGRP) framework using a nonstationary univariate generalized extreme value (UGEV) model to estimate short-term COR in urban corridors.
-High-resolution trajectories from the Argoverse-2 dataset, covering 28 locations along Miami’s Biscayne Boulevard, were analyzed to extract extreme near misses (V–V and V–I).
-The model incorporates dynamic variables and road features as covariates, with partial pooling across locations to address unobserved heterogeneity.
-The results show that the HBSGRP–UGEV framework outperforms fixed-parameter alternatives, reducing the DIC by up to 7.5% for V–V and 3.1% for V–I near-misses.
-Predictive validation using ROC–AUC confirms strong performance: 0.89 for V–V segments, 0.82 for V–V intersections, 0.79 for V–I segments, and 0.75 for V–I intersections.
-Model interpretation reveals that relative speed and distance dominate V–V risks at intersections and segments, with deceleration critical in segments, while V–I risks are driven by speed, boundary proximity, and steering/heading adjustments.
-These findings highlight the value of a statistically rigorous, geometry-sensitive, and spatially adaptive modeling approach for proactive corridor-level safety management, supporting real-time interventions and long-term design strategies aligned with Vision Zero.
+Real-time corridor-wide crash occurrence risk (COR) prediction is challenging because existing near-miss EVT models oversimplify collision geometry, neglect vehicle–infrastructure (V–I) interactions, and fail to capture spatial heterogeneity in traffic and roadway conditions.
+To address these gaps, this study introduces a geometry-aware two-dimensional time-to-collision (2D-TTC) indicator for capturing V–V and V–I near misses, and a Hierarchical Bayesian spatiotemporal grouped random parameter (HBSGRP) UGEV model for estimating short-term COR in urban corridors.
+High-resolution trajectories from the Argoverse-2 dataset were analyzed across 28 sites on Miami’s Biscayne Boulevard to extract extreme near-miss events.
+The model incorporates vehicle dynamics and roadway features as covariates, with partial pooling across segments and intersections to capture corridor-wide heterogeneity.
+Results show that the HBSGRP–UGEV framework outperforms fixed-parameter models, reducing DIC by up to 7.5% for V–V and 3.1% for V–I near misses.
+Predictive validation using ROC–AUC confirms strong accuracy (0.89 for V–V segments, 0.82 for intersections, 0.79 for V–I segments, and 0.75 for intersections).
+Random-parameter effects showed that relative speed, distance, and deceleration dominated V–V near misses at segments (only distance for V–I), while intersections were dominated by relative speed and distance for V–V, with no significant effects for V–I..
+These findings demonstrate the value of a geometry-aware, spatially adaptive framework for proactive corridor safety management, supporting both real-time interventions and long-term Vision Zero goals.
 keywords:
@@
 1 Introduction
-The increasing complexity of urban road networks, rising traffic volumes, and the coexistence of human-driven vehicles (HDVs) and autonomous vehicles (AVs) require innovative safety assessment approaches.
-Urban corridors are particularly vulnerable, where a single incident can disrupt mobility, trigger cascading congestion, and impose severe economic losses.
+The increasing complexity of urban corridors, rising traffic volumes, and the coexistence of human-driven vehicles (HDVs) and autonomous vehicles (AVs) demand innovative safety assessment approaches.
+Urban corridors are particularly vulnerable, as even a single crash can disrupt mobility, trigger cascading congestion, and result in substantial economic losses.
 In 2023, traffic crashes in the United States claimed 40901 lives, with a fatality rate of 1.26 per 100 million vehicle miles traveled (VMT) (National Center for Statistics and Analysis, 2025).
-The total annual cost of crashes is estimated to exceed \$1.85 trillion, including \$460 billion in direct economic losses and \$1.4 trillion in reduced quality of life (Blincoe et al., 2023).
-Despite extensive research on crash patterns, risk estimation, and mitigation strategies (Lanzaro et al., 2023; Kamel et al., 2023, 2024; Kamel and Sayed, 2024; Singh et al., 2024; Ghoul and Sayed, 2025; Singh et al., 2025; Anis et al., 2025b), accurate prediction of crash risk in urban corridors remains methodologically underdeveloped.
-Most existing models focus exclusively on vehicle–vehicle (V–V) near-misses, largely due to the availability of vehicle trajectory data, while systematically excluding vehicle–infrastructure (V–I) interactions such as lane edges, curbs, barriers, and crosswalks.
-However, V–I events account for a substantial proportion of severe run-off-road and roadway departure crashes (Islam et al., 2021).
-This omission makes corridor-level safety models inherently incomplete.
-Recent advances in AV sensing and high-definition (HD) mapping, such as the Argoverse-2 dataset (Wilson et al., 2023), now allow the precise integration of roadway boundaries with vehicle trajectories, allowing both V–V and V–I near-misses to be modeled with geometric fidelity.
-Crash risk prediction strategies based on historical retrospective crash data (Arun et al., 2021; Lee et al., 2017; Pei et al., 2011) suffer from well-documented limitations (Lord and Mannering, 2010; Lord et al., 2021; Mannering et al., 2020), including the rarity and stochastic nature of crashes, underreporting (Arun et al., 2021; Tarko, 2018), temporal delays, and spatial mismatches with real-time conditions.
-These challenges motivate a paradigm shift towards proactive trajectory-based surrogate safety measures (SSMs) (Ali et al., 2023a; Arun et al., 2021; Mahmud et al., 2017).
-Using high-resolution vehicle trajectory data, SSMs reflect the underlying near-miss dynamics (Li et al., 2024; Tarko, 2018; Davis et al., 2011).
-Among SSMs, time-based indicators such as Time-to-Collision (TTC), Post-Encroachment Time (PET), and Modified TTC (MTTC) are widely used (Hayward, 1971; Perkins and Harris, (1967; Minderhoud and Bovy, 2001; Ozbay et al., 2008; Allen et al., 1978; Venthuruthiyil and Chunchu, 2022).
-However, these formulations often assume constant velocity and linear trajectories, leading to biased results in boundary-constrained environments.
-Recent research has extended TTC to two-dimensional (2D) indicators of higher-fidelity (Li et al., 2024; Anis et al., 2025b), incorporating latitude and longitude position, heading, and acceleration.
-However, even these advances typically employ simplified circular approximations, which can influence risk estimates in urban contexts where vehicle shape and alignment significantly influence the likelihood and severity of near misses.
-To support a reliable short-term crash occurrence risk (COR) estimate, SSM frameworks must enhance geometric fidelity to more accurately represent real-world interactions between vehicles and infrastructure.
-This study builds on these efforts by introducing a high-fidelity, geometry-aware 2D-TTC indicator that explicitly incorporates vehicle geometry, dynamic states, and roadway boundaries, ensuring realistic and efficient detection of both V–V and V–I near misses.
-While SSMs capture frequent near-misses, a robust COR estimation framework requires extrapolating extreme near-misses to rare event outcomes.
-To date, several studies have proposed probabilistic (Saunier and Sayed, 2008), causal (Davis et al., 2011), and extreme value theory (EVT) (Fu and Sayed, 2021a, b; Zheng et al., 2021b; Ali et al., 2023a; Zheng et al., 2021a; Songchitruksa and Tarko, 2006; Zheng et al., 2014; Tarko, 2012).
-EVT provides a principled statistical foundation for this task, modeling the tail behavior of the extreme near-miss distribution.
-Two EVT approaches dominate: the Generalized Extreme Value (GEV) distribution using Block Maxima (BM) sampling and the Generalized Pareto Distribution (GPD) using Peaks over Threshold (POT) sampling (Coles et al., 2001).
-The BM approach is particularly suitable for real-time analysis due to its short block structure, lower threshold sensitivity, and compatibility with discrete-time samples (Fu and Sayed, 2022a).
-Early applications of EVT in traffic safety employed stationary GEV models (Wang et al., 2018; Guo et al., 2019; Orsini et al., 2019; Alozi and Hussein, 2022; Hussain et al., 2022) that assumed constant parameterization across the study area.
-More recently, non-stationary GEV models have been introduced, allowing covariates to influence location and scale parameters (Tahir and Haque, 2024; Ali et al., 2022; Fu and Sayed, 2022a; Zheng et al., 2019; Nazir et al., 2023; Kar et al., 2024; Ali et al., 2023b; Alozi and Hussein, 2022; Zheng et al., 2014; Kar et al., 2023) and thereby capture heterogeneity across traffic states (Lanzaro et al., 2023; Kamel et al., 2023, 2024; Kamel and Sayed, 2024; Singh et al., 2024; Ghoul and Sayed, 2025; Singh et al., 2025; Anis et al., 2025b).
-Bayesian hierarchical extensions have also addressed data sparsity and unobserved heterogeneity (Zheng and Sayed, 2019; Fu and Sayed, 2022b).
-However, most applications remain geographically narrow and functionally limited to either intersections (Fu and Sayed, 2023, 2021c, 2022b) or segments (Kumar and Mudgal, 2024; Kamel et al., 2023; Kamel and Sayed, 2024; Anis et al., 2025b; Kamel et al., 2024).
-Although some studies have incorporated vehicle dynamics (Kumar and Mudgal, 2024; Fu and Sayed, 2022b; Anis et al., 2025b) and examined heterogeneity across traffic flow directions or roadway geometries (Kamel et al., 2023, 2024), these frameworks generally omit explicit integration of roadway geometry and directional heterogeneity at the corridor level.
-Such integration requires a modeling structure capable of jointly representing vehicle dynamics (which vary temporally) and roadway features (which vary spatially across segments and intersections).
-Therefore, a critical gap remains for a comprehensive Bayesian hierarchical GEV framework capable of jointly modeling vehicle dynamics, roadway geometry, flow directionality, and multilevel heterogeneity.
-This framework is crucial for accurately estimating the COR at the corridor level, which reflects the complex spatiotemporal dynamics of urban traffic environments.
-This paper makes the following unique contributions:
-  • Develops a corridor-level COR framework that simultaneously incorporates both V–V and V–I interactions, addressing a critical omission in previous studies.
-  • Introduces a geometry-aware high-fidelity 2D-TTC indicator that realistically captures extreme near-miss events by accounting for vehicle geometry, dynamic states, and roadway boundaries.
-  • Advances a hierarchical Bayesian spatiotemporal grouped random parameter (HBSGRP) Univariate GEV model that integrates roadway geometry, traffic flow directionality, and multilevel heterogeneity, enabling corridor-wide yet location-sensitive crash risk estimation
-Together, these contributions establish the first infrastructure-aware, statistically rigorous framework for a proactive COR framework in urban corridors, leveraging Arogovers-2 AV trajectory and HD map data to support both operational interventions and long-term safety planning.
+The total annual cost of crashes exceeds \$1.85 trillion, including \$460 billion in direct economic losses and \$1.4 trillion in reduced quality of life (Blincoe et al., 2023).
+Despite decades of research, corridor-wide crash risk prediction remains methodologically limited.
+Crash risk prediction has traditionally relied on retrospective crash data (Arun et al., 2021; Lee et al., 2017; Pei et al., 2011).
+While such approaches have advanced understanding of crash patterns, they face well-documented limitations (Lord and Mannering, 2010; Lord et al., 2021; Mannering et al., 2020), including the rarity and stochastic nature of crashes, underreporting (Arun et al., 2021; Tarko, 2018), temporal delays, and spatial mismatches with real-time traffic conditions.
+These drawbacks highlight the need for proactive approaches that anticipate rather than react to crash risk.
+To address these challenges, researchers have shifted toward proactive trajectory-based surrogate safety measures (SSMs) (Ali et al., 2023a; Arun et al., 2021; Mahmud et al., 2017).
+SSMs use frequent near-miss events as proxies for crashes, enabling the study of traffic conflict dynamics at a much higher temporal resolution (Li et al., 2024; Tarko, 2018; Davis et al., 2011).
+Among these, time-based indicators such as Time-to-Collision (TTC), Post-Encroachment Time (PET), and Modified TTC (MTTC) are widely used (Hayward, 1971; Perkins and Harris, (1967; Minderhoud and Bovy, 2001; Ozbay et al., 2008; Allen et al., 1978; Venthuruthiyil and Chunchu, 2022).
+However, their assumptions of constant velocity and linear paths bias results in boundary-constrained settings.
+Recent two-dimensional (2D) extensions (Li et al., 2024; Anis et al., 2025b) incorporate position, heading, and acceleration, yet still rely on simplified circular vehicle approximations.
+This study addresses these limitations by introducing a geometry-aware 2D-TTC that accounts for vehicle dimensions, dynamics, and roadway boundaries to realistically capture both vehicle–vehicle (V–V) and vehicle–infrastructure (V–I) near misses.
+Building on this foundation, recent studies have integrated SSMs with extreme value theory (EVT) to estimate crash occurrence risk (COR) in urban corridors (Lanzaro et al., 2023; Kamel et al., 2023, 2024; Kamel and Sayed, 2024; Singh et al., 2024; Ghoul and Sayed, 2025; Singh et al., 2025; Anis et al., 2025b).
+EVT extrapolates from frequent near misses to rare crash-prone extremes, offering a statistically principled framework for proactive safety analysis.
+Yet most SSM–EVT applications remain limited, focusing narrowly on V–V interactions and excluding V–I conflicts such as lane edges, curbs, and barriers.
+This omission is critical because V–I events account for a large share of severe roadway departure crashes (Islam et al., 2021).
+Recent advances in AV sensing and high-definition (HD) mapping, such as the Argoverse-2 dataset (Wilson et al., 2023), now enable precise alignment of roadway boundaries with vehicle trajectories, creating new opportunities to model both V–V and V–I near misses with geometric fidelity.
+While SSMs capture frequent near misses, robust COR estimation requires extrapolating extreme near misses to rare crash outcomes.
+EVT (Fu and Sayed, 2021a, b; Zheng et al., 2021b; Ali et al., 2023a; Zheng et al., 2021a; Songchitruksa and Tarko, 2006; Zheng et al., 2014; Tarko, 2012) has become the most widely used method for this task, with two dominant formulations: the Generalized Extreme Value (GEV) distribution using Block Maxima (BM) sampling, and the Generalized Pareto Distribution (GPD) using Peaks-over-Threshold (POT) sampling (Coles et al., 2001).
+The BM approach is particularly well-suited to real-time applications due to its short block structure, reduced threshold sensitivity, and compatibility with discrete-time samples (Fu and Sayed, 2022a).
+Early applications of EVT in traffic safety used stationary GEV models (Wang et al., 2018; Guo et al., 2019; Orsini et al., 2019; Alozi and Hussein, 2022; Hussain et al., 2022), which assumed constant parameters across study areas.
+More recent work has developed nonstationary GEV formulations, allowing covariates to influence distribution parameters (Tahir and Haque, 2024; Ali et al., 2022; Fu and Sayed, 2022a; Zheng et al., 2019; Nazir et al., 2023; Kar et al., 2024; Ali et al., 2023b; Alozi and Hussein, 2022; Zheng et al., 2014; Kar et al., 2023) and capture heterogeneity across traffic states.
+(Lanzaro et al., 2023; Kamel et al., 2023, 2024; Kamel and Sayed, 2024; Singh et al., 2024; Ghoul and Sayed, 2025; Singh et al., 2025; Anis et al., 2025b).
+Bayesian hierarchical extensions further address data sparsity and unobserved heterogeneity (Zheng and Sayed, 2019; Fu and Sayed, 2022b).
+Despite these advances, most applications remain geographically narrow and functionally restricted to either intersections (Fu and Sayed, 2023, 2021c, 2022b) or segments (Kumar and Mudgal, 2024; Kamel et al., 2023; Kamel and Sayed, 2024; Anis et al., 2025b; Kamel et al., 2024).
+Although some studies have integrated vehicle dynamics (Kumar and Mudgal, 2024; Fu and Sayed, 2022b; Anis et al., 2025b) and examined heterogeneity across traffic flow directions or roadway geometries (Kamel et al., 2023, 2024), explicit corridor-wide models that jointly integrate vehicle dynamics, roadway geometry, and directional heterogeneity remain rare.
+To address this gap, this paper makes three contributions: (1) it develops a corridor-wide COR framework that incorporates both V–V and V–I interactions, (2) it introduces a geometry-aware 2D-TTC indicator that captures extreme near-miss dynamics with higher geometric fidelity, and (3) it advances a hierarchical Bayesian spatiotemporal grouped random parameter (HBSGRP) GEV model that integrates roadway geometry, traffic directionality, and multilevel heterogeneity for corridor-wide yet location-sensitive crash risk estimation.
+Together, these contributions establish the first infrastructure-aware, statistically rigorous framework for proactive corridor-wide crash risk prediction, leveraging Argoverse-2 trajectory and HD map data to support both real-time interventions and long-term safety planning.
 The remainder of this paper introduces the proposed 2D-TTC indicators, the HBSGRP–UGEV modeling framework, and the data preparation used in the analysis.
@@
 Scenarios are annotated with agent IDs, AV driving mode, and map-relative lane-following behaviors.
-The sensor suite includes seven ring cameras and two front-facing stereo cameras (20 Hz) supported by dual 64-beam VLP-32C LiDAR sensors, producing  107k 3D points per frame.
+The sensor suite comprises seven ring cameras and two front-facing stereo cameras (20 Hz), supported by dual 64-beam VLP-32C LiDAR sensors, which produce  107k 3D points per frame.
 HD maps contain lane centerlines, crosswalks, medians, and drivable areas registered to local coordinates.
@@
 Scenarios are 11s duration and are sampled at 10 Hz, producing 110 time steps per vehicle for fine-grained interaction modeling.
-At the scenario level, metadata include a unique scenario ID, the recording city, and the focal agent, typically the ego AV.
+At the scenario level, metadata includes a unique scenario ID, the recording city, and the focal agent, typically the ego AV.
 Each scenario then contains multiple agent tracks, where each track corresponds to a road user and specifies the object type (e.g., vehicle, pedestrian, cyclist), scoring category (focal, scored, or unscored), and a time-ordered sequence of states.
@@
 For each track, frame-level kinematic attributes were retrieved, including $(x,y)$ position, heading, velocity components, and timestamps.
-Because positions were reported in local, city-specific frames, all coordinates were transformed to globally referenced WGS84.
+Because positions were reported in local, city-specific frames, all coordinates were transformed into a globally referenced WGS84 system.
 This involved (i) projecting each city’s local origin into its UTM zone (e.g., Zone 17 for Miami, Pittsburgh, and Detroit), and (ii) offsetting local $(x,y)$ positions and inverse-projecting them to latitude–longitude using pyproj.
@@
 Lane configurations vary between 2 and 3 lanes per direction, with average widths ranging from 3.1m (S7) to 3.6m (S6), indicating subtle but important differences in cross-sectional design.
-Driveway and minor intersection density also differs across the corridor: segments such as S3 and S4 contain four access points, while others like S5 and S6 have few or none, creating heterogeneous exposure to side-entry near-misses.
-Median type alternates between raised (e.g., S1–S2) and flush paved (e.g., S3–S9), influencing turning maneuvers and midblock encroachments.
+Driveway and minor intersection density also differ across the corridor: segments such as S3 and S4 contain four access points, while others like S5 and S6 have few or none, creating heterogeneous exposure to side-entry near-misses.
+The median type alternates between raised (e.g., S1–S2) and flush-paved (e.g., S3–S9), influencing turning maneuvers and mid-block encroachments.
 Overall, the table highlights that intersections serve as near-miss-prone nodes with concentrated but lower-volume interactions, whereas segments provide richer samples of continuous traffic flow, characterized by greater geometric variability.
@@
 Prior research has emphasized the role of vehicle kinematics in crash risk (Desai et al., 2021; Jun et al., 2007; Tak et al., 2015); here, additional geometric covariates, such as lane configuration and boundary proximity, are introduced.
-For example, high steering angles, often indicating abrupt lateral maneuvers—were included as predictors, given their link to boundary incursions and sudden evasive actions that may elevate near-miss risk (Gilbert et al., 2021).
+For example, high steering angles, often indicating abrupt lateral maneuvers, were included as predictors, given their link to boundary incursions and sudden evasive actions that may elevate near-miss risk (Gilbert et al., 2021).
 Fig.
@@
 5 Modeling estimation result
-This study develops a series of Bayesian hierarchical UGEV models to estimate COR from extreme 2D-TTC near-miss events observed along Biscayne Boulevard in Miami.
+This study develops a series of Bayesian hierarchical UGEV models to estimate corridor-wide COR from extreme 2D-TTC near-miss events observed along Biscayne Boulevard in Miami.
 Extreme near-misses are identified using the BM approach, wherein the minimum 2D-TTC within each temporal block is treated as the most critical interaction.
-The theoretical foundations and implementation of this sampling methodology are detailed in prior research (2025b).
+The theoretical basis and implementation of this sampling methodology are detailed in prior work (2025b).
 Each V–V and V–I interaction is treated as an independent block.
-Because extreme events are sparse at individual sites, direct location-specific estimation can yield unstable posteriors with wide uncertainty intervals.
-To address this, a grouping strategy was adopted, pooling samples across functionally similar sites (intersections, segments) while still preserving local heterogeneity.
-This clustering approach improves estimation stability and aligns with prior network-level applications of UGEV models in traffic safety (Zheng et al., 2014; Fu and Sayed, 2021c).
-The primary analytical framework is the HBSGRP–UGEV model, while a simpler fixed-effect hierarchical model (HBSFP) was estimated as a benchmark.
-Accordingly, two models were developed for each near-miss type: (1) HBSFP and (2) HBSGRP.
-The grouped random parameter specification allows selected coefficients to vary across spatial clusters, thereby capturing latent differences linked to roadway geometry and interaction dynamics.
-Model performance was evaluated using the DIC (Spiegelhalter et al., 2002; El-Basyouny and Sayed, 2012b, a), where lower values indicate stronger predictive performance (Zheng et al., 2019).
+Since extreme events are relatively sparse at individual sites, direct location-specific estimation often yields unstable posteriors with wide uncertainty intervals.
+To address this issue, a corridor-wide grouping strategy was adopted, pooling samples across functionally similar sites (intersections, segments) while preserving local heterogeneity.
+The HBSGRP–UGEV model is the primary analytical framework, with a simpler fixed-effect hierarchical model (HBSFP) also estimated for comparison.
+For each near-miss type (V–V and V–I), two models were therefore developed: (1) HBSFP and (2) HBSGRP.
+The grouped random parameter formulation in HBSGRP allows selected coefficients to vary across spatial clusters, thereby capturing latent heterogeneity associated with interaction dynamics and roadway characteristics.
+Model performance was evaluated using the DIC (Spiegelhalter et al., 2002; El-Basyouny and Sayed, 2012b, a), with lower values indicating superior predictive fit (Zheng et al., 2019).
 Estimation was conducted using MultiBUGS, a parallelized extension of OpenBUGS tailored for high-dimensional Bayesian inference.
-Its support for block updating and parallel execution makes it well-suited for capturing spatial dependence, random effects, and extreme value behavior within complex data structures.
-Parameters were estimated by Markov Chain Monte Carlo (MCMC) simulation, two parallel MCMC chains of 50000 iterations were run, with the first 20000 discarded as burn-in and the remaining 30000 retained for posterior inference.
-Convergence was verified both visually (trace plots) and quantitatively (Brooks–Gelman–Rubin diagnostic), with all monitored parameters achieving BGR values below 1.1, indicating satisfactory convergence (Gelman and Rubin, 1992; Brooks and Gelman, 1998; El-Basyouny and Sayed, 2009).
-This estimation procedure yielded robust posterior summaries of means, standard deviations, and 95% credible intervals, even under correlation structures induced by spatial clustering and random effects.
-The study incorporated a comprehensive set of covariates into the location ($\mu$) and scale ($\sigma$) parameters of the non-stationary UGEV distribution to reflect contextual and behavioral variability in COR.
-For both V–V and V–I interactions, dynamic covariates included relative speed, distance, acceleration, deceleration, headway difference, steering angle difference, jerk and traffic volume, factors capturing key dimensions of vehicle dynamics and near-miss severity.
-Geometric and infrastructure related covariates were also included: lane number, lane width, driveway intensity, median type, and lane position for segment models; and a categorical indicator of turning movement for intersection models.
-All covariates were specified in both $\mu$ and $\sigma$ components, enabling the models to capture variation in both the central tendency and dispersion of extreme 2D-TTC.
+Its support for block updating and parallel execution makes it well-suited to modeling spatial dependence, random effects, and extreme value behavior within complex data structures.
+Parameters were estimated by Markov Chain Monte Carlo (MCMC) simulation; two parallel MCMC chains of 50000 iterations were run, with the first 20000 discarded as burn-in and the remaining 30000 retained for posterior inference.
+Convergence was assessed both visually (trace plots) and quantitatively (Brooks–Gelman–Rubin diagnostic), with all monitored parameters achieving BGR values below 1.1, confirming reliable convergence (Gelman and Rubin, 1992; Brooks and Gelman, 1998; El-Basyouny and Sayed, 2009).
+This procedure yielded robust posterior summaries of means, standard deviations, and 95% credible intervals, even under correlation structures induced by spatial clustering and random effects.
+The modeling framework incorporated a comprehensive set of covariates into the location ($\mu$) and scale ($\sigma$) parameters of the nonstationary UGEV distribution to capture contextual and behavioral variability in COR.
+For both V–V and V–I interactions, dynamic covariates included relative (speed, distance, acceleration, deceleration, headway, steering angle), jerk, and traffic volume, thereby capturing multiple dimensions of vehicle dynamics.
+Geometric and infrastructure related covariates were also included: lane number, lane width, driveway intensity, median type, and lane position of vehicle.
+All covariates were specified in both the $\mu$ and $\sigma$ components, enabling the models to capture variation in both the central tendency and dispersion of extreme 2D-TTC.
 In contrast, the shape parameter ($\xi$) was held constant across all locations.
-This choice is deliberate: $\xi$ governs tail heaviness in the UGEV distribution but is notoriously difficult to estimate from sparse BM samples.
-Allowing $\xi$ to vary with covariates often induces instability and overfitting (Coles et al., 2001; Cooley et al., 2006).
-Fixing $\xi$ improves convergence and yields more interpretable posterior distributions.
+This decision reflects a tradeoff between flexibility and stability: although $\xi$ governs tail heaviness in the UGEV distribution, it is notoriously difficult to estimate from sparse BM samples.
+Allowing $\xi$ to vary with covariates often induces instability and overfitting (Coles et al., 2001; Cooley et al., 2006), whereas fixing it improves convergence and yields more interpretable posterior distributions.
 5.1 Posterior results and model performance
 Tables 5 through 8 report the posterior summaries for both HBSFP and HBSGRP models, estimated separately for V–V and V–I near-misses at intersections and segments.
-Across all cases, the HBSGRP specification consistently outperformed the HBSFP baseline, confirming its ability to capture unobserved heterogeneity and locatione-level variability.
+Across all cases, the HBSGRP structure consistently outperformed the HBSFP baseline model, confirming its capacity to capture unobserved heterogeneity and location-level variability.
 The DIC values for the HBSGRP models were 508.9 (V–V) and 3179.6 (V–I) at intersections, and 4189.0 (V–V) and 6550.0 (V–I) at segments.
-These correspond to relative reductions in DIC of 3.41% and 0.54% at intersections, and 7.49% and 3.11% at segments, compared with their HBSFP counterparts.
-Such reductions exceed the widely accepted threshold of five DIC units, indicating statistically meaningful improvements in fit (El-Basyouny and Sayed, 2012b, a).
-To further assess model robustness, WAIC and LOOIC were also computed.
-Both measures produced rankings consistent with DIC, reinforcing the superior performance of the HBSGRP structure.
-By allowing selected coefficients to vary across spatial groups, the HBSGRP models capture latent heterogeneity in both the central tendency and dispersion of near-miss severity, which fixed-effect structures fail to represent.
+These correspond to relative reductions of 3.41% and 0.54% at intersections, and 7.49% and 3.11% at segments, compared with their HBSFP counterparts.
+Such improvements exceed the widely accepted threshold of five DIC units, indicating substantively better model fit (El-Basyouny and Sayed, 2012b, a).
+To further evaluate robustness, WAIC and LOOIC were also computed, both of which produced rankings consistent with DIC, thereby reinforcing the superior performance of the HBSGRP.
+By allowing selected coefficients to vary across spatial groups, the HBSGRP models capture latent heterogeneity in both the central tendency and dispersion, a capability absent in fixed-effect structures.
 Collectively, these results demonstrate that grouped random parameter models provide a more flexible and accurate framework for modeling COR dynamics in complex urban traffic environments.
 5.1.1 Intersection-level
-The intersection-level HBSGRP–UGEV model was designed to capture the heterogeneity of extreme 2D-TTC distributions arising from V-V interactions at signalized locations.
-In the model specification stage, several covariates were tested as random effects; however, most roadway geometry variables exhibited negligible variation across sites and did not improve posterior fit.
-To maintain parsimony and interpretability, random effects were therefore restricted to the intercept and two dynamic covariates (relative speed and relative distance), while other vehicle dynamics and geometric features were retained as fixed effects, following recommendations from (Anis et al., 2025a).
+The intersection-level HBSGRP–UGEV model was designed to capture the heterogeneity of extreme 2D-TTC distributions from V-V interactions at signalized intersections.
+In the modeling stage, all covariates were considered as random effects; however, roadway characteristics and few vehicle dynamics were exhibited negligible variation across sites and did not improve posterior fit.
+To maintain parsimony and interpretability, random effects were restricted to the intercept and two dynamic covariates (relative speed and distance), while other covariates were retained as fixed effects, following recommendations from (Anis et al., 2025a).
 This choice reflects theoretical expectations: vehicle dynamics are highly variable across intersections depending on signal phasing, approach alignment, and driver behavior, whereas roadway structure tends to exert a more stable influence.
@@
     Indicates statistical significance at the 95% level (interval excludes 0)
-The fixed-parameter benchmark (HBSFP) provides a baseline against which the grouped random specification can be compared.
-Consistent reductions in DIC, WAIC, and LOOIC confirm the superior fit of the HBSGRP model, underscoring the explanatory value of allowing relative speed and spacing effects to vary across intersection clusters.
-The direction and magnitude of these coefficients provide valuable insights into the mechanisms underlying COR.
-The random-parameter specification highlights the spatially varying influence of relative speed and relative distance on both the location and scale components of the HBSGRP distribution.
-For $\mu$, relative speed shows a positive effect across intersections, indicating that larger speed differentials between interacting vehicles lead to sharply reduced 2D-TTC values and more severe near-miss conditions.
-In contrast, relative distance is negatively associated with $\mu$, meaning that shorter headways systematically increase near-miss severity by limiting the buffer available for evasive action.
-For $\sigma$, both covariates are associated with reduced variability in near-miss severity.
-Relative speed has a negative coefficient, suggesting that high-speed interactions produce consistently critical outcomes, dangerous but less variable, because drivers have little opportunity to recover through braking or swerving.
-Similarly, relative distance exerts a negative effect on $\sigma$, reflecting that short headways yield uniformly critical near-miss with little dispersion in outcomes.
-Together, these results confirm the physical intuition that speed amplifies severity while distance compresses variability, producing deterministic near miss outcomes under constrained conditions.
+Consistent reductions in DIC, WAIC, and LOOIC confirm the superior fit of the HBSGRP model, highlighting the added value of allowing speed and spacing effects to vary across intersections.
+The direction and magnitude of these coefficients provide insights into the mechanisms underlying COR.
+Relative speed and distance emerge as the dominant dynamics, influencing both the location ($\mu$) and scale ($\sigma$) parameters.
+For $\mu$, relative speed has a positive effect, indicating that larger speed differentials between vehicles sharply reduce 2D-TTC and intensify near-miss severity.
+In contrast, relative distance shows a negative effect, meaning shorter headways systematically elevate severity by limiting available buffer space for evasive action.
+For $\sigma$, both covariates carry negative signs, implying that high speeds and short distances compress variability, yielding consistently critical outcomes with little room for driver recovery through braking or swerving.
+Taken together, these results confirm that speed amplifies severity while short headways exacerbate risk and reduce variability, producing deterministic near-miss outcomes under constrained conditions.
 Fig.
-5.1.1 further illustrate this spatial heterogeneity across intersection, while most locations show consistent trends, the I2 site exhibits unusually wide CI for the $\mu$ effects and CI exceed zero for the $\sigma$, across intercept, relative speed, and relative distance.
-This instability is directly attributable to I2’s small BM sample size, which weakens posterior precision and introduces greater uncertainty in estimated group-level effects.
-Despite this, the overall pattern across intersections remains robust: relative speed consistently intensifies near-miss severity, while shorter distances exacerbate risk and constrain outcome variability.
-Beyond the random effects, several covariates were modeled as fixed parameters due to their relatively stable contributions across intersections.
-Their interpretation reveals distinct roles of longitudinal dynamics, lateral control, maneuver type, and exposure effects in shaping near-miss severity and variability.
-Table 5 indicate that relative acceleration shows a significant negative effect on both the location and scale parameters.
-This suggests that acceleration is generally associated with smoother and less variable near-miss, consistent with situations where vehicles accelerate out of a green phase or clear an intersection without abrupt interaction.
-By contrast, relative deceleration exerts a positive effect on severity ($\mu$) but a negative effect on variability ($\sigma$).
-This indicates that abrupt braking sharply intensifies near-misses but also produces more uniform outcomes, reflecting the deterministic nature of emergency deceleration in high-risk encounters.
-Jerk, which captures the rate of change in acceleration, has the opposite pattern: a negative effect on $\mu$ but a positive effect on $\sigma$.
-Although jerk slightly reduces average near-miss severity (perhaps due to hesitation or driver corrections), it simultaneously increases unpredictability in outcomes, highlighting the unstable nature of rapid throttle/brake oscillations.
-Lateral vehicle dynamics also contribute meaningfully to near-miss risk estimation.
-Steering difference emerges as a significant contributor in the HBSGRP specification, showing a positive association with $\mu$ and a negative association with $\sigma$.
-This suggests that abrupt steering maneuvers intensify near-miss severity but lead to more consistent outcomes within constrained intersection geometries.
-Heading difference also shows a significant positively association with $\mu$ and a negatively with $\sigma$.
-Indicating that even small misalignments in trajectory lead to predictable patterns of near-miss outcomes, consistent with the rigid geometry of signalized approaches.
-The categorical indicator for left-turn movements is significant and has a positive effect on both $\mu$ and $\sigma$.
-This aligns with the well-documented complexity of left-turn phases, which involve interactions with opposing through traffic and diverse driver decision strategies.
-The result suggests that left turns not only elevate near-miss severity but also introduce greater variability in outcomes, reflecting the wide range of intersection designs and driver behaviors under turning conditions.
-Traffic volume is statistically insignificant in both $\mu$ and $\sigma$.
-Nonetheless, it was retained in the model to ensure conceptual completeness, since it represents exposure.
-The lack of effect implies that short-duration 2D-TTC near-miss events are governed more by instantaneous vehicle interactions than by aggregate traffic demand.
+5.1.1 illustrates the spatial heterogeneity across intersections.
+While most locations show consistent trends, the I2 site exhibits unusually wide credible intervals for the $\mu$ effects and intervals exceeding zero for the $\sigma$ across the intercept, relative speed, and relative distance.
+This instability is attributable to I2’s small BM sample size, which weakens posterior precision and increases uncertainty in group-level effects.
+Despite localized uncertainty, the dominant intersection-level trend persists, with relative speed and distance exerting stable and significant effects on near-miss outcomes.
+Beyond random effects, several covariates were modeled as fixed parameters due to their stable contributions across intersections.
+These effects can be interpreted in terms of driver behavior.
+These factors highlight the roles of longitudinal dynamics, lateral control, maneuver type, and exposure in shaping near-miss outcomes.
+Table 5 shows that relative acceleration has a significant negative effect on both the location and scale parameters, suggesting smoother and less variable near misses when vehicles accelerate through green phases or clear intersections without abrupt interaction.
+By contrast, relative deceleration increases severity ($\mu$) but reduces variability ($\sigma$), capturing emergency braking events.
+These situations sharply intensify near misses but also produce uniform outcomes, since drivers in crisis have limited response options.
+Jerk, representing the rate of change in acceleration, shows the opposite pattern: a negative effect on $\mu$ but a positive effect on $\sigma$.
+This suggests that rapid throttle–brake oscillations, often signs of hesitation or overcorrection, slightly lower average severity but introduce greater unpredictability, underscoring the instability of abrupt corrective maneuvers.
+Lateral vehicle dynamics also play an important role in near-miss risk estimation.
+Steering difference is a significant factor in the HBSGRP model, showing a positive association with $\mu$ and a negative association with $\sigma$.
+This indicates that abrupt steering maneuvers intensify near-miss severity but produce more consistent outcomes within constrained intersection geometries.
+Heading difference shows a similar pattern, with a positive effect on $\mu$ and a negative effect on $\sigma$.
+This suggests that even small misalignments in vehicle trajectory elevate severity while reducing variability, reflecting the rigid geometric constraints of signalized approaches.
+The categorical indicator for left-turn movements is significant, with positive effects on both $\mu$ and $\sigma$.
+This finding is consistent with the well-documented complexity of left-turn phases, which involve interactions with opposing through traffic and diverse driver decision strategies.
+Left turns therefore not only elevate near-miss severity but also increase variability, reflecting differences in intersection design and driver behavior under turning conditions.
+By contrast, traffic volume is statistically insignificant for both parameters but was retained in the model to represent exposure.
+Its lack of effect suggests that short-duration extreme 2D-TTC events are driven more by instantaneous vehicle interactions than by aggregate demand.
 Finally, the estimated shape parameter ($\xi$) is negative and significant across all intersections (Fig.
@@
 Importantly, group-level variability in $\xi$ suggests that intersection design and control strategies influence not only the frequency but also the extremal nature of near-miss outcomes.
-The HBSGRP model for V–I near-misses specifies the intercept as a random parameter in both the location and scale components, capturing baseline heterogeneity in near-miss severity and variability across intersections.
-Relative speed (location) and relative distance (scale) were also treated as random effects (Fig.
-5.1.1), though their 95% CI overlapped zero.
-They were retained on theoretical grounds, given their inclusion improve overall metrics.
-All other covariates were modeled as fixed effects, representing more stable behavioral and geometric influences.
-Fit statistics (Table 6) show modest but consistent improvements in DIC, WAIC, and LOOIC over the HBSFP benchmark, suggesting that the grouped random parameter structure adds nuance but yields smaller performance gains compared to the V–V case.
+The HBSGRP model for V–I near misses specified the intercept as a random parameter in both the location and scale components, capturing baseline heterogeneity in severity and variability across intersections.
+Relative speed (in $\mu$) and relative distance (in $\sigma$) were also modeled as random effects (Fig.
+5.1.1), although their 95% credible intervals overlapped zero.
+They were nevertheless retained on theoretical grounds, as their inclusion improved overall model fit.
+This specification reflects behavioral intuition: heterogeneity in speed effects is more pronounced in severity (location), whereas heterogeneity in clearance effects is more evident in variability (scale).
+All other covariates were treated as fixed effects, reflecting their comparatively stable behavioral and geometric influences.
+Fit statistics (Table 6) show modest but consistent improvements in DIC, WAIC, and LOOIC relative to the HBSFP specification, suggesting that the grouped random parameter structure adds explanatory nuance but produces smaller performance gains than in the V–V case.
 Table 6: Posterior estimates for V-I near misses at intersections
@@
 In the location parameter, several interpretable effects emerge.
-Relative distance is strongly negative and significant, confirming that reduced clearance to the roadway edge sharply lowers 2D-TTC, elevating severity.
-Relative deceleration also carries a significant negative sign, which contrasts with its positive effect in V–V near-miss.
-This suggests that braking near infrastructure reflects proactive yielding or evasive action, thereby limiting severity rather than escalating it.
+Relative distance is strongly negative and significant, confirming that reduced clearance to the roadway edge sharply lowers 2D-TTC and elevates severity.
+Relative deceleration also enters with a significant negative sign, contrasting with its positive effect in V–V near misses.
+This suggests that braking near infrastructure often reflects proactive yielding or evasive maneuvers, which mitigate severity rather than intensify it.
 Heading and steering differences are positive and significant, indicating that lateral deviations (e.g., turning toward or away from curbs) increase near-miss severity.
-Left-turn movement, however, shows a significant negative association, suggesting that left-turn trajectories may provide greater clearance to boundaries, reducing average severity compared to right turns.
-This finding requires careful interpretation, as it may partly reflect geometric factors such as wider turning radii or reduced pedestrian/boundary overlap in left-turn phasing.
-Relative speed, though included as a random effect, is not significant in the pooled analysis, highlighting that V–I risks are less sensitive to speed differentials than V–V encounters.
-The scale parameter results add further insight into variability.
-The negative random intercept suggests baseline differences across intersections, with some sites producing more predictable outcomes than others.
-Relative speed exhibits a positive and significant association with scale, implying that faster approaches to infrastructure yield greater variability in severity—contrasting with the V–V case, where speed reduced dispersion.
-This difference reflects the inherently diverse driver responses to boundary proximity: some brake aggressively, while others take wider trajectories or maintain speed.
-Relative acceleration is also positive and significant, reinforcing the idea that heterogeneous driver behavior during acceleration near boundaries widens outcome variability.
-In contrast, relative deceleration has a significant negative effect on scale, meaning that strong braking compresses the severity distribution into consistently critical outcomes.
-Jerk is similarly positive, confirming that abrupt longitudinal control amplifies variability.
-Lateral dynamics again exhibit distinct roles: heading and steering differences reduce variability, suggesting that once a driver commits to a lateral maneuver, outcomes become more predictable.
-Finally, left-turn movement increases variability despite lowering average severity, consistent with heterogeneous geometric layouts and driver strategies in left-turn contexts.
+Left-turn movement, by contrast, shows a significant negative association, suggesting that left-turn trajectories provide greater clearance to boundaries and reduce severity compared with right turns.
+This pattern may reflect geometric characteristics such as wider turning radii or reduced overlap with curbs and crosswalks during left-turn phasing.
+Finally, relative speed—although modeled as a random effect—does not reach significance in the pooled analysis, reinforcing that V–I severity is less sensitive to speed differentials than V–V encounters.
+The scale parameter results provide additional insight into variability.
+The negative random intercept indicates baseline differences across intersections, with some sites producing more predictable outcomes than others.
+Relative speed shows a positive and significant association with $\sigma$, implying that faster approaches to infrastructure increase variability in severity—in contrast to the V–V case, where speed compressed dispersion.
+This divergence reflects the diverse ways drivers respond to boundary proximity: some brake aggressively, while others maintain speed or adjust trajectory.
+Relative acceleration is also positive and significant, reinforcing that heterogeneous acceleration behavior near boundaries broadens outcome variability.
+By contrast, relative deceleration has a significant negative effect, indicating that decisive braking compresses the distribution into consistently critical outcomes.
+Jerk is similarly positive, confirming that abrupt throttle–brake oscillations amplify variability.
+Lateral dynamics play the opposite role: heading and steering differences both reduce variability, suggesting that once a lateral maneuver is committed, outcomes become more predictable within intersection geometry.
+Finally, left-turn movement increases variability despite lowering mean severity, consistent with heterogeneous layouts and driver strategies in left-turn contexts.
 The shape parameter remains negative and significant across intersections, supporting the Weibull domain for V–I near misses and confirming that these events have a bounded upper tail.
-This implies a physical or behavioral limit to how severe boundary near-misses can become, constrained by roadway geometry and driver response capacity.
-Spatial variation in the shape parameter underscores the role of intersection design and infrastructure placement in shaping the extremal properties of near-miss distributions.
-Overall, the intersection-level results suggest that V–I near-misses are driven more by clearance, braking behavior, and lateral maneuvers than by speed differentials.
-Unlike V–V near-misses, where high speeds produce deterministic severity, V–I cnear-misses are marked by greater variability, reflecting the complex interplay between infrastructure geometry, driver avoidance strategies, and intersection control.
-5.2 Segment-level
-The HBSGRP model for segment-level V–V near misses introduces random parameters within a hierarchical framework (Table 7).
-Allowing the intercept and three time-varying vehicle-based measures, relative speed, relative deceleration, and relative distance, to vary across segments (See Fig.
-5.2).
-This specification acknowledges that longitudinal dynamics in uninterrupted midblock contexts are shaped not only by vehicle behavior but also by geometric features such as lane configuration, median presence, and driveway density.
-All remaining covariates, including relative acceleration, jerk, heading and steering differences, traffic volume, lane attributes, and access density, were modeled as fixed effects due to their comparatively stable influence across segments.
-The HBSGRP specification produces substantial improvements in model fit over the HBSFP baseline, confirming the value of partial pooling in capturing heterogeneity across the nine directional segments analyzed.
-For the location parameter ($\mu$), random effects highlight systematic heterogeneity across the nine directional segments.
-Relative speed exerts a significant positive effect, confirming that higher closing speeds consistently compress available 2D-TTC and elevate near-miss severity.
-However, the posterior distributions in Fig.
-5.2 reveal that the magnitude of this effect varies substantially across segments.
-Segments with uninterrupted, high-speed flows (e.g., S11) show stronger effects, while others with denser access points (e.g., S6, S7 and S13) exhibit wider credible intervals.
-Relative deceleration also carries a negative sign, though with more overlap in its credible intervals across segments, suggesting that emergency braking may sometimes mitigate severity in smoother-flow environments, while in higher-speed segments it intensifies near-misses.
-Notably, in some lower-sample segments, credible intervals for the location intercept approach or cross zero, reflecting weaker estimation stability under sparse BM events.
-Relative distance displays a consistently negative effect across segments, meaning that shorter headways reliably reduce 2D-TTC, yet its posterior variation underscores differences in spacing behavior along divided versus undivided facilities.
+This indicates a physical and behavioral ceiling on severity, constrained by roadway geometry and driver response capacity.
+Spatial variation in $\xi$ further underscores the influence of intersection design and infrastructure placement on the extremal properties of near-miss distributions.
+Taken together, the intersection-level results suggest that V–I near misses are shaped more by clearance, braking, and lateral maneuvers than by speed differentials.
+In contrast, V–V near misses are dominated by high speeds, which produce deterministic severity, whereas V–I outcomes exhibit greater variability driven by the interplay of infrastructure geometry, driver avoidance strategies, and intersection control.
+5.1.2 Segment-level
+The HBSGRP model for segment-level V–V near misses introduces random parameters within a hierarchical structure (Table 7).
+Specifically, the intercept and three time-varying vehicle dynamics—relative speed, relative deceleration, and relative distance—were allowed to vary across segments (Fig.
+5.1.2).
+This specification recognizes that longitudinal interactions in uninterrupted midblock contexts are shaped by both driver behavior and roadway characteristics (lane configuration, median type, driveway density).
+All remaining covariates—relative acceleration, jerk, heading and steering differences, traffic volume, lane attributes, and access density—were treated as fixed effects given their comparatively stable influence across segments.
+The HBSGRP model yields substantially better fit than the fixed-effect alternative, confirming the value of partial pooling for capturing heterogeneity across the nine directional segments.
 Refer to caption Refer to caption
@@
     Indicates statistical significance at the 95% level (interval excludes 0)
-Among the fixed effects, relative acceleration shows a significant negative coefficient in both models, implying that acceleration phases are systematically linked to higher near-miss severity on segments.
-This contrasts with intersection contexts, where acceleration often reflected flow resumption.
-Here, aggressive acceleration during overtaking or merging maneuvers reduces spacing and increases risk.
-Jerk is also negative, reinforcing the role of abrupt longitudinal control in producing critical outcomes.
-Lateral control indicators—heading and steering differences—show positive signs, suggesting that lateral adjustments modestly increase 2D-TTC where lane-changing flexibility exists.
-Driveway density exhibits a strong positive effect on $\mu$, which may appear counterintuitive, but is likely explained by anticipatory driver caution in access-rich environments.
-Conversely, undivided medians sharply reduce $\mu$, indicating that exposure to opposing flows creates consistently severe interactions.
-For the scale parameter ($\sigma$), the random effects for relative speed, deceleration, and distance are all negative, showing that extreme dynamics compress outcome variability, leading to uniformly critical near misses.
-The plots confirm this: high-speed and short-spacing conditions cluster tightly around lower 2D-TTC values, with narrow posteriors in segments such as S12 and S15.
-In contrast, segments like S10 display wider CI for random scale effects, again reflecting limited BM sample sizes.
-Fixed effects add further nuance: jerk and heading difference increase scale, suggesting that volatile acceleration or lateral corrections introduce heterogeneity in outcomes, while lane number and lane width reduce scale, indicating that greater maneuvering space fosters more predictable near-misses.
-Driveway density increases dispersion, consistent with the variability introduced by frequent merging and crossing maneuvers.
-Median type again exerts a strong negative influence, reinforcing that the absence of raised medians compresses severity outcomes into consistently critical ranges.
-Finally, the shape parameter ($\xi$) remains negative and significant across segment groups, confirming a bounded upper tail consistent with the Weibull domain.
-This indicates physical and behavioral limits to maximum 2D-TTC values during critical segment near-misses, constrained by speed environment, geometry, and driver reaction times.
-Posterior variation in the shape intercept suggests that undivided segments experience steeper extremal limits, while divided facilities allow marginally greater spread in outcomes.
-Overall, the segment-level results highlight the primacy of longitudinal dynamics in shaping corridor risk, with acceleration, braking, and headway dominating near-miss severity, while roadway design elements such as medians, lane width, and access density govern both severity and variability.
-The superior fit of the HBSGRP model underscores the importance of capturing these segment-specific dynamics, as fixed-parameter structures fail to account for the spatially contingent influence of speed and spacing across midblock contexts.
-The HBSGRP specification for segment-level V–I near-misses provides a detailed view of how longitudinal dynamics and roadway design influence the severity and variability of extreme 2D-TTC outcomes (Table 8, Fig.
-5.2).
-In this model, the intercept and relative distance are treated as group-level random effects in both the location and scale components of the HBSGRP model, allowing for spatial heterogeneity in how spacing between interacting vehicles shapes COR across segments.
-All other predictors—including relative speed, acceleration, deceleration, jerk, heading difference, steering angle difference, traffic volume, lane attributes, driveway density, median type, and left-lane presence, are modeled as fixed effects, reflecting their expected stability across segment environments.
+For the location parameter ($\mu$), the random effects reveal meaningful differences across segments that align with expected driving dynamics.
+Relative speed shows a consistently positive effect, confirming that faster closing speeds sharply reduce 2D-TTC and intensify near-miss severity.
+The magnitude of this effect, however, varies across roadway contexts Fig.
+5.1.2.
+Segments with uninterrupted high-speed flows (e.g., S11) show the strongest impacts, while access-rich segments (e.g., S6, S7, and S13) display wider credible intervals, reflecting greater variability in interactions.
+Relative deceleration carries a negative sign overall, but its effect is less stable, suggesting that braking sometimes helps dissipate conflicts in smoother-flow environments, whereas in higher-speed segments the same braking often signals critical near-miss events.
+Relative distance exhibits a consistently negative association, indicating that shorter headways reliably elevate severity by restricting the available buffer for evasive maneuvers.
+The posterior variation across segments suggests that spacing behavior differs between divided and undivided facilities.
+Finally, in some segments with sparse block maxima samples, the intercept’s credible intervals approach zero, highlighting weaker estimation stability under limited data conditions.
+Among the fixed effects, relative acceleration shows a significant negative association with $\mu$, indicating that acceleration phases are systematically linked to more severe near-misses on segments.
+This pattern contrasts with intersections, where acceleration often reflects flow resumption under green signals.
+Along segments, however, aggressive acceleration during overtaking or merging reduces spacing and elevates risk.
+Jerk also exhibits a negative effect, underscoring how abrupt longitudinal control contributes to critical outcomes.
+Lateral control indicators—heading and steering differences—display positive coefficients, suggesting that lateral adjustments provide modest increases in 2D-TTC where lane-changing flexibility is available.
+Driveway density is positively associated with $\mu$, a result that may seem counterintuitive but likely reflects anticipatory caution by drivers in access-rich environments.
+In contrast, undivided medians show a strong negative effect, confirming that exposure to opposing traffic flows produces consistently more severe interactions.
+For the scale parameter ($\sigma$), the random effects for relative speed, deceleration, and distance are all negative, indicating that extreme longitudinal dynamics compress outcome variability and yield uniformly critical near-miss conditions.
+This pattern is evident in Fig.
+5.1.2, where high-speed and short-spacing conditions cluster tightly around lower 2D-TTC values, producing narrow posteriors in segments such as S12 and S15.
+By contrast, segments like S10 display wider credible intervals, reflecting weaker estimation stability under sparse BM samples.
+Among the fixed effects, jerk and heading difference increase $\sigma$, suggesting that volatile acceleration or lateral corrections introduce greater heterogeneity in near-miss outcomes.
+In contrast, lane number and lane width reduce $\sigma$, implying that additional maneuvering space fosters more predictable interactions.
+Driveway density increases dispersion, consistent with the variability created by frequent merging and crossing maneuvers.
+Median type again shows a strong negative effect, reinforcing that undivided facilities compress severity outcomes into consistently critical ranges.
+Finally, the shape parameter ($\xi$) remains negative and statistically significant across segment groups, confirming a bounded upper tail consistent with the Weibull domain.
+This finding indicates physical and behavioral constraints on extreme near-miss severity, governed by speed environment, roadway geometry, and driver reaction capacity.
+Overall, the segment-level results emphasize the central role of longitudinal dynamics—acceleration, braking, and headway—in determining near-miss severity, while roadway features such as medians, lane width, and access density primarily shape variability.
+The strong performance of the HBSGRP model underscores its value in capturing these segment-specific dynamics, which fixed-parameter structures fail to represent, particularly in accounting for the spatially contingent effects of speed and spacing in midblock contexts.
+The HBSGRP model for segment-level V–I near-misses offers insights into how longitudinal dynamics and roadway design jointly shape the severity and variability of extreme 2D-TTC outcomes (Table 8, Fig.
+5.1.2).
+In this model, the intercept and relative distance are specified as group-level random effects in both the location and scale components, capturing spatial heterogeneity in how spacing influences near-miss severity across segments.
+All other covariates, including relative speed, acceleration, deceleration, jerk, heading difference, steering angle difference, traffic volume, lane attributes, driveway density, median type, and left-lane presence—are modeled as fixed effects to represent their more consistent contributions across segment environments.
 (a) $\beta_{\mu,0[1,18]}$
@@
     Indicates statistical significance at the 95% level (interval excludes 0)
-For the location parameter, shorter relative distances strongly reduce 2D-TTC values, confirming that closer spacing consistently produces more critical near-miss outcomes.
-The random effect for distance further indicates that influence varies by segment group, suggesting sensitivity to local geometric or operational conditions.
-Relative speed displays a notable divergence: while the fixed-parameter benchmark shows a negative coefficient, the grouped model produces a small but positive estimate, implying that in certain contexts, higher speeds are associated with marginally longer 2D-TTC values.
-This reversal highlights that segment-level heterogeneity in speed environments, such as free-flow arterials versus constrained multilane facilities, can fundamentally alter the role of speed in shaping severity.
-Relative acceleration emerges as a significant positive effect in the grouped specification, suggesting that acceleration phases along segments may often represent clearance or recovery maneuvers rather than heightened near-miss risk.
-In contrast, relative deceleration maintains a negative coefficient across both models, reflecting the reality that braking during close-following conditions is tied to critical reductions in 2D-TTC, likely triggered by sudden evasive actions.
-Several geometric and operational features are also important for location.
-Heading and steering differences exhibit significant positive coefficients, indicating that lateral adjustments yield small gains in clearance and mitigate near-miss severity.
-Volume, lane number, and lane width also emerge as positive contributors in the HBSGRP specification, suggesting that higher-capacity, wider facilities afford drivers greater maneuvering space, leading to slightly larger 2D-TTC values.
-The effect of driveway density, however, is particularly striking: while insignificant in the fixed-parameter model, the grouped specification produces a large negative estimate with a wide CI.
-This instability suggests that driveway effects are highly localized—segments with dense access points may concentrate more severe near-miss outcomes, but the data are sparse and sensitive to segment-specific characteristics.
-Median type is also significant, with undivided facilities consistently producing lower 2D-TTC values, reflecting heightened exposure to opposing traffic flows.
+For the location parameter, reduced clearance to roadway boundaries exerts a strong negative effect on 2D-TTC, confirming that closer spacing to infrastructure consistently produces more critical near-miss outcomes.
+The random effect for distance further indicates that this influence varies across segment groups, reflecting sensitivity to local geometric and operational conditions such as median presence or driveway density.
+Relative speed shows a notable divergence: while the fixed-parameter model suggests higher speeds intensify severity, the grouped specification produces a small positive estimate, implying that in some contexts faster travel may coincide with greater clearance from boundaries (e.g., drivers maintaining lane center on free-flow arterials).
+This reversal highlights how heterogeneous speed environments fundamentally alter the role of speed in shaping V–I severity.
+Relative acceleration is positive and significant, suggesting that acceleration phases often represent clearance from roadside constraints (e.g., moving away from curbs or barriers) rather than heightened risk.
+By contrast, relative deceleration retains a negative effect across both models, indicating that braking near boundaries is typically associated with abrupt evasive actions, sharply reducing 2D-TTC and producing more critical near-miss conditions.
+Several geometric and operational features also shape V–I near-miss severity at the segment level.
+Heading and steering differences exhibit significant positive coefficients, indicating that lateral adjustments toward or away from boundaries provide modest clearance gains and mitigate severity.
+Traffic volume, lane number, and lane width also emerge as positive contributors in the HBSGRP specification, suggesting that wider, higher-capacity segments allow greater separation from roadside features, producing slightly larger 2D-TTC values.
+The effect of driveway density, however, is particularly notable: while insignificant in the fixed-parameter model, the grouped specification yields a large negative estimate with a wide CI.
+This suggests that dense roadside access points intensify boundary conflicts by forcing frequent merging and turning maneuvers, though the data remain sensitive to localized characteristics.
+Median type is also significant, with undivided facilities consistently producing lower 2D-TTC values, reflecting heightened exposure to opposing flows and limited recovery space near boundaries.
 The left-lane variable shows a sign reversal, with the fixed model suggesting higher severity in interior lanes and the grouped model indicating greater severity in outer lanes.
-This again illustrates heterogeneity across segment groups, where the positional context of a lane relative to roadside boundaries and merging activity shifts the locus of risk.
-The scale parameter provides complementary insights into the dispersion of near-miss severity.
-The random effect for relative distance is positive, showing that tighter spacing increases not only severity but also variability in outcomes, as different drivers adopt divergent strategies under minimal headway conditions.
-Relative speed also has a positive association with scale, indicating that higher speeds amplify heterogeneity in outcomes across V–I interactions, unlike in V–V segment near-misses, where speed reduces dispersion.
-Acceleration, by contrast, has a negative coefficient in the grouped model, suggesting that acceleration phases yield more predictable outcomes, consistent with smoother-flow recovery dynamics.
-Relative deceleration also reduces variability, implying that braking in near-crash contexts converges toward uniformly critical outcomes.
-Geometric attributes reveal nuanced effects on variability.
-Lane width exerts a positive effect, suggesting that while wider lanes modestly reduce average severity, they also introduce greater variability in outcomes by enabling diverse driver responses.
-The undivided median type strongly increases scale, underscoring that undivided facilities not only heighten the average risk but also broaden the range of severity outcomes, likely due to variation in driver avoidance maneuvers.
-The left-lane effect is positive in scale, indicating that severity outcomes in these lanes are more heterogeneous, perhaps reflecting the diversity of driver strategies in navigating outer versus inner lanes.
-Driveway density, although statistically unstable, shows a wide and imprecise effect that suggests variability is strongly contingent on site-specific access management patterns.
+This pattern highlights heterogeneity across segment groups, where lane position relative to curbs, medians, or roadside access alters the likelihood of boundary-related near misses.
+The scale parameter provides complementary insights into the variability of boundary-related near misses.
+The random effect for relative distance is positive, indicating that reduced clearance to roadside features not only increases severity but also broadens variability, as drivers adopt divergent strategies when negotiating minimal spacing with curbs, medians, or barriers.
+Relative speed also shows a positive association with scale, suggesting that higher speeds amplify heterogeneity in boundary interactions—contrasting with V–V near misses, where speed reduces dispersion—because drivers respond differently to infrastructure proximity under fast-moving conditions.
+Acceleration, by contrast, has a negative coefficient in the grouped model, implying that acceleration phases near boundaries often reflect smoother clearance maneuvers, yielding more predictable outcomes.
+Relative deceleration also reduces variability, indicating that braking close to infrastructure compresses outcomes into uniformly critical events.
+Geometric attributes further shape variability.
+Lane width exerts a positive effect, showing that while wider lanes provide greater average clearance, they also increase variability by enabling a broader range of lateral positioning relative to boundaries.
+The undivided median type strongly increases scale, underscoring that the absence of a raised median not only heightens risk but also expands outcome variability due to heterogeneous driver avoidance maneuvers.
+The left-lane variable is also positive, indicating that boundary proximity in outer lanes produces more heterogeneous near-miss outcomes compared to interior lanes.
+Finally, driveway density, though statistically unstable, suggests that variability is strongly site-dependent, as frequent access points create unpredictable driver–infrastructure interactions.
 The shape parameter remains negative and significant across both specifications, confirming that the upper tail of the 2D-TTC distribution is bounded.
-This supports the Weibull characterization of extremes, implying that even in segment-level V–I interactions, physical and behavioral limits constrain how severe the most critical near-miss events can be.
-Variation in shape across segment groups highlights the role of local geometry, access density, and operational features in shaping the extremal properties of near-miss distributions.
-Together, the segment-level analyses of V–V and V–I near misses underscore that roadway geometry and vehicle dynamics interact to shape both the severity and variability of extreme near-misses.
-For V–V events, relative speed and distance dominate as key random effects, with speed consistently amplifying severity and distance compressing variability, reflecting deterministic outcomes under close-following conditions.
-For V–I events, however, the role of speed is more heterogeneous—sometimes protective, sometimes risk-enhancing—depending on the segment context.
-Access-related variables, particularly driveway density and median type, emerge as critical for V–I outcomes, underscoring the importance of access management and median design in shaping roadside safety.
-In both near-miss types, the negative and significant shape parameter indicates bounded extremes, reinforcing the physical and behavioral limits to crash risk escalation.
+This supports the Weibull domain for extremes and indicates that even in V–I segment interactions, physical and behavioral limits constrain how severe the most critical boundary-related near-misses can become.
+Variation in the shape parameter across segment groups underscores the role of local geometry, access density, and operational conditions in shaping the extremal behavior of these events.
+Segment-level analyses show that roadway geometry and vehicle dynamics jointly shape near-miss severity and variability.
+For V–V events, relative speed and distance dominate—speed amplifies severity while distance compresses variability, producing more deterministic outcomes under close-following conditions.
+For V–I events, the effect of speed is heterogeneous, sometimes reducing severity when clearance to boundaries is maintained, but elevating it when high speeds bring vehicles closer to curbs or medians.
+Access-related factors, particularly driveway density and median type, are critical for V–I outcomes, underscoring the role of access management and median design in roadside safety.
+Across both interaction types, the negative shape parameter indicates bounded extremes, reflecting physical and behavioral limits on crash risk escalation.
 6 Model validation
-The dataset for the study corridor provides precise geolocation for each extreme near-miss event and readily available annual corridor-level observed crashes.
-The dataset for the study corridor provides precise geolocation for each extreme near-miss event from Eqn.
-32.
-To address this limitation, validation was conducted as an event-level case–control approach, designed to assess whether the model assigns higher predicted probabilities to more severe near-miss events.
-Posterior exceedance probabilities were computed from Eqn.
-30 for severity thresholds ranging from –0.1 to –0.9 seconds, and these probabilities were then used to evaluate classification performance using the area under the receiver operating characteristic curve (ROC–AUC).
-This metric is particularly well-suited to rare-event prediction because it is independent of a specific threshold.
+The study corridor dataset provided precise geolocation for each extreme near-miss event (Eqn.
+32) along with annual corridor-wide crash totals.
+Although annual crash data offer a useful aggregate benchmark, they lack precise geolocation of individual crashes and therefore cannot be directly matched to near-miss predictions at the event level.
+To address this limitation, validation was instead designed as an event-level case–control exercise, assessing whether the model assigns higher exceedance probabilities to more severe near-miss events.
+Posterior exceedance probabilities were computed (Eqn.
+30) across severity thresholds from –0.1 to –0.9 seconds, and classification performance was evaluated using ROC–AUC, a threshold-independent metric well suited to rare-event prediction.
 (a) V–V Segment
@@
 Fig. 13.: ROC curves for four HBSGRP–UGEV models across severity thresholds from $\omega=-0.1$s to $\omega=-0.9$s
-ROC curves were generated for the four HBSGRP–UGEV variants, covering both V-V, V–I interactions at segment and intersection levels.
+ROC curves were generated for the four HBSGRP–UGEV variants, covering both V-V and V-I interactions at the segment and intersection levels.
 The results (Figure 6) reveal clear distinctions across near-miss types and roadway contexts.
-The segment-level V-V model consistently achieved the highest accuracy, with mean AUC values of about 0.89 and very little variation across thresholds.
-This indicates stable and robust predictive performance, reflecting the ability of grouped random parameters to capture high-speed longitudinal near-miss dynamics along uninterrupted midblock segments.
-The V-I model at the segment level achieved slightly lower accuracy, averaging around 0.79, and exhibited moderate sensitivity to the choice of severity threshold, a finding consistent with the greater variability in driver responses when vehicles approach roadside boundaries.
-At intersections, the V-V model demonstrated moderately strong predictive accuracy, with an average AUC of nearly 0.82 and relatively low volatility across thresholds.
-This indicates that grouped speed and spacing dynamics remain reliable even under more complex signalized and turning conditions.
-By contrast, the intersection-level V-I model performed weakest, averaging an AUC of approximately 0.75 and showing large fluctuations across thresholds, which reflects the complexity of boundary-related near-misses in intersections where turning trajectories, curb proximity, traffic signals, and pedestrian activity all interact.
-Taken together, these findings confirm that the HBSGRP–UGEV framework provides strong and reliable performance for V-V interactions, particularly in segment environments, while predictions for V-I interactions are more sensitive to threshold selection and local variability.
-This suggests that segment-level V-V estimates are the most operationally reliable for proactive safety monitoring, whereas intersection-level V-I near-misses may require further refinement.
-Incorporating more detailed information on curb geometry, pedestrian exposure, and high-resolution vehicle–infrastructure proximity measures could improve stability and predictive accuracy in future applications.
+The V–V segment model achieved the highest predictive accuracy (AUC $\approx$ 0.89), reflecting the influence of dynamic variability in driver behavior.
+Relative speed, deceleration, and distance emerged as significant random parameters, capturing heterogeneity in conflict severity across segments.
+At the same time, fixed effects such as heading difference, steering difference, jerk, lane width, and driveway density contributed additional explanatory power, indicating that both dynamic interactions and roadway design shape V–V near-miss risks.
+The V–I model at the segment level achieved moderately strong accuracy (AUC $\approx$ 0.79) but displayed greater sensitivity to severity thresholds.
+Only relative distance emerged as a significant random parameter, while other vehicle dynamics (e.g., speed, deceleration, etc.) and roadway features (e.g., lane width, driveway density, median type, and etc.) contributed as fixed effects.
+This suggests that V–I risks are shaped primarily by localized conditions, with dynamic variability playing a comparatively smaller role than in V–V interactions.
+At intersections, the V–V model achieved moderately strong accuracy (AUC $\approx$ 0.82) with low threshold sensitivity.
+Random parameters for relative speed and relative distance captured cross-location heterogeneity in approach and gap-acceptance conditions, while fixed effects— including deceleration/acceleration, steering and heading differences, jerk, and a left-turn indicator—explained lateral maneuvering and control effects.
+Together, these terms allow the model to represent both longitudinal and lateral conflict dynamics under signalized conditions, yielding stable performance.
+In contrast, the intersection-level V-I model performed weakest (AUC $\approx$ 0.75) and showing large fluctuations across thresholds, which reflects the complexity of boundary-related near-misses in intersections where turning trajectories, curb proximity, traffic signals, and pedestrian activity all interact.
+Taken together, the results demonstrate that the HBSGRP–UGEV framework yields robust and consistent predictive accuracy for V–V interactions, particularly along corridor segments, where conflict dynamics are more regular and less influenced by intersectional variability.
+By contrast, V–I models—especially at intersections—showed lower stability and greater sensitivity to threshold specification, reflecting the complexity of boundary-related near-miss shaped by turning flows, curb geometry, and multimodal activity.
+These findings suggest that segment-level V–V estimates are the most operationally reliable for proactive monitoring, while intersection-level V–I models may require further refinement.
+Incorporating richer representations of curb design, pedestrian exposure, and high-resolution V-I proximity measures could strengthen stability and predictive performance in future applications.
 7 Summary and conclusions
-This study addresses the persistent challenge of accurately estimating short-term crash occurrence risk (COR) in complex, mixed urban traffic environments, where conventional surrogate safety models often overlook detailed vehicle geometry, roadway context, and site-specific behavioral variability.
-Existing approaches are limited by oversimplified near-miss representations, a predominant focus on vehicle–vehicle (V–V) interactions, and insufficient integration of spatial heterogeneity in statistical modeling.
-To overcome these gaps, we propose a high-fidelity, geometry-aware, and statistically robust framework for real-time COR estimation that introduces a novel two-dimensional time-to-collision (2D-TTC) indicator.
-This metric accounts for vehicle dimensions, dynamic states, and precise geometric positioning, enabling accurate detection of extreme near-miss events from both V–V and vehicle–infrastructure (V–I) interactions.
-Extreme events are translated into probabilistic risk estimates through a non-stationary Univariate Generalized Extreme Value (UGEV) framework implemented within a Hierarchical Bayesian Spatial Grouped Random Parameter (HBSGRP) structure.
-A corridor-based grouping strategy was employed to address event sparsity, pooling near-miss samples across intersections and directional segments while preserving local heterogeneity.
-Execution of the framework involved extracting high-resolution vehicle trajectories from the Argoverse-2 motion forecasting dataset, integrating them with HD maps and supplementary roadway information, and identifying near-miss events.
-The model incorporated both vehicle dynamics covariate (relative speed, acceleration, deceleration, jerk, heading difference, steering difference, and traffic volume) and geometric or operational covariates (lane number, lane width, driveway density, median type, turning movement, and vehicle lane position).
+This study addresses the persistent challenge of estimating short-term crash occurrence risk (COR) in urban corridor-wide traffic environments, where conventional surrogate safety models often overlook detailed vehicle geometry, roadway context, and site-specific behavioral variability.
+Existing approaches have typically focused on vehicle–vehicle (V–V) interactions and neglected vehicle–infrastructure (V–I) conflicts, limiting their applicability at the corridor scale.
+To overcome these gaps, we introduced a geometry-aware two-dimensional time-to-collision (2D-TTC) indicator that incorporates vehicle dimensions, dynamic states, and roadway boundaries, enabling the detection of extreme near-miss events from both V–V and V–I interactions.
+Extreme events were modeled through a nonstationary Univariate Generalized Extreme Value (UGEV) distribution embedded within a Hierarchical Bayesian Spatial Grouped Random Parameter (HBSGRP) framework.
+This structure allowed partial pooling across intersections and directional segments while preserving local heterogeneity, addressing data sparsity and contextual variation.
+High-resolution vehicle trajectories from the Argoverse-2 dataset were integrated with HD maps to extract near-miss events.
+Vehicle dynamics (relative speed, acceleration, deceleration, jerk, heading and steering differences, and traffic volume) and roadway features (lane width, driveway density, median type, turning movements, and lane position) were incorporated as covariates, with grouped random parameters specified for key dynamic variables.
 Grouped random parameters were specified for temporally varying dynamics, relative speed, relative deceleration, and relative distance, to capture their context-dependent effects.
 For benchmarking, a baseline hierarchical Bayesian fixed-parameter (HBSFP) model was also estimated.
-Results demonstrate that HBSGRP consistently outperformed the HBSFP across both intersections and segments, achieving lower DIC, WAIC, and LOOIC values and providing greater estimation stability.
-Relative speed emerged as the dominant predictor of near-miss severity in intersections, while relative distance was more influential for segment-level V–V events.
-Deceleration played a critical role in segment-level V–V near-miss, underscoring the importance of longitudinal control.
+Results showed that the HBSGRP model consistently outperformed its fixed-parameter counterpart, yielding lower DIC, WAIC, and LOOIC values and demonstrating greater stability.
+Random-parameter vehicle dynamics revealed that relative speed, deceleration, and distance dominated V–V near misses at the segment level, whereas only relative distance was significant for V–I near misses.
+At intersections, V–V near misses were dominated by relative speed and distance, while no dynamic variables were significant for V–I near misses.
 Fixed covariates such as heading difference, steering difference, lane width, driveway density, and median type also showed significant effects, demonstrating the ability of the model to capture both dynamic behaviors and roadway design influences.
 Validation results further confirmed model credibility.
-ROC–AUC analysis showed strong predictive performance, with V–V segment models achieving the highest accuracy (AUC $\approx$ 0.89), followed by intersection V–V ($\approx$ 0.82).
-V–I models performed reasonably well (segment $\approx$ 0.79; intersection $\approx$ 0.75) but displayed greater sensitivity to severity threshold choice.
-These findings highlight that while V–V interactions are modeled with high stability, V–I near-misses especially at intersections, remain more complex and variable.
-Despite its contributions, the study has several limitations.
-First, potential bias may arise from the dependence on AV onboard sensors, which may affect traffic state accuracy and near-miss detection.
-Second, the 2D-TTC formulation, while tractable, simplifies nonlinear vehicle dynamics, which could understate risks in extreme evasive maneuvers.
-Third, the current framework does not explicitly model near-miss severity levels.
-Future research should extend the framework to multivariate UGEV models incorporating complementary surrogate measures such as post-collision velocity change ($\Delta V$) and explicitly address vulnerable road user (VRU) near-misses.
-Broader validation across multiple cities, transferability calibration, and comparison with peaks-over-threshold (POT) GPD methods will further strengthen robustness and generalizability.
-As autonomous vehicle deployment increases, access to larger and richer trajectory datasets will enhance the statistical power of extreme value analysis and enable more nuanced safety assessments.
-In practice, the proposed HBSGRP–UGEV framework offers a scalable, data-driven tool for proactive safety monitoring at the corridor level.
-By enabling early identification of high-risk conditions and supporting real-time traffic management, it provides a pathway toward Vision Zero objectives, reducing fatalities and serious injuries while promoting safe, efficient, and equitable mobility in urban environments.
+ROC–AUC validation confirmed robust predictive accuracy, with V–V segment models performing best (AUC $\approx$ 0.89), followed by intersection V–V ($\approx$ 0.82)), while V–I models showed moderate performance ($\approx$ 0.79 for segments, $\approx$ 0.75 for intersections).
+These findings indicate that while V–V risks can be reliably modeled, V–I interactions, particularly at intersections, remain more variable and complex.
+While the proposed framework advances corridor-wide COR modeling, several aspects warrant further development.
+First, reliance on AV onboard sensors may introduce measurement bias in traffic state estimation and near-miss detection.
+Second, the 2D-TTC formulation simplifies nonlinear vehicle dynamics, which could understate risks during extreme evasive maneuvers.
+Third, the framework does not yet capture near-miss severity stratification.
+Future research offers opportunities to enrich this framework.
+Extending it to multivariate HBSGRP-UGEV models with complementary surrogate measures such as post-collision velocity change ($\Delta V$) would provide deeper insights into crash severity dynamics.
+Incorporating vulnerable road user (VRU) near-misses, validating across diverse urban networks, calibrating for transferability, and benchmarking against GPD methods will further enhance robustness and generalizability.
+As AV deployment expands, increasingly rich trajectory datasets will strengthen the statistical power of HBSGRP–UGEV and open new opportunities for nuanced, multimodal safety assessments.
+Within this evolving data landscape, the proposed framework provides a scalable and adaptable foundation for proactive corridor-wide safety monitoring.
+By enabling early identification of high-risk conditions and supporting real-time traffic management, the framework aligns with the broader Vision Zero agenda, advancing the long-term goal of safer, more efficient, and equitable urban mobility.
 Acknowledgement:

### v4

@@
 Abstract
-Real-time corridor-wide crash occurrence risk (COR) prediction is challenging because existing near-miss EVT models oversimplify collision geometry, neglect vehicle–infrastructure (V–I) interactions, and fail to capture spatial heterogeneity in traffic and roadway conditions.
-To address these gaps, this study introduces a geometry-aware two-dimensional time-to-collision (2D-TTC) indicator for capturing V–V and V–I near misses, and a Hierarchical Bayesian spatiotemporal grouped random parameter (HBSGRP) UGEV model for estimating short-term COR in urban corridors.
+Real-time, corridor-wide prediction of crash-occurrence risk (COR) is challenging because existing near-miss EVT models oversimplify collision geometry, neglect vehicle–infrastructure (V–I) interactions, and fail to adequately account for spatial heterogeneity in traffic and roadway conditions.
+To address these gaps, this study develops a geometry-aware 2D-TTC near-miss extraction and integrates it with a hierarchical Bayesian structure grouped random parameters (HBSGRP–UGEV) to estimate short-term COR in urban corridors.
+Building on prior grouped EVT formulations while explicitly accommodating both V–V and V–I near-miss processes within a unified corridor-wide modeling framework.
 High-resolution trajectories from the Argoverse-2 dataset were analyzed across 28 sites on Miami’s Biscayne Boulevard to extract extreme near-miss events.
 The model incorporates vehicle dynamics and roadway features as covariates, with partial pooling across segments and intersections to capture corridor-wide heterogeneity.
-Results show that the HBSGRP–UGEV framework outperforms fixed-parameter models, reducing DIC by up to 7.5% for V–V and 3.1% for V–I near misses.
+Results show that the HBSGRP–UGEV framework outperforms fixed-parameter models, reducing DIC by up to 7.5% (V–V) and 3.1% (V–I).
 Predictive validation using ROC–AUC confirms strong accuracy (0.89 for V–V segments, 0.82 for intersections, 0.79 for V–I segments, and 0.75 for intersections).
-Random-parameter effects showed that relative speed, distance, and deceleration dominated V–V near misses at segments (only distance for V–I), while intersections were dominated by relative speed and distance for V–V, with no significant effects for V–I..
+Grouped random-parameter results indicate that relative (speed, distance, and deceleration) dominate V–V near-miss risk on segments, whereas V–I segment risk is primarily associated with relative distance; at intersections, V–V risk is driven by relative (speed and distance), while V–I dynamics exhibit no statistically significant effects.
 These findings demonstrate the value of a geometry-aware, spatially adaptive framework for proactive corridor safety management, supporting both real-time interventions and long-term Vision Zero goals.
@@
 1 Introduction
-The increasing complexity of urban corridors, rising traffic volumes, and the coexistence of human-driven vehicles (HDVs) and autonomous vehicles (AVs) demand innovative safety assessment approaches.
+The increasing complexity of urban corridors, rising traffic volumes, and the coexistence of human-driven vehicles (HDVs) and autonomous vehicles (AVs) require innovative safety assessment approaches.
 Urban corridors are particularly vulnerable, as even a single crash can disrupt mobility, trigger cascading congestion, and result in substantial economic losses.
-In 2023, traffic crashes in the United States claimed 40901 lives, with a fatality rate of 1.26 per 100 million vehicle miles traveled (VMT) (National Center for Statistics and Analysis, 2025).
+In 2023, traffic crashes in the US claimed 40901 lives, with a fatality rate of 1.26 per 100 million vehicle miles traveled (VMT) (National Center for Statistics and Analysis, 2025).
 The total annual cost of crashes exceeds \$1.85 trillion, including \$460 billion in direct economic losses and \$1.4 trillion in reduced quality of life (Blincoe et al., 2023).
@@
 Crash risk prediction has traditionally relied on retrospective crash data (Arun et al., 2021; Lee et al., 2017; Pei et al., 2011).
-While such approaches have advanced understanding of crash patterns, they face well-documented limitations (Lord and Mannering, 2010; Lord et al., 2021; Mannering et al., 2020), including the rarity and stochastic nature of crashes, underreporting (Arun et al., 2021; Tarko, 2018), temporal delays, and spatial mismatches with real-time traffic conditions.
+Although such approaches have advanced understanding of crash patterns, they face well-documented limitations (Lord and Mannering, 2010; Lord et al., 2021; Mannering et al., 2020), including the rarity and stochastic nature of crashes, underreporting (Arun et al., 2021; Tarko, 2018), temporal delays, and spatial mismatches with real-time traffic conditions.
 These drawbacks highlight the need for proactive approaches that anticipate rather than react to crash risk.
 To address these challenges, researchers have shifted toward proactive trajectory-based surrogate safety measures (SSMs) (Ali et al., 2023a; Arun et al., 2021; Mahmud et al., 2017).
-SSMs use frequent near-miss events as proxies for crashes, enabling the study of traffic conflict dynamics at a much higher temporal resolution (Li et al., 2024; Tarko, 2018; Davis et al., 2011).
+SSMs use frequent near-miss events as proxies for crashes, enabling the study of traffic conflict dynamics at a much higher temporal resolution (e.g., Li et al., 2024; Tarko, 2018; Davis et al., 2011).
 Among these, time-based indicators such as Time-to-Collision (TTC), Post-Encroachment Time (PET), and Modified TTC (MTTC) are widely used (Hayward, 1971; Perkins and Harris, (1967; Minderhoud and Bovy, 2001; Ozbay et al., 2008; Allen et al., 1978; Venthuruthiyil and Chunchu, 2022).
-However, their assumptions of constant velocity and linear paths bias results in boundary-constrained settings.
-Recent two-dimensional (2D) extensions (Li et al., 2024; Anis et al., 2025b) incorporate position, heading, and acceleration, yet still rely on simplified circular vehicle approximations.
-This study addresses these limitations by introducing a geometry-aware 2D-TTC that accounts for vehicle dimensions, dynamics, and roadway boundaries to realistically capture both vehicle–vehicle (V–V) and vehicle–infrastructure (V–I) near misses.
-Building on this foundation, recent studies have integrated SSMs with extreme value theory (EVT) to estimate crash occurrence risk (COR) in urban corridors (Lanzaro et al., 2023; Kamel et al., 2023, 2024; Kamel and Sayed, 2024; Singh et al., 2024; Ghoul and Sayed, 2025; Singh et al., 2025; Anis et al., 2025b).
-EVT extrapolates from frequent near misses to rare crash-prone extremes, offering a statistically principled framework for proactive safety analysis.
-Yet most SSM–EVT applications remain limited, focusing narrowly on V–V interactions and excluding V–I conflicts such as lane edges, curbs, and barriers.
+However, their assumptions of constant velocity and linear paths can bias near-miss characterization, especially in turning and boundary-constrained settings.
+Recent two-dimensional (2D) extensions of TTC (Li et al., 2024; Anis et al., 2025b) that incorporate position, heading, yaw rate, speed, and acceleration often still rely on simplified vehicle shape approximations (e.g., circular).
+Although Hou et al.
+2014 compared alternative vehicle-shape representations (e.g., circle, rectangle, and hybrid forms) for simplified TTC computation in simulation, a corridor-scale, high-fidelity 2D-TTC that propagates geometry-aware footprints and jointly extracts V–V and V–I near-misses from trajectories remains limited.
+Recent studies have integrated computed SSMs with the extreme value theory (EVT) framework to estimate crash occurrence risk (COR) in urban corridors (Lanzaro et al., 2023; Kamel et al., 2023, 2024; Kamel and Sayed, 2024; Singh et al., 2024; Ghoul and Sayed, 2025; Singh et al., 2025; Anis et al., 2025b).
+Since EVT can extrapolate from frequent near misses or SSMs to rare crash-prone extremes, it offers a statistically principled framework for proactive safety analysis.
+Yet most near-miss and EVT frameworks remain limited, focusing narrowly on V–V interactions and excluding V–I near-misses such as lane edges, curbs, and barriers.
 This omission is critical because V–I events account for a large share of severe roadway departure crashes (Islam et al., 2021).
@@
 The BM approach is particularly well-suited to real-time applications due to its short block structure, reduced threshold sensitivity, and compatibility with discrete-time samples (Fu and Sayed, 2022a).
-Early applications of EVT in traffic safety used stationary GEV models (Wang et al., 2018; Guo et al., 2019; Orsini et al., 2019; Alozi and Hussein, 2022; Hussain et al., 2022), which assumed constant parameters across study areas.
-More recent work has developed nonstationary GEV formulations, allowing covariates to influence distribution parameters (Tahir and Haque, 2024; Ali et al., 2022; Fu and Sayed, 2022a; Zheng et al., 2019; Nazir et al., 2023; Kar et al., 2024; Ali et al., 2023b; Alozi and Hussein, 2022; Zheng et al., 2014; Kar et al., 2023) and capture heterogeneity across traffic states.
+Early applications of EVT in traffic safety used stationary GEV models (Wang et al., 2018; Guo et al., 2019; Orsini et al., 2019; Alozi and Hussein, 2022; Hussain et al., 2022), which assumed constant parameters across the study area.
+Over time, nonstationary GEV formulations have been developed, allowing covariates to influence distribution parameters (Tahir and Haque, 2024; Ali et al., 2022b; Fu and Sayed, 2022a; Zheng et al., 2019; Nazir et al., 2023; Kar et al., 2024; Ali et al., 2023b; Alozi and Hussein, 2022; Zheng et al., 2014; Kar et al., 2023) and thereby represent variation across traffic states and locations.
 (Lanzaro et al., 2023; Kamel et al., 2023, 2024; Kamel and Sayed, 2024; Singh et al., 2024; Ghoul and Sayed, 2025; Singh et al., 2025; Anis et al., 2025b).
 Bayesian hierarchical extensions further address data sparsity and unobserved heterogeneity (Zheng and Sayed, 2019; Fu and Sayed, 2022b).
-Despite these advances, most applications remain geographically narrow and functionally restricted to either intersections (Fu and Sayed, 2023, 2021c, 2022b) or segments (Kumar and Mudgal, 2024; Kamel et al., 2023; Kamel and Sayed, 2024; Anis et al., 2025b; Kamel et al., 2024).
-Although some studies have integrated vehicle dynamics (Kumar and Mudgal, 2024; Fu and Sayed, 2022b; Anis et al., 2025b) and examined heterogeneity across traffic flow directions or roadway geometries (Kamel et al., 2023, 2024), explicit corridor-wide models that jointly integrate vehicle dynamics, roadway geometry, and directional heterogeneity remain rare.
-To address this gap, this paper makes three contributions: (1) it develops a corridor-wide COR framework that incorporates both V–V and V–I interactions, (2) it introduces a geometry-aware 2D-TTC indicator that captures extreme near-miss dynamics with higher geometric fidelity, and (3) it advances a hierarchical Bayesian spatiotemporal grouped random parameter (HBSGRP) GEV model that integrates roadway geometry, traffic directionality, and multilevel heterogeneity for corridor-wide yet location-sensitive crash risk estimation.
-Together, these contributions establish the first infrastructure-aware, statistically rigorous framework for proactive corridor-wide crash risk prediction, leveraging Argoverse-2 trajectory and HD map data to support both real-time interventions and long-term safety planning.
+A growing body of EVT-based safety studies now incorporates vehicle dynamics and roadway context (Fu and Sayed, 2022b; Kumar and Mudgal, 2024; Anis et al., 2025b; Singh et al., 2025) and, in some cases, distinguishes traffic-flow directions (Kamel et al., 2024).
+However, direction-wise parametrization of vehicle-dynamics effects remains limited, particularly for corridor segments where interaction mechanisms can differ systematically by travel direction and access density.
+Moreover, most EVT applications focus on intersection settings (Fu and Sayed, 2023, 2021c, 2022b; Ghoul et al., 2023a; Kumar and Mudgal, 2024), only a small subset extends to corridor or network-scale inference (Ghoul et al., 2023b; Kamel et al., 2023, 2024; Kamel and Sayed, 2024; Singh et al., 2024; Anis et al., 2025b; Singh et al., 2025).
+To address these gaps, this paper makes two contributions.
+First, it develops a corridor-wide COR framework that incorporates both V–V and V–I interactions by implementing a geometry-aware 2D-TTC formulation that forward-projects rectangular vehicle footprints and uses HD-map boundaries to define V–I exposure, enabling joint extraction of V–V and V–I extremes from AV trajectories.
+Second, it adapts and extends a hierarchical Bayesian structure grouped random-parameters (HBSGRP) UGEV framework for corridor-scale modeling, integrating fixed roadway and time-varying vehicle-dynamics covariates within an interaction-window BM framework to capture multilevel heterogeneity across intersections and directional segments.
+Together, these contributions establish an infrastructure-aware and spatially adaptive framework for proactive corridor-wide crash-risk estimation that links high-fidelity near-miss geometry with hierarchical EVT inference, supporting both real-time intervention and longer-term safety planning.
 The remainder of this paper introduces the proposed 2D-TTC indicators, the HBSGRP–UGEV modeling framework, and the data preparation used in the analysis.
@@
 2 Near-miss detection framework
-This section presents the modeling framework for computing a geometry-aware, high-fidelity 2D-TTC to support COR estimation.
-The framework integrates a continuous-time kinematic bicycle model, vehicle geometry representation, and a spatial near-miss detection algorithm, allowing capture of V-V and V-I interactions.
-Together, these components enable the accurate forecasting of vehicle trajectories and identification of potential near misses.
-The resulting architecture forms the computational backbone of the SSM analysis, ensuring a realistic representation of near-miss events in mixed-traffic environments.
+This section presents the computational framework for estimating a geometry-aware, high-fidelity 2D-TTC to support real-time COR inference.
+The proposed workflow follows projection-based TTC concepts that forward-simulate oriented rectangular footprints, but is implemented directly on high-frequency AV trajectories and HD-map boundary primitives.
+The framework integrates a continuous-time kinematic bicycle model, explicit vehicle footprint geometry, and a spatial proximity algorithm to capture both V–V and V–I interactions.
+Vehicles are represented as oriented bounding boxes, and V–I exposure is defined using the curb/median boundaries of the vehicle’s current carriageway extracted from the HD map, enabling consistent joint detection of near misses along the corridor.
+Together, these components provide a deterministic, high-resolution basis for extracting extreme near-miss severity used in the subsequent EVT-based COR model.
 2.1 Vehicle dynamics model
 The vehicle dynamics module provides the predictive foundation of the framework.
-This adopts the kinematic bicycle model to represent vehicle footprint, as it is suitable for short-term near-miss analysis, where steering and acceleration dominate, and lateral slip can be neglected.
+Which adopts the kinematic bicycle model for short-horizon interaction forecasting, where steering and acceleration dominate, and lateral tire slip can be neglected.
 The motion is represented by a four-dimensional (4D) dynamic state vector.
-At any time $t$, the state of vehicle $i$ is given by:
-$\mathbf{x}_{i}(t)=\begin{bmatrix}x_{i}(t)\\[2.84526pt] y_{i}(t)\\[2.84526pt] \theta_{i}(t)\\[2.84526pt] v_{i}(t)\end{bmatrix}$ (1)
+At any time $t$, the state of vehicle $i$ is given by Eq.
+(1):
+$\mathbf{x}_{i}(t)=\begin{bmatrix}x_{i}(t)\\[2.84526pt] y_{i}(t)\\[2.84526pt] v_{i}(t)\end{bmatrix}$ (1)
 Here, $x_{i}(t),y_{i}(t)$ denote the vehicle’s global center coordinates, $\theta_{i}(t)$ its heading angle, and $v_{i}(t)$ its speed.
 Over the prediction window, each vehicle is assumed to maintain constant acceleration $a_{i}$ and steering angle $\delta_{i}$, simplifying short-horizon trajectory integration.
-The motion then follows the standard kinematic bicycle equations:
+The motion then follows the standard kinematic bicycle Eqs.
+(2-5):
 $\dot{x}_{i}(t)$ $=v_{i}(t)\cos\theta_{i}(t)$ (2)
@@
 The steering angle $\delta_{i}$ and acceleration $a_{i}$ are treated as fixed inputs over the lookahead period.
-Equations (2-5) describe the rates of change in position, orientation, and velocity, determining the vehicle’s trajectory.
+Eqs.
+(2-5) describe the rates of change in position, orientation, and velocity, determining the vehicle’s trajectory.
 To model interactions, individual states are concatenated into an eight-dimensional (8D) joint state vector, defined as:
@@
 $\dot{X}(t)=F\bigl(X(t)\bigr)=\begin{bmatrix}v_{A}\cos\theta_{A}&v_{A}\sin\theta_{A}&\frac{v_{A}}{L_{A}}\tan\delta_{A}&a_{A}&v_{B}\cos\theta_{B}&v_{B}\sin\theta_{B}&\frac{v_{B}}{L_{B}}\tan\delta_{B}&a_{B}\end{bmatrix}^{\top}$ (7)
-The state-space vector field in Eqn.7 gives the instantaneous rate of change of the joint state $\dot{X}(t)$.
+The state-space vector field in Eq.
+(7) gives the instantaneous rate of change of the joint state $\dot{X}(t)$.
 Its forward evolution is computed using the fourth-order Runge–Kutta (RK4) method, which balances accuracy and efficiency for nonlinear vehicle dynamics.
@@
 This integration is repeated for each timestep $n\in\{0,\dots,N-1\}$, producing a sequence of predicted states $\{{X}_{n}\}$ over the horizon.
-These trajectories form the basis for geometric collision checks between vehicles and roadway boundaries.
+These predicted states are then converted to oriented footprints for geometric proximity checks against other vehicles and HD-map boundaries.
 2.2 Vehicle geometry representation
@@
 Repeating it for all corners at each timestep produces the vehicle’s time-varying geometric footprint, which serves as input for near-miss detection.
-These footprints enable precise evaluation of proximity-based events, both V-V and V-I, ensuring accurate 2D-TTC estimation while reducing false positives from geometric simplifications.
+This footprint construction is applied at every RK4 step, yielding a time-varying oriented bounding box suitable for both V–V and V–I proximity evaluation.
 2.3 V-V near-miss detection
-Vehicles are modeled as rigid rectangles, and their interactions are determined by predicted states and orientations.
-At each timestep of the RK4 integration, the global positions of all four corners are computed using the kinematic and rotational transformations.
-A collision is defined when a vehicle’s corner lies within a proximity threshold of a corner of another vehicle.
+Vehicles are modeled as rigid rectangles, and predicted states and orientations determine their interactions.
+At each RK4 integration timestep, the global positions of all four corners are computed using the kinematic and rotational transformations.
+A V–V near-miss is detected when the predicted footprints exhibit near-contact under a numerical tolerance $\epsilon$, implemented through corner-based proximity checks across all corner pairs.
 Let ${p}_{A}^{(j)}(t)$ and ${p}_{B}^{(k)}(t)$ denote the global coordinates of corners $j$ and $k$ of vehicles A and B at time $t$.
@@
 ${p}_{A}^{(j)}(t)={p}_{B}^{(k)}(t)$ (13)
-An exact match between corners of two vehicles is highly unlikely due to floating-point arithmetic and continuous motion.
+An exact match between the corners of two vehicles is highly improbable, given floating-point arithmetic and continuous motion.
 Instead, a practical criterion is applied: a collision is flagged if the difference in both the x and y coordinates of any corner pair is less than or equal to a small threshold $\epsilon$:
@@
 A V–V near-miss is flagged if any pair satisfies the proximity condition, and the corresponding collision time $t_{c}$ is recorded for 2D-TTC computation.
-This discrete-time approach enables accurate detection of spatial overlap while avoiding the oversimplifications of circular vehicle approximations.
+This discrete-time footprint-based check aligns with high-frequency trajectory sampling and avoids the geometric bias introduced by circular approximations in oblique and turning interactions.
 2.4 V-I near-miss detection
-In addition to V-V interactions, the framework also detects near misses between vehicles and roadway infrastructure, including curbs, medians, or barriers.
+In addition to V-V interactions, the framework also detects near-misses between vehicles and roadway infrastructure, including curbs, medians, and barriers.
 Capturing these V–I events is critical for identifying edge encroachments or departures that often precede run-off-road crashes.
-Road boundaries are represented as static polylines of $M$ discrete points $(X_{\ell},Y_{\ell})$ consistent with HD map formats in datasets such as Argoverse-2.
+In this study, the scope is defined as the curb and median boundaries of the vehicle’s current carriageway as encoded in the HD map; opposing-direction curbs, minor side-street sidewalks, and internal lane markings are beyond the scope.
+To do so, curb and boundaries are represented as static polylines of $M$ discrete points $(X_{\ell},Y_{\ell})$ consistent with HD map formats in datasets such as Argoverse-2.
 At each timestep, the global positions of the four vehicle corners ${p}_{i}^{(j)}(t)$ are obtained from the predicted state.
@@
 These events are then passed to the 2D-TTC framework to support extreme risk estimation.
-2.5 Spatiotemporal COR
+2.5 COR estimation
 The near-miss detection framework integrates vehicle dynamics, geometric transformations, and spatial proximity checks to identify both V–V and V–I events.
@@
 Fig. 1.: Computation of 2D-TTC for near-miss scenarios
-While this geometric framework enables precise localization of near misses, it does not capture the probabilistic structure or temporal evolution of COR.
-To address this, the extracted blockwise minima of 2D-TTC are modeled using the EVT model.
-For each block, the model estimates the probability that 2D-TTC falls below a critical safety threshold, which defines the COR metric as a short-term indicator of crash risk.
-Block duration is chosen to balance temporal sensitivity with statistical stability across diverse contexts.
-This framework, therefore, combines deterministic near-miss detection with probabilistic EVT modeling to enable localized, data-driven assessment of crash-prone conditions.
-The complete methodological pipeline, from trajectory simulation to dynamic risk estimation, is summarized in Fig.
+While the above procedure deterministically detects and localizes near-misses, COR requires a probabilistic mapping from near-miss severity to short-term crash risk.
+Accordingly, BM, defined as the minimum 2D-TTC observed across the selected window, is modelled using EVT.
+COR is then defined as the estimated probability that 2D-TTC falls below a pre-specified critical threshold.
+Here, this block definition preserves short-lived maneuver dynamics at the native trajectory resolution while yielding a sufficient set of approximately independent extremes for stable inference across locations.
+The end-to-end pipeline from trajectory simulation to COR estimation is summarized in Fig.
 2.5.
@@
 In this study, extreme events are represented by block maxima of 2D-TTC derived from high-resolution vehicle trajectories (Algorithm 1).
-EVT allows risk estimation without requiring large volumes of historical crash data (Zheng and Sayed, 2020; Fu and Sayed, 2022a).
+EVT enables risk estimation without requiring large volumes of historical crash data (e.g., Zheng and Sayed, 2020; Fu and Sayed, 2022a).
 Two EVT approaches are commonly applied: the Block Maxima (BM) method, which fits maxima or minima from fixed-duration blocks using the Generalized Extreme Value (GEV) distribution (Coles et al., 2001), and the Peaks Over Threshold (POT) method, which models threshold exceedances using the Generalized Pareto Distribution (GPD).
-Although POT can be data-efficient, it is highly sensitive to threshold choice and less compatible with discrete sensor-derived datasets.
-By contrast, BM aligns naturally with real-time AV data, avoids subjective thresholds, and supports scalable, time-series analysis (Fu and Sayed, 2022a; Songchitruksa and Tarko, 2006).
+Although POT can be data-efficient, it is highly sensitive to threshold selection and typically requires declustering and handling of dependencies when applied to dense trajectory data.
+By contrast, BM aligns naturally with real-time AV unique-segmented data, avoids subjective thresholds, and supports scalable time-series analysis (Fu and Sayed, 2022a; Songchitruksa and Tarko, 2006).
 Consequently, this study adopts the BM–GEV framework (see Fig.
@@
 The GEV has been widely applied in traffic safety research.
-Early studies modeled near-missets with univariate and multivariate GEV formulations (Orsini et al., 2020; Zhang et al., 2019; Zheng et al., 2014).
+Early studies modeled near-misses using univariate and multivariate GEV formulations (e.g., citeporsini2020large, zhang2019real, zheng2014freeway).
 Recent advances incorporate nonstationary structures, linking GEV parameters to covariates such as vehicle dynamics (Anis et al., 2025b; Fu and Sayed, 2021c, 2022b).
 Bayesian hierarchical models have also addressed sparse data and unobserved heterogeneity across vehicles, sites, and traffic contexts.
-The increasing availability of high-resolution AV trajectory and HD map data has highlighted the need for flexible GEV formulations that jointly capture V–V and V–I interactions.
-Despite this progress, most UGEV models still neglect the joint influence of roadway geometry and vehicle dynamics, two critical but distinct sources of heterogeneity.
-Roadway geometry (e.g., lane count, curvature, median type) is spatially structured and relatively stable within segments or intersections, while vehicle dynamics (e.g., speed, acceleration) vary temporally and spatially, driving short-term behavioral volatility.
-Capturing this multilevel heterogeneity is crucial for estimating corridor and network level crash risk.
-To address this gap, this study develops a hierarchical Bayesian Spatiotemporal Grouped Random Parameter (HBSGRP) UGEV model.
-Unlike standard random-parameter models, which assign variation independently to each unit, grouped random parameters capture shared variation across clusters (e.g., segment types, corridor directions) while allowing for within-group variability.
+With increasing availability of AV trajectories and HD maps, there is growing motivation to combine geometry-aware near-miss extraction (including V–I exposure) with hierarchical EVT to support corridor-scale risk inference.
+Despite this progress, most UGEV models still neglect the joint influence of fixed and time-variant parameters, which are critical but distinct sources of heterogeneity.
+For instance, (e.g., lane count, median type) is relatively stable within segments or intersections, whereas vehicle dynamics (e.g., speed, deceleration, spacing) vary within and across interaction windows and drive short-term volatility.
+Capturing this heterogeneity is crucial for estimating corridor or network-level crash risk.
+To address this gap, this study employs a hierarchical Bayesian structure grouped random parameters (HBSGRP) UGEV model.
+Unlike standard random-parameter models, which assign variation independently to each unit, grouped random parameters capture shared variation across clusters (e.g., segments, intersections) while allowing for within-group variability.
 This structure provides both flexibility and parsimony, making it well-suited for high-resolution, corridor-scale crash risk modeling (Washington et al., 2020; Cai et al., 2018; Islam et al., 2023).
 3.1 Univariate generalized extreme value (UGEV)
-As established in the previous section, the BM-based UGEV model is employed to characterize the tail behavior of 2D-TTC, which represents the most critical near-miss events in both V–V and V–I interactions.
-To implement this framework, continuous vehicle trajectories are partitioned into fixed-duration windows (e.g., 11 s).
-Within each window, the minimum 2D-TTC is extracted and negated, consistent with the upper-tail modeling convention of UGEV (Coles et al., 2001).
-The resulting sequence of BM serves as the input for parameter estimation.
-Throughout this study, we focus on maxima, denoted as $M_{n}$, under the assumption that each block consists of $n$ independent and identically distributed (i.i.d.) random variables, $X_{1},X_{2},\ldots,X_{n}$.
+The BM-based UGEV model is used to characterize the tail behavior of extreme (e.g., 2D-TTC), which represents the most critical events (near-miss: V–V and V–I interactions).
+To implement this framework, continuous vehicle trajectories are separated into fixed-duration windows.
+Within each window, the minimum 2D-TTC is extracted and negated to align with the upper-tail modeling convention (Coles et al., 2001).
+The resulting BM sequence serves as input for parameter estimation.
+Throughout this study, we focus on maxima, denoted by $M_{n}$, under the assumption that each block consists of $n$ independent and identically distributed (i.i.d., $X_{1},X_{2},\ldots,X_{n}$).
 To align with conventional UGEV formulations, we consider the negative values of the original variables derived from Algorithm 1, thereby defining the block maximum as $M_{n}=\max\{-X_{1},-X_{2},\ldots,-X_{n}\}$.
-Under suitable regularity conditions, if there exist sequences of constants $a_{n}\in\mathbb{R}$ and $b_{n}>0$, the normalized block maxima $M_{n}^{*}=(M_{n}-a_{n})/b_{n}$ converge in distribution to a non-degenerate limit $G(x)$ as $n\to\infty$, then: $\Pr\left(\frac{M_{n}-a_{n}}{b_{n}}\leq X\right)\xrightarrow[n\to\infty]{}f(x)$, where $G(x)$ is a non-degenerate distribution function.
+Under suitable regularity conditions, if there exist sequences of constants $a_{n}\in\mathbb{R}$ and $b_{n}>0$, the normalized block maxima $M_{n}^{*}=(M_{n}-b_{n})/a_{n}$ converge in distribution to a non-degenerate limit $G(x)$ as $n\to\infty$, then: $\Pr\left(\frac{M_{n}-a_{n}}{b_{n}}\leq X\right)\xrightarrow[n\to\infty]{}f(x)$, where $G(x)$ is a non-degenerate distribution function.
 According to EVT, the limiting distribution $G(x)$ must belong to one of three families: Gumbel, Fréchet, or Weibull, each of which can be captured by the parametric form of the GEV distribution (Coles et al., 2001).
@@
 The value of $\xi_{x}$ determines tail heaviness: $\xi_{x}>0$ corresponds to the heavy-tailed Fréchet, $\xi_{x}<0$ to the bounded Weibull, and $\xi_{x}=0$ to the light-tailed Gumbel case.
-3.1.1 Hierarchical Bayesian Spatiotemporal Grouped Random Parameter (HBSGRP) model
-To capture structured heterogeneity in extreme outcomes across the corridor, we employ a Hierarchical Bayesian Spatiotemporal Grouped Random Parameter (HBSGRP) model.
-Unlike conventional random-parameter models that assign variation to individual units, the HBSGRP formulation introduces variation at the group level.
-This structure allows systematic differences across predefined observational groups while maintaining parsimony and interpretability.
-In this specification, covariates are divided into two sets: fixed effects $X$, which apply uniformly across all observations, and group-varying effects $Z$, whose coefficients differ across groups.
-For a dataset divided into $K$ groups, with observations $i$ belonging to group $k(i)$, both the intercept and selected slopes are allowed to vary by group.
-This enables the capture of localized relationships while preserving overall model coherence.
-The model is expressed within a three-layer hierarchical Bayesian framework.
-The data layer specifies the likelihood of observed block maxima of 2D-TTC from V–V and V–I interactions, assuming a UGEV distribution.
-The process layer links covariates to UGEV parameters through fixed and group-level random effects, with optional latent terms (e.g., Gaussian processes) to capture residual spatial–temporal dependence.
-The prior layer defines prior distributions for all parameters and hyperparameters, supporting regularization and uncertainty quantification.
+3.1.1 Hierarchical Bayesian Structure Grouped Random Parameters (HBSGRP) model
+To capture structured heterogeneity in extreme outcomes across the corridor, we employ the HBSGRP model.
+In this study, groups $k=1,\dots,K$ correspond to corridor elements (intersections or directional segments), so each block extreme $X_{k,i}$ is associated with a specific spatial context.
+Covariates are divided into (i) group-level fixed effects $Y$, which describe relatively stable roadway context (e.g., lane configuration, median type), and (ii) interaction-level dynamics $W$, which vary across blocks (e.g., relative speed, relative distance).
+The key feature of HBSGRP is that selected dynamic effects are allowed to vary by group, enabling systematic differences in interactions across corridor elements while maintaining corridor-wide coherence through partial pooling.
+The model is formulated as a three-layer hierarchical Bayesian structure: a data layer that models block extremes using a UGEV distribution, a process layer that introduces grouped random parameters (partial pooling across corridor elements), and a prior layer that regularizes the model and quantifies uncertainty.
+Under a latent Gaussian formulation, the UGEV parameters are linked to covariates through identity link functions that combine global fixed effects with group-specific random effects, allowing the model to represent corridor-wide structure while preserving localized heterogeneity.
 Bayesian inference proceeds by updating prior beliefs with observed data to obtain the posterior distribution. Formally, the joint posterior of parameters $\Theta$, given data $X$, is:
@@
 Lastly, $q_{\text{prior}}(\Psi)$ encompasses the prior distributions of hyperparameters that control the model’s flexibility and penalize over-complexity.
+Eqs.
+(2329) therefore summarize the full HBSGRP–UGEV hierarchy used in this study.
 The likelihood function for the data layer is specified as each extreme value sample $X_{k,i}$, observed for event $i$ within group $k$, is modeled as:
-$q_{\text{data}}(X_{k,i}\mid\Theta)=\prod_{k=1}^{K}\prod_{i=1}^{N_{k}}\frac{1}{\sigma_{k,i}}\exp\left\{-\left[1+\xi_{k,i}\left(\frac{X_{k,i}-\mu_{k,i}}{\sigma_{k,i}}\right)\right]^{-1/\xi_{k,i}}\right\}\left[1+\xi_{k,i}\left(\frac{X_{k,i}-\mu_{k,i}}{\sigma_{k,i}}\right)\right]^{-1-1/\xi_{k,i}}$ (23)
-The location ($\mu_{k,i}$), scale ($\sigma_{k,i}$), and shape ($\xi_{k,i}$) parameters are expressed as functions of group-level fixed-effect covariates $X_{k,j}$ (e.g., road geometry, speed limit) and sample-specific random-effect covariates $W_{k,i,j}$ (e.g., speed, acceleration, time-to-collision) as follows:
-$\mu_{k,i}=\beta_{\mu,0}+\sum_{j=1}^{J}\beta_{\mu,j}X_{k,j}+\sum_{j=1}^{J^{\prime}}\gamma_{\mu,j,k}W_{k,i,j}$ (24)
-$\vartheta_{k,i}=\beta_{\sigma,0}+\sum_{j=1}^{J}\beta_{\sigma,j}X_{k,j}+\sum_{j=1}^{J^{\prime}}\gamma_{\sigma,j,k}W_{k,i,j},\quad\sigma_{k,i}=\exp(\vartheta_{k,i})$ (25)
-$\xi_{k,i}=\beta_{\xi,0}+\sum_{j=1}^{J}\beta_{\xi,j}X_{k,j}$ (26)
-In the above equations, $X_{k,j}$ defined as a group-level fixed-effect covariate $j$, $W_{k,i,j}$is a sample-specific random-effect covariate $j$.
-$\beta_{\mu,0},\beta_{\sigma,0},\beta_{\xi,0}$ as fixed intercepts for respective parameters.
-$\beta_{\mu,j},\beta_{\sigma,j},\beta_{\xi,j}$: fixed-effect coefficients.
-$\gamma_{\mu,j,k},\gamma_{\sigma,j,k}$: random-effect coefficients, varying across groups $k$.
-Therefore, the likelihood for the process layer, modeling within-group variability, is defined as:
+$q_{\text{data}}(X_{k,i}\mid\Theta)=\prod_{k=1}^{K}\prod_{i=1}^{N_{k}}\frac{1}{\sigma_{k}}\exp\left\{-\left[1+\xi_{k,i}\left(\frac{X_{k,i}-\mu_{k,i}}{\sigma_{k,i}}\right)\right]^{-1/\xi_{k,i}}\right\}\left[1+\xi_{k,i}\left(\frac{X_{k,i}-\mu_{k,i}}{\sigma_{k,i}}\right)\right]^{-1-1/\xi_{k}}$ (23)
+where $\mu_{k,i}$, $\sigma_{k,i}$, and $\xi_{k,i}$ are the location, scale, and shape parameters of the UGEV distribution for observation $i$ in group $k$.
+This layer expresses how block maxima of 2D-TTC are modeled as extreme outcomes that differ across groups and blocks.
+Parameters are expressed as functions of group-level fixed-effect covariates $Y_{k,j}$ (e.g., lane no, lane width, median type, etc.) and sample-specific random-effect covariates $W_{k,i,j}$ (e.g., speed, acceleration, distance, etc.) as follows:
+$\mu_{k,i}=\beta_{\mu,0}+\sum_{j=1}^{J}\beta_{\mu,j}Y_{k,j}+\sum_{j=1}^{J^{\prime}}\gamma_{\mu,j,k}W_{k,i,j}$ (24)
+$\vartheta_{k,i}=\beta_{\sigma,0}+\sum_{j=1}^{J}\beta_{\sigma,j}Y_{k,j}+\sum_{j=1}^{J^{\prime}}\gamma_{\sigma,j,k}W_{k,i,j},\quad\sigma_{k,i}=\exp(\vartheta_{k,i})$ (25)
+$\xi_{k,i}=\beta_{\xi,0}$ (26)
+In these expressions, $Y_{k,j}$ denotes the $j$-th group-level fixed-effect covariate for group $k$, and $W_{k,i,j}$ denotes the $j$-th observation-level covariate for event $i$ within group $k$.
+The parameters $\beta_{\mu,0}$, $\beta_{\sigma,0}$, and $\beta_{\xi,0}$ are global intercepts for the location, log-scale, and shape parameters, respectively.
+The coefficients $\beta_{\mu,j}$, $\beta_{\sigma,j}$, and $\beta_{\xi,j}$ are global fixed-effect coefficients associated with geometric and static traffic variables, whereas $\gamma_{\mu,j,k}$ and $\gamma_{\sigma,j,k}$ are grouped random-effect coefficients that vary across groups $k$ and are associated with dynamic, time-varying covariates.
+Taken together, Eqs.
+(23)–(26) show how the HBSGRP model combines corridor-level geometric characteristics (through fixed effects on $Y_{k,j}$) with time-varying driving conditions (through group-specific random effects on $W_{k,i,j}$), making the formulation well suited for corridor-level risk analysis.
+Therefore, the contribution of the process layer, modeling within-group variability, is defined as
 $q_{\text{process}}(\Theta\mid\Psi)=\prod_{k=1}^{K}\prod_{i=1}^{N_{k}}\frac{1}{\sqrt{2\pi\tau_{\mu}^{2}}}\exp\left\{-\frac{1}{2\tau_{\mu}^{2}}(\mu_{k,i}-\bar{\mu}_{k})^{2}\right\}\\ \times\frac{1}{\sqrt{2\pi\tau_{\vartheta}^{2}}}\exp\left\{-\frac{1}{2\tau_{\vartheta}^{2}}(\vartheta_{k,i}-\bar{\vartheta}_{k})^{2}\right\}\times\frac{1}{\sqrt{2\pi\tau_{\xi}^{2}}}\exp\left\{-\frac{1}{2\tau_{\xi}^{2}}(\xi_{k,i}-\bar{\xi}_{k})^{2}\right\}$ (27)
 where $\bar{\mu}_{k}$, $\bar{\vartheta}_{k}$, and $\bar{\xi}_{k}$ are group-specific average parameters.
+The variance components $\tau_{\mu}^{2}$, $\tau_{\vartheta}^{2}$, and $\tau_{\xi}^{2}$ control the amount of between-group heterogeneity and induce shrinkage toward group means, thereby encoding the “grouped random parameters” aspect of the HBSGRP model and allowing different corridor elements to share information while retaining distinct risk profiles.
 Prior distributions are assigned to all model parameters. The priors for fixed and random effect coefficients are as follows:
@@
 $\tau_{\mu}^{2}\sim\text{IG}(\alpha_{\mu},\beta_{\mu}),\quad\tau_{\vartheta}^{2}\sim\text{IG}(\alpha_{\vartheta},\beta_{\vartheta}),\quad\tau_{\xi}^{2}\sim\text{IG}(\alpha_{\xi},\beta_{\xi})$ (29)
-where $\text{IG}(\alpha,\beta)$ denotes an inverse gamma prior distribution ensuring positivity constraints for variance parameters.
+where $\text{IG}(\alpha,\beta)$ denotes an inverse gamma prior distribution ensuring positivity constraints for variance parameters.These prior and hyperprior specifications provide regularization for both fixed and grouped random effects, thereby completing the hierarchical definition of the HBSGRP model.
 3.2 COR estimation
-Building on the HBSGRP–UGEV estimates, the framework computes COR by evaluating the probability that 2D-TTC for both V–V and V–I interactions falls below a critical threshold within each block.
-Unlike severity-stratified approaches, this framework estimates the total number of latent extreme near-miss events regardless of outcome severity, allowing for a holistic assessment of potential COR under mixed traffic conditions.
-Let $k,i$ denote the block $i$ within the group $k$, where groups correspond to spatial or functional corridor classifications (e.g., intersections or directional segments).
-The minimum 2D-TTC in each block is modeled with the HBSGRP-UGEV distribution, parameterized by $\mu_{k,i}$, $\sigma_{k,i}$, and $\xi_{k,i}$.
-Let $\omega$ be the critical 2D-TTC threshold below which a near-miss is considered crash-prone.
-The probability of a crash-level event in block $i$ of group $k$ is:
-$P_{k,i}^{\text{crash}}=\Pr(2D-TTC_{k,i}\leq\omega)=1-\exp\left\{-\left[1-\xi_{k,i}\left(\frac{\omega-\mu_{k,i}}{\sigma_{k,i}}\right)\right]^{-1/\xi_{k,i}}\right\},\quad\text{where }1-\xi_{k,i}\left(\frac{\omega-\mu_{k,i}}{\sigma_{k,i}}\right)>0$ (30)
-Assuming a homogeneous Poisson process for exceedances, the expected COR for the group $k$ is obtained by aggregating exceedance probabilities across blocks and normalizing by exposure duration $T$.
-Here $y_{k,i}$ is the number of near-misses (V-V or V-I) in block $i$, and $N_{k}$ the number of blocks in group $k$.
-The expected COR for group $k$ is:
-$COR_{k}=\frac{1}{T}\sum_{i=1}^{N_{k}}y_{k,i}\cdot P_{k,i}^{\text{crash}}=\frac{1}{T}\sum_{i=1}^{N_{k}}y_{k,i}\cdot\left[1-\exp\left\{-\left[1-\xi_{k,i}\left(\frac{\omega-\mu_{k,i}}{\sigma_{k,i}}\right)\right]^{-1/\xi_{k,i}}\right\}\right]$ (31)
-The total expected crash frequency across the study area is then:
-$CF_{\text{total}}=\sum_{k=1}^{K}COR_{k}=\frac{1}{T}\sum_{k=1}^{K}\sum_{i=1}^{N_{k}}y_{k,i}\cdot\left[1-\exp\left\{-\left[1+\xi_{k,i}\left(\frac{\omega-\mu_{k,i}}{\sigma_{k,i}}\right)\right]^{-1/\xi_{k,i}}\right\}\right]$ (32)
-This formulation integrates exceedance probabilities from the HBSGRP–UGEV model with observed V–V and V–I near-miss counts, maintaining a direct link between real-world interaction intensity and modeled crash propensity.
-By separating probability from exposure, the framework yields interpretable, proactive crash frequency estimates aligned with EVT-based safety assessment.
+Building on the HBSGRP–UGEV model, COR is quantified by mapping block-level extreme near-miss severity to the probability of a crash-level outcome.
+For each corridor element $k$ (intersection or directional segment), the trajectory stream is partitioned into fixed-duration interaction blocks, and the minimum 2D–TTC in each block is extracted and transformed to the UGEV scale (Section 3.1.1).
+These block extremes form a sequence $X_{k,i}$, which is modeled using a nonstationary UGEV distribution with parameters $\mu_{k,i}$, $\sigma_{k,i}$, and $\xi_{k,i}$.
+On the transformed scale, a crash-level event corresponds to the physical boundary at which the $\omega=0$ (i.e., 2D–TTC =0=0).
+Therefore, the probability of a crash occurring within the interaction window $i$ in group $k$ is defined as the upper-tail probability of the fitted UGEV distribution evaluated at $\omega=0$:
+$P_{k,i}^{\text{crash}}=\Pr(X_{k,i}\geq\omega)=1-G(\omega;\mu_{k,i},\sigma_{k,i},\xi_{k,i})=1-\exp\left\{-\left[1+\xi_{k,i}\left(\frac{\omega-\mu_{k,i}}{\sigma_{k,i}}\right)\right]^{-1/\xi_{k,i}}\right\},\omega=0$ (30)
+Aggregating these block-level crash probabilities over an exposure period $T$ yields the expected COR for corridor element $k$:
+$COR_{k}=\frac{1}{T}\sum_{i=1}^{N_{k}}P_{k,i}^{\text{crash}}$ (31)
+where $N_{k}$ is the number of blocks observed for group $k$. The total expected crash frequency across the corridor is then obtained by summing over all groups:
+$CF_{\text{total}}=\sum_{k=1}^{K}COR_{k}$ (32)
+Negative TTC thresholds (e.g., 0.1-0.1 s to 3-3 s) are not interpreted as crash boundaries.
+Instead, they are used solely to probe increasingly severe regions of the near-miss distribution and to compare relative crash-prone risk across locations.
+Accordingly, COR should be interpreted as a model-based measure of crash-prone extreme conflict risk extrapolated from near-miss interactions, rather than as a direct count or prediction of observed crashes.
 4 Data description
-The effectiveness of real-time COR estimation is critically dependent on the fidelity and coverage of trajectory data.
-Traditional sensing technologies, such as roadside cameras, loop detectors, and fixed LiDAR, have supported safety analysis by capturing speed, spacing, and acceleration patterns (Islam et al., 2021; Yuan et al., 2019; Li et al., 2020).
-However, these systems suffer from narrow fields of view, occlusion, overlapping detections, and high maintenance costs, which limit scalability and data quality (St-Aubin et al., 2013).
-Recent advances in AV sensing offer a promising alternative.
-AV datasets provide high-frequency, multimodal trajectories with unprecedented spatial precision, allowing for detailed analysis of kinematics, spacing, and maneuvers.
-This richness allows for realistic near-miss detection in complex traffic environments.
-Despite this potential, AV data remain underutilized in generalized frameworks for real-time COR estimation.
-For example, studies using the Waymo Open Dataset have focused mainly on localized near-miss events without incorporating roadway geometry or V-I interactions, limiting broader applicability (Anis et al., 2025b).
-To address these challenges, this study uses the Argoverse-2 dataset (Wilson et al., 2023), which was collected in 2021 across six U.S.
-cities (Austin, Detroit, Miami, Palo Alto, Pittsburgh, and D.C.) using SAE Level 4 AVs.
-Argoverse-2 comprises more than 250k driving scenarios that span 113 roadway segments, with each 11-second scenario sampled at 10 Hz.
-Compared to Waymo Open (Ettinger et al., 2021), Lyft Level-5 (Houston et al., 2021), and nuScenes (Caesar et al., 2020), Argoverse-2 is uniquely curated to emphasize safety-critical interactions, such as intersection near-misses, lane changes, and pedestrian crossings, conditions most indicative of crash risk.
-A defining feature of Argoverse-2 is its scene-adaptive sampling, which prioritizes high-interaction episodes rather than routine free-flow traffic.
-Scenarios are annotated with agent IDs, AV driving mode, and map-relative lane-following behaviors.
+The effectiveness of real-time COR estimation depends on the spatiotemporal fidelity and contextual coverage of trajectory data.
+Traditional sensing technologies (e.g., roadside cameras, loop detectors, fixed LiDAR) can support safety analysis by capturing speed, spacing, and acceleration patterns (Islam et al., 2021; Yuan et al., 2019; Li et al., 2020).
+However, these systems often face limited spatial coverage, occlusion, overlapping detections, and deployment/maintenance burdens, which constrain scalability and data quality (St-Aubin et al., 2013).
+In contrast, AV datasets provide high-frequency, multi-agent trajectories with precise localization, enabling consistent reconstruction of short-horizon kinematics and interaction geometry needed for near-miss-based COR modeling.
+Despite this potential, AV data remain underutilized in generalized real-time COR frameworks.
+For example, studies using the Waymo Open Dataset have mainly analyzed localized conflicts and often omit HD-map-aligned roadway boundaries needed to represent V–I exposure, limiting transferability to corridor-scale inference (Anis et al., 2025b).
+To address these challenges, this study uses the Argoverse-2 dataset (Wilson et al., 2023), collected in 2021 across six U.S.
+cities (Austin, Detroit, Miami, Palo Alto, Pittsburgh, and Washington, D.C.) using SAE Level 4 AVs.
+Argoverse-2 contains >>250k driving scenarios spanning 113 roadway segments, with each 11-second scenario sampled at 10 Hz.
+Compared to Waymo Open (Ettinger et al., 2021), Lyft Level-5 (Houston et al., 2021), and nuScenes (Caesar et al., 2020), Argoverse-2 is curated to emphasize interaction-rich episodes (e.g., intersection encounters, lane changes, and complex crossings) that are informative for extreme-conflict modeling.
+A defining feature is its scene-adaptive sampling, which prioritizes high-interaction moments rather than routine free-flow conditions.
+Because this design intentionally over-samples kinematically and socially unusual behavior (Wilson et al., 2023), the extracted near-misses and inferred COR should be interpreted as conditional on safety-critical traffic states, rather than as unconditional averages over typical corridor operations.
+Scenarios include agent identifiers, AV status, and map-based behavioral annotations.
 The sensor suite comprises seven ring cameras and two front-facing stereo cameras (20 Hz), supported by dual 64-beam VLP-32C LiDAR sensors, which produce  107k 3D points per frame.
-HD maps contain lane centerlines, crosswalks, medians, and drivable areas registered to local coordinates.
+HD maps contain lane centerlines, crosswalks, medians, and drivable areas, all registered to local coordinates.
 Table 1: Representative open-source AV datasets with applications in safety and crash-risk analysis
@@
 Argoverse-2 was selected after benchmarking open-source AV datasets (Table 1).
-Waymo Open offers a greater total mileage but lower near-miss density and no AV-behavior labels.
-Lyft Level-5 is confined to a single city and lacks scene-level filtering.
-nuScenes samples at 2 Hz, limiting the capture of rapid interactions.
-In contrast, Argoverse-2 balances spatial coverage, sensor fidelity, near-miss-rich scenarios, and behavioral annotations, making it well-suited to this study’s goals.
-This study used motion-forecasting and HD-map modules of Argoverse-2 to extract vehicle trajectories, compute near-miss indicators, and encode roadway geometry.
-4.1 Data Preprocessing
-To operationalize the proposed framework, both the motion forecasting and HD map modules of the Argoverse-2 dataset were processed through a structured pipeline.
-The preprocessing consisted of four key stages: (i) extraction of multi-agent trajectories, (ii) transformation of coordinates from local map space to global geographic space, (iii) integration with HD map features to contextualize V-I interactions, and (iv) selection of a high-risk urban corridor for detailed analysis.
-4.1.1 Motion forecasting dataset
-Each motion forecasting scenario follows a hierarchical structure with three levels: scenario metadata, agent tracks, and frame-level states.
-Scenarios are 11s duration and are sampled at 10 Hz, producing 110 time steps per vehicle for fine-grained interaction modeling.
-At the scenario level, metadata includes a unique scenario ID, the recording city, and the focal agent, typically the ego AV.
-Each scenario then contains multiple agent tracks, where each track corresponds to a road user and specifies the object type (e.g., vehicle, pedestrian, cyclist), scoring category (focal, scored, or unscored), and a time-ordered sequence of states.
-Frame-level states capture position, velocity, and heading at each timestep, forming the basis for computing 2D-TTC.
-To implement the framework, we developed a Python-based batch pipeline using the Argoverse-2 SDK.
-The pipeline parsed motion forecasting scenarios across all cities and extracted tracks for all agents—not just the focal vehicle—to retain critical multi-agent interactions.
-For each track, frame-level kinematic attributes were retrieved, including $(x,y)$ position, heading, velocity components, and timestamps.
-Because positions were reported in local, city-specific frames, all coordinates were transformed into a globally referenced WGS84 system.
-This involved (i) projecting each city’s local origin into its UTM zone (e.g., Zone 17 for Miami, Pittsburgh, and Detroit), and (ii) offsetting local $(x,y)$ positions and inverse-projecting them to latitude–longitude using pyproj.
-This geospatial normalization provided consistent referencing across cities and ensured compatibility with HD map overlays.
-Batch scripts automatically traversed scenario folders, performed attribute extraction and coordinate conversion, and exported unified CSV datasets containing both original city-frame and converted lat/lon values.
-These datasets were then used for spatial filtering, interaction detection, and overlay validation.
-The processed files formed the basis for V–V and V–I near-miss detection.
-For each valid scenario, pairwise distances and relative approach rates were computed for all vehicle pairs, while V-I distances were derived from buffered polygons of lane edges and curbs in the HD map.
-Scenarios with no substantive interaction (e.g., straight-lane cruising without lane changes or turns) were excluded, ensuring the dataset emphasized behaviorally informative, high-risk conditions consistent with proactive safety modeling.
-Non-vehicle agents (e.g., pedestrians, cyclists) were also excluded at this stage, though the pipeline is extensible to incorporate vulnerable road users in future work.
-4.1.2 HD map
+Waymo Open offers greater overall mileage but does not explicitly target interaction-dense segments at the same level, Lyft Level-5 is limited to a single city, and nuScenes samples at 2 Hz, reducing sensitivity to rapid interactions.
+Argoverse-2 offers a practical balance of sampling rate, multi-agent density, HD-map detail, and interaction-centric scenario selection, making it well-suited for geometry-aware near-miss extraction and corridor-scale COR estimation.
+This study used vehicle trajectories and HD-map details, which provide multiple agent tracks labeled by object type (e.g., vehicle, pedestrian, cyclist).
+To operationalize the proposed framework, we processed these modules through a four-stage pipeline: (i) extraction of multi-agent trajectories and necessary information, (ii) coordinate transformation from local map space to global geographic space, (iii) integration with HD-map boundary features to define V–I exposure, and (iv) selection of a high-risk urban corridor for detailed analysis.
+4.1 Motion forecasting dataset
+Each Argoverse-2 motion-forecasting scenario has a three-level structure: scenario metadata, agent tracks, and frame-level states.
+Scenarios are 11 s long and provide a birds-eye-view 2D centroid and heading for each tracked object at 10 Hz (110 time steps), enabling fine-grained interaction modeling.
+Scenario metadata includes a unique scenario ID, recording city, and a focal agent (typically the ego AV).
+Each scenario contains multiple agent tracks with object type (e.g., vehicle, pedestrian, cyclist), scoring category (focal/scored/unscored), and a time-ordered sequence of states.
+Frame-level states provide position, velocity components, and heading, which form the basis for computing 2D-TTC.
+We directly use these trajectories as provided by Argoverse-2; tracks are produced by Argo’s internal multi-sensor fusion and map-aligned ego poses, which help keep trajectories tightly registered to the HD map.
+To implement the proposed framework, we developed a Python batch-processing pipeline using the Argoverse-2 SDK.
+The pipeline parses scenarios and extracts tracks for all agents (not only the focal vehicle) to preserve the multi-agent interaction context.
+For each track, we retrieve frame-level kinematic attributes, including $(x,y)$ position, heading, velocity components, and timestamps.
+For the present analysis, we retain only motor-vehicle agents and filter out pedestrians and cyclists; the same pipeline can be extended to incorporate vulnerable road users in future work.
+Because positions are reported in city-specific local coordinate frames, all coordinates are transformed to a globally referenced WGS84 system.
+This is done by projecting each city’s local origin into the appropriate UTM zone (e.g., Zone 17 for Miami, Pittsburgh, and Detroit), offsetting local $(x,y)$ values, and inverse-projecting to latitude/longitude using pyproj.
+Batch scripts traverse scenario folders, extract attributes and convert coordinates, and export unified CSV files containing both the original city-frame coordinates and the converted lat/lon values.
+To verify that this transformation preserves trajectory geometry, we conducted a visual geospatial validation for the Biscayne Boulevard corridor by overlaying converted trajectories and HD-map features on satellite imagery (Fig.
+4.2).
+Vehicle paths closely follow mapped lane centerlines and intersection geometry, indicating residual registration error is small relative to lane width and is acceptable for conflict localization.
+In addition to the motion-forecasting trajectories, we use the Argoverse-2 sensor annotation files $(\texttt{annotations.feather})$, which provide 3D bounding-box dimensions for each tracked object at each timestamp (including length_m and width_m).
+The motion-forecasting tracks are joined to these annotations using the shared scenario identifier, timestamp_ns, and track_uuid.
+For each vehicle track, the corresponding length_m and width_m values are used as the plan-view dimensions of the oriented rectangular footprint in the geometry-aware 2D-TTC computation.
+The processed files form the input for V–V and V–I near-miss detection.
+For each valid scenario, we evaluate V-V interactions by computing pairwise relative spacing and approach rates across vehicle pairs, while V–I proximity is evaluated relative to HD-map-derived lane edges/curb boundaries.
+Scenarios without substantive interaction (e.g., sustained straight-lane cruising without turning, merging, or lateral maneuvers) are excluded to emphasize behaviorally informative, high-interaction conditions consistent with proactive extreme-conflict modeling.
+4.2 HD map
 The Argoverse-2 HD map dataset provides vectorized representations of roadway geometry and drivable space, enabling infrastructure-aware risk estimation.
@@
 Each local $(x,y)$ coordinate was then offset by this origin and inverse-projected into latitude–longitude, ensuring spatial consistency across cities.
-Once transformed, all HD map components: lane boundaries, centerlines, and crosswalks—were indexed using their global coordinates and converted into shapefiles or GeoPandas geometries.
+Once transformed, all HD map components: lane boundaries, centerlines, and crosswalks—were indexed by their global coordinates and converted into shapefiles or GeoPandas geometries.
 This geospatial transformation enabled efficient spatial querying, collision proximity analysis, and visual overlay with satellite imagery and trajectory data.
@@
 From the processed dataset, a high-density urban corridor in Miami was present to showcase the transformation (see Fig.
-4.1.2).
+4.2).
 This illustrates the transformation of AV data into real-world coordinates, its alignment with HD maps and satellite imagery, and the resulting traffic patterns that highlight interaction hotspots across intersections and midblocks.
 Fig.
-3.: AV data transformed to real-world coordinates and aligned: (a) trajectories with lane boundaries and crosswalks, (b) validation with satellite imagery, and (c–f) directional traffic patterns highlighting interaction hotspots.
-4.1.3 Trajectory preprocessing and derivation of vehicle dynamics
+3.: AV data transformed to real-world coordinates and aligned: (a) trajectories with lane boundaries and crosswalks, (b) validation with satellite imagery, and (c–f) directional traffic patterns highlighting interaction hotspots
+4.3 Trajectory preprocessing and derivation of vehicle dynamics
 Before computing 2D-TTC, a series of preprocessing steps was applied to ensure the precision, continuity, and physical plausibility of vehicle motion data from the Argoverse-2 dataset.
-These steps included the extraction of raw trajectory data, trajectory smoothing and filtering, and the derivation of dynamic variables such as speed, acceleration, and steering angle.
+These steps included extracting raw trajectory data, smoothing and filtering the trajectories, and deriving dynamic variables such as speed, acceleration, and steering angle.
 First, all multi-agent trajectories were transformed from local city-specific frames to global WGS84 coordinates, enabling consistent alignment with HD map geometries.
@@
 To suppress such noise while preserving curvature and lane-change dynamics, cubic B-spline smoothing was applied to the position coordinates (Eilers and Marx, 1996; Choi et al., 2024).
-This method effectively removed local fluctuations while retaining key geometric features (see Fig.4.1.3).
+This method effectively removed local fluctuations while retaining key geometric features (see Fig.
+ref fig:4).
 Stationary or irrelevant agents, such as parked vehicles and bicycles, were filtered out to reduce computational load and focus only on interaction-relevant entities.
@@
 4.4 Study area
-Under realistic and heterogeneous urban traffic conditions, a representative high-volume corridor was selected along Biscayne Boulevard in Miami, Florida.
-This arterial extends from Downtown Miami through the Upper East Side toward the northern city limits, serving as a critical connector to major activity centers, including the Miami International Airport region.
-It is characterized by multimodal interactions, geometric complexity, frequent turning maneuvers, and diverse traffic control features.
-The corridor was chosen for its rich coverage in the Argoverse-2 motion forecasting dataset, HD map, and its suitability for extracting sufficient samples for extreme near-miss scenarios.
+Under realistic and heterogeneous urban traffic conditions, we selected a high-volume study corridor along Biscayne Boulevard in Miami, Florida.
+This arterial links Downtown Miami to the Upper East Side and northern city limits and is characterized by dense access activity, frequent turning and lane-changing maneuvers, and diverse traffic control.
+The corridor was chosen because it has high scenario density in Argoverse-2 and high-fidelity HD map coverage, and because its mapped curb and median boundaries provide explicit roadway-edge geometry needed to consistently define V–I near-misses.
 As shown in Fig.
-4.2, the study area was extracted using ArcGIS Pro based on scenario density, HD map fidelity, and heterogeneity of vehicle maneuvers.
-The extracted network includes a continuous corridor with multiple intersections, midblock segments, and bidirectional traffic operations.
-For analytical consistency, the corridor was subdivided into nine directional segments (S1–S9) and ten intersections (I1–I10).
-Segment boundaries were defined midblock to maintain flow continuity, while intersections were delineated to capture near-miss-prone zones such as turning paths, crosswalks, and lane merges.
-Each subregion was mapped with lane boundaries, crosswalks, and vehicle trajectories to preserve geometric fidelity and directional alignment.
-Insets in Fig.
-4.2 illustrate the spatial configuration of the segmented units, which serve as the basis for grouped modeling and localized risk estimation.
+4.4, the study area was extracted in ArcGIS Pro based on scenario density, HD map fidelity, and maneuver diversity.
+The resulting network forms a continuous corridor with bidirectional operations, multiple intersections, and midblock sections.
+For analytical consistency, the corridor was subdivided into 10 intersections (I1–I10) and 18 directional segments organized into nine segment pairs (S1–S9).
+Segment limits were defined midblock to preserve operational continuity, while intersection extents were delineated to capture conflict-prone subareas (turning paths, merges, and channelization transitions).
+These subregions define the hierarchical modeling structure and enable direction-aware inference.
 Fig. 5.: Segmented study corridor along Biscayne Boulevard, Miami, with lane boundaries and crosswalks. Insets show direction-specific segments (S1–S9) and intersections (I1–I10)
-Table 2 provides a detailed profile of the 28 study subregions, consisting of 10 intersections and 18 directional segments.
-Intersections generally show fewer scenarios and unique vehicles (e.g., I2 with 47 scenarios and 220 vehicles, I4 with 96 scenarios and 568 vehicles) compared to adjacent midblock segments, reflecting shorter dwell times but higher near-miss intensity.
-In contrast, segments accumulate a significantly higher number of scenarios and vehicles due to sustained travel.
-For example, S7 (NE 19-17 St) records 308 northbound and 332 southbound scenarios involving more than 4000 vehicles, while S8 (NE 17–15 St) captures nearly 4400 vehicles in 608 scenarios.
+Table 2 summarizes the 28 study subregions.
+Intersections generally include fewer scenarios and unique vehicles than adjacent segments (e.g., I2: 47 scenarios/220 vehicles; I4: 96 scenarios/568 vehicles), reflecting shorter dwell time within intersection boundaries.
+By contrast, segments accumulate substantially more scenarios and vehicles due to sustained travel (e.g., S7 and S8 each include several hundred scenarios and >>4000 unique vehicles)
 Table 2: Profile of study locations along Biscayne Boulevard
@@
 I10 NE 13 St 232 1139
-Lane configurations vary between 2 and 3 lanes per direction, with average widths ranging from 3.1m (S7) to 3.6m (S6), indicating subtle but important differences in cross-sectional design.
-Driveway and minor intersection density also differ across the corridor: segments such as S3 and S4 contain four access points, while others like S5 and S6 have few or none, creating heterogeneous exposure to side-entry near-misses.
-The median type alternates between raised (e.g., S1–S2) and flush-paved (e.g., S3–S9), influencing turning maneuvers and mid-block encroachments.
-Overall, the table highlights that intersections serve as near-miss-prone nodes with concentrated but lower-volume interactions, whereas segments provide richer samples of continuous traffic flow, characterized by greater geometric variability.
-Building on the spatial segmentation and detailed subregion profiles outlined above, the next step involves quantifying dynamic safety conditions using the proposed 2D-TTC framework (Algorithm 1).
-For each block, extreme near-miss events were identified by forward simulating vehicle trajectories under the geometric and traffic conditions described in Table 2.
-Both V–V and V–I interactions were modeled to reflect the heterogeneous near-miss environments of intersections and segments.
-Trajectories were projected over a 3s horizon, and spatial overlaps with surrounding traffic or roadway boundaries were detected as potential near-miss events.
+Lane configurations range from 2 to 3 lanes per direction, with average lane widths of 3.1 m (S7) to 3.6 m (S6).
+Access density also varies across segments (e.g., multiple driveways/minor access points in S3–S4 versus few or none in S5–S6), creating heterogeneous exposure to lateral disturbances and side-entry maneuvers.
+Median treatments alternate between raised and flush-paved sections, which can affect turning opportunities and the potential for edge encroachment.
+Building on the spatial segmentation and subregion profiles above, we next quantify time-varying safety conditions using the proposed 2D-TTC pipeline (Algorithm 1).
+Because this study adopts the BM sampling strategy, there is no global consensus on the optimal block duration.
+Prior BM–GEV studies based on video data often employ long, time-based blocks (e.g., several minutes) and select block sizes based on goodness-of-fit criteria.
+In contrast, AV trajectory datasets consist of short scenarios (e.g., Argoverse-2: 17 s; Waymo: 20 s; KITTI: 1–2 min) and high-frequency interaction windows.
+Consequently, this study adopts an interaction-window BM strategy, consistent with real-time near-miss modeling in our prior work (Anis et al., 2025b), in which blocks are defined by unique interaction units (Ali et al., 2022a).
+Specifically, each continuously observed interaction window, either a V–V or a V–I interaction within a scenario, is treated as a block, and the block extreme is defined as the minimum 2D-TTC over the 11 s (110-frame) window.
+For each interaction window, the near-miss process is evaluated at every frame by forward-simulating motion using an RK4-based kinematic projection and subsequently computing 2D-TTC.
+Thus, each retained extreme summarizes 110 high-frequency kinematic states and 110 interaction assessments, providing substantial within-block information despite the short scenario duration.
+Because multiple vehicles are present in a single 11s scenario, multiple interaction windows occur simultaneously.
+We therefore distinguish between (i) candidate near-miss detected across vehicles and frames within a scenario and (ii) the single extreme block retained for EVT per interaction window.
+Defining blocks by interaction windows also avoids repeatedly selecting adjacent frame-wise minima from the same persistent interaction, which would overweight long-lasting conflicts and amplify temporal dependence.
+For each retained block, extreme near-miss events are identified by forward-simulating vehicle trajectories under the geometric and traffic conditions summarized in Table 2.
+Trajectories are projected over a 3s horizon, and spatial overlaps with surrounding traffic or roadway boundaries are evaluated as potential near-miss events.
+To focus the BM sample on genuinely unsafe edge-encroachment behavior, we retain V–I encounters only when the interaction window contains short 2D-TTC values with moderate-to-high approach speeds, filtering out routine low-speed turning proximity that can appear as boundary approach under constant-input projection.
 Representative outcomes from this process are shown in Fig.
-4.2, illustrating the spatial dynamics of predicted V–V and V–I near-misses within the Miami study corridor.
+4.4, illustrating the spatial dynamics of predicted V–V and V–I near-misses within the Miami study corridor.
 These event visualizations provide the empirical basis for block-level 2D-TTC extraction across the 10 intersections (I1–I10), enabling localized evaluation of crash-prone conditions.
-Table 3: Summary statistics for V–V and V–I near-misses at Intersection
+Finally, all microscopic covariates (e.g., relative speed, relative acceleration/deceleration, jerk, heading/steering differences) are extracted from the same interaction window and aligned to the extreme event, either evaluated at the frame where the extreme 2D-TTC occurs.
+Table 3: Summary statistics for V–V and V–I near-misses at Intersections
 Metric V–V V–I
@@
 Continuous variables
-TTC 1.617 0.11 2.99 0.863 1.304 0.10 2.99 0.889
-Relative speed 5.257 0.00 23.724 4.446 8.753 0.00 22.276 4.035
-Relative acc/dec -0.535 -13.706 10.343 2.227 0.161 -5.925 13.199 0.702
-Relative distance 11.673 2.419 46.857 6.844 1.520 0.60 3.80 1.110
-Jerk -5.550 -13.609 9.315 2.652 -7.900 -10.800 2.800 14.700
-Heading difference 1.077 0.00 6.251 1.410 -0.045 -3.138 3.141 1.647
-Steering difference 0.409 0.00 3.142 0.666 -0.006 -1.571 1.571 0.130
-Volume 13.462 2.00 31.00 5.969 10.461 1.00 31.00 5.392
-Vehicle length 4.78 3.18 6.92 0.49 5.1 3.95 12.65 1.30
-Vehicle width 2.22 1.76 3.16 0.30 2.04 1.36 3.08 0.23
+TTC (s) 1.617 0.11 2.99 0.863 1.304 0.10 2.99 0.889
+Relative speed (m/s) 5.257 0.00 23.724 4.446 8.753 0.00 22.276 4.035
+Relative acc./dec. (m/s2) -0.535 -13.706 10.343 2.227 0.161 -5.925 13.199 0.702
+Relative distance (m) 11.673 2.419 46.857 6.844 1.520 0.60 3.80 1.110
+Jerk ($\mathrm{m/s^{3}}$) -5.550 -13.609 9.315 2.652 -7.900 -10.800 2.800 14.700
+Heading difference (rad) 1.077 0.00 6.251 1.410 -0.045 -3.138 3.141 1.647
+Steering difference (rad) 0.409 0.00 3.142 0.666 -0.006 -1.571 1.571 0.130
+Traffic volume (veh) 13.462 2.00 31.00 5.969 10.461 1.00 31.00 5.392
+Vehicle length (m) 4.78 3.18 6.92 0.49 5.1 3.95 12.65 1.30
+Vehicle width (m) 2.22 1.76 3.16 0.30 2.04 1.36 3.08 0.23
 Categorical variable: Near-miss location (Proportion %)
@@
 Fig. 6.: Representative intersection-level near-miss detection using the 2D-TTC framework: (a) V–V and (b) V–I
-Descriptive statistics of intersection-level V–V and V–I near-misses are summarized in Table 3.
-The continuous indicators include 2D-TTC, relative speed, acceleration, distance, jerk, and angular deviations (heading and steering), along with vehicle dimensions and exposure (volume).
-On average, 2D-TTC values for both near-miss types range between 1.3 and 1.6 s, with extreme cases falling below 0.1 s, reflecting highly critical near-miss conditions.
-V–I near-misses are associated with higher relative speeds and tighter distance gaps, indicating more abrupt encroachment against roadway boundaries.
-Conversely, V–V near-misses exhibit larger heading and steering deviations, along with higher jerk variability, suggesting unstable maneuvers such as sudden turns or evasive swerves.
-The categorical distribution of near-misses further illustrates key behavioral differences.
-Nearly three-quarters (72%) of V–V events occurred during through movements, with fewer during left (26%) or right (9%) turns.
-In contrast, V–I near-misses were almost entirely turning-related, with 46% occurring in left turns and 54% in right turns, and none during through movements.
-This pattern highlights that infrastructure-related near-misses typically occur during lateral maneuvers at intersections, where vehicles cross lane boundaries or encroach on curbs and medians.
-Beyond intersections, the analysis was extended to directional roadway segments (S1–S9) using the same 2D-TTC framework (See Fig.
-4.2).Table 4 summarizes descriptive statistics of both V–V and V–I near misses at the segment level.
-For V–V near-misses, average TTC values were longer (1.83 s) than those observed at intersections, suggesting greater temporal headway between vehicles in segment settings.
-However, the extreme minima (0.11 s) and high maximum relative speeds (up to 33.9 m/s) highlight the presence of aggressive overtaking and high-speed convergence.
-Relative distances also exhibited a much broader range (up to 141.5 m), indicating that near-misses often arise in more dispersed longitudinal interactions rather than tightly constrained intersection maneuvers.
+Descriptive statistics for intersection-level V–V and V–I near-misses are reported in Table 3.
+The continuous indicators include 2D-TTC and interaction dynamics (relative speed, relative acceleration/deceleration, relative distance, jerk, and heading/steering differences), along with vehicle dimensions and exposure (volume).
+Mean 2D-TTC values are approximately 1–2s (V–V: 1.62s; V–I: 1.30s), while the most safety-critical cases lie in the lower tail, with minima near 0.10–0.11s.
+Table further quantifies the lower-tail structure of 2D-TTC, showing that crash-prone near-miss risk is concentrated in a narrow extreme range.
+To ensure that extremely small 2DTTC values from geometry-oriented rectangle footprints (e.g., spurious corner overlap), the corresponding scenarios were manually reviewed by visualizing the time-evolving oriented bounding boxes and headings; the critical cases reflected plausible relative motion and geometry rather than numerical overlap errors.
+V–I near-misses tend to exhibit higher relative speeds and smaller distance gaps, consistent with abrupt boundary approaches, whereas V–V near-misses show larger heading/steering deviations and higher jerk variability, consistent with unstable maneuvers (e.g., abrupt turns or evasive swerves).
+Categorical summaries further highlight differences: most V–V events occur during through movements, whereas V–I near-misses are concentrated in lateral/turning maneuvers, consistent with edge approach under turning and merge dynamics.
+The analysis was also conducted for directional segments (S1–S9) using the same 2D-TTC pipeline (See Fig.
+4.4).
+Table 4 reports segment-level descriptive statistics.
+For V–V near-misses, mean TTC values are longer than at intersections (1.83 s), suggesting larger headways in midblock traffic, while very small TTC minima and high relative-speed maxima indicate occasional aggressive overtaking and rapid convergence.
+Relative distances span a wider range than at intersections, reflecting more dispersed longitudinal interactions in segment settings.
 Table 4: Summary statistics for V–V and V–I near-misses at Segments
@@
 Continuous variables
-TTC 1.829 0.11 2.99 0.805 1.272 0.10 2.99 0.912
-Relative speed 4.513 0.00 33.909 4.278 10.262 0.000 28.067 4.212
-Relative acc/dec -0.871 -23.014 18.336 3.139 0.164 -14.067 23.343 1.400
-Relative distance 15.361 1.931 141.548 12.773 1.720 0.32 6.56 1.02
-Jerk -7.121 -17.502 8.423 1.552 -5.70 -12.72 3.556 4.77
-Heading difference 0.368 0.000 6.256 0.990 -0.032 -3.102 3.138 1.586
-Steering difference 0.577 0.000 3.142 0.756 -0.001 -1.571 1.571 0.167
-Volume 16.801 1.000 65.000 11.009 12.816 1.000 65.000 7.906
-Vehicle Length 4.74 3.0 7.85 1.11 4.79 2.95 8.65 0.87
-Vehicle width 2.12 1.36 3.08 0.23 2.22 1.76 4 0.18
+TTC (s) 1.829 0.11 2.99 0.805 1.272 0.10 2.99 0.912
+Relative speed ($\mathrm{m/s}$) 4.513 0.00 33.909 4.278 10.262 0.000 28.067 4.212
+Relative acc./dec. (m/s2) -0.871 -23.014 18.336 3.139 0.164 -14.067 23.343 1.400
+Relative distance (m) 15.361 1.931 141.548 12.773 1.720 0.32 6.56 1.02
+Jerk ($\mathrm{m/s^{3}}$) -7.121 -17.502 8.423 1.552 -5.70 -12.72 3.556 4.77
+Heading difference (rad) 0.368 0.000 6.256 0.990 -0.032 -3.102 3.138 1.586
+Steering difference (rad) 0.577 0.000 3.142 0.756 -0.001 -1.571 1.571 0.167
+Traffic volume (veh) 16.801 1.000 65.000 11.009 12.816 1.000 65.000 7.906
+Vehicle Length (m) 4.74 3.0 7.85 1.11 4.79 2.95 8.65 0.87
+Vehicle width (m) 2.12 1.36 3.08 0.23 2.22 1.76 4 0.18
+Driveway intensity ($\mathrm{m^{-1}}$) 0.009 0.000 0.015 0.005 0.009 0.000 0.015 0.005
+Lane width (m) 3.28 2.99 3.96 0.14 3.27 2.99 3.96 0.141
 Lane number 2.73 2 4 0.578 2.51 2.0 4.0 0.52
-Lane width 3.28 2.99 3.96 0.14 3.27 2.99 3.96 0.141
-Driveway intensity 0.009 0.000 0.015 0.005 0.009 0.000 0.015 0.005
 Categorical variable: (Proportion %)
@@
 Fig. 7.: Representative segment-level near-miss detection using the 2D-TTC framework: (a) V–V and (b) V–I
-For V–I near-misses, the mean TTC was shorter (1.27s), with relative speeds peaking at 28.1 m/s, underscoring the abrupt nature of vehicle–boundary encounters.
-Compared with intersections, segment-based V–I near-misses involved longer approach distances but also showed greater instability in jerk and heading, reflecting frequent lateral drifts toward lane edges or curbs.
-These events often occurred near driveways and minor access points, where vehicles deviated from their intended path, creating encroachment risks.
+For V–I near-misses, the mean TTC was shorter with greater relative speeds, underscoring the abrupt nature of vehicle–infrastructure encounters.
+Compared with intersections, segment-level V–I near-misses tend to occur at longer approach distances but exhibit greater variability in jerk and heading, consistent with lateral drift or corrective steering as vehicles move toward lane edges, curbs, or medians.
+These events frequently cluster near driveways and minor access points, where side-entry activity and path adjustments increase the likelihood of encroachment.
+Because the V–I definition is restricted to same-carriageway curbs/medians and the retained encounters are filtered by 2D-TTC, the resulting V–I sample reflects elevated boundary-encroachment risk rather than generic turning or lane-change exposure.
+Segment-level access exposure is captured by driveway intensity, defined as the density of driveways and minor access points per unit segment length.
 The categorical breakdown reinforces these patterns.
@@
 Furthermore, median type also influenced outcomes: 76% of V–V near-misses and 86% of V–I near-misses occurred along undivided sections, underscoring the vulnerability of open medians and flush-paved designs in segment environments.
-Taken together, these segment-level findings reveal that while intersections are hotspots for tightly constrained V-V near-misses, segments carry distinctive risks tied to high-speed encounters and infrastructure encroachments.
-The observed distributions underscore the importance of treating intersections and segments as functionally distinct safety domains when modeling near-miss dynamics and designing targeted countermeasures.
+Taken together, these segment-level results indicate that, while intersections concentrate tightly constrained V–V conflicts, segments exhibit distinct risk mechanisms associated with higher-speed interactions and boundary encroachments.
+This supports treating intersections and segments as functionally different safety domains when modeling near-miss dynamics and designing countermeasures.
 (a) Intersection
@@
 Fig. 8.: BM near-miss frequencies by location type and direction
-After identifying and compiling the BM near-miss events for each observation unit (segments and intersections), the frequencies of extreme V–V and V–I events were systematically aggregated and visualized in Fig.
-4.2.
-This plot illustrates the spatial distribution of extreme near-miss frequencies across 28 locations, comprising 10 intersections and 18 segments, organized into intersection zones, primary-direction segments, and opposite-direction segments.
-This spatial aggregation confirms the integrity of the BM dataset and provides the foundation for modeling.
-The study extends prior approaches by incorporating both V–V and V–I near misses into the HBSGRP–UGEV framework, explicitly integrating roadway geometry and infrastructure features alongside vehicle dynamics.
-Prior research has emphasized the role of vehicle kinematics in crash risk (Desai et al., 2021; Jun et al., 2007; Tak et al., 2015); here, additional geometric covariates, such as lane configuration and boundary proximity, are introduced.
-For example, high steering angles, often indicating abrupt lateral maneuvers, were included as predictors, given their link to boundary incursions and sudden evasive actions that may elevate near-miss risk (Gilbert et al., 2021).
+After extracting block extremes per valid interaction window, the BM extremes are aggregated by location and summarized in Fig.
+4.4.
+The figure reports extreme near-miss frequencies across 28 locations.
+To contextualize exposure, Tables 34 also report frame-level volumes, which varied across scenarios.
+Because the ego-centric LiDAR has a limited sensing range, the observable traffic volume varies across frames and interaction windows and is therefore used as a local exposure measure to contextualize 2DTTC-based extremes.
+The study extends prior EVT-based near-miss work by jointly modeling V–V and V–I extremes under the HBSGRP–UGEV framework and incorporating roadway geometry and infrastructure context alongside vehicle dynamics.
+In addition to kinematic predictors emphasized in earlier studies (Desai et al., 2021; Jun et al., 2007; Tak et al., 2015), we include geometric covariates, such as lane configuration and boundary context; for example, steering-related measures capture abrupt lateral maneuvers associated with boundary incursions (Gilbert et al., 2021).
 Fig.
-4.2 highlights several notable spatial patterns.
-Opposite-direction segments (S82, S72, S62) exhibited the highest V–I frequencies, suggesting greater lateral variability and exposure to curb or median near-misses.
-Primary-direction segments (S11, S21, S81) showed elevated V–V near-miss frequencies, consistent with midblock environments dominated by longitudinal interactions such as close-following, lane changes, or overtaking.
-Among intersections, I7, I8, and I9 emerged as critical hotspots, with elevated frequencies of both near-miss types.
-This dual concentration reflects the operational complexity of intersections, where near-missing maneuvers, pedestrian flows, and signal control interact to amplify crash potential.
-The distribution and magnitude of near-miss counts across all locations validate the effectiveness of the BM approach in capturing rare but safety-critical events.
-Specifically, nearly all study sites meet or exceed the recommended minimum threshold of 30 BM events for reliable GEV-based inference (Zheng et al., 2014), with the exception of intersection I2.
-This ensures that the dataset is statistically robust for subsequent HBSGRP–UGEV analysis.
-The spatial heterogeneity revealed in Fig.
-4.2 further underscores the need to incorporate direction-specific geometric and behavioral covariates within the modeling framework.
-These results provide compelling justification for developing localized, high-resolution safety models that integrate both roadway configuration and observed vehicle dynamics, thereby enabling precise and context-sensitive real-time risk assessments in complex urban environments.
+4.4 also reveals clear spatial patterns: opposite-direction segments (S82, S72, S62) show the highest V–I extreme frequencies, while primary-direction segments (S11, S21, S81) exhibit elevated V–V frequencies consistent with midblock longitudinal interactions.
+Among intersections, I7–I9 emerge as hotspots for both near-miss types, reflecting the compounded operational complexity of signal control, turning movements, and interaction density.
+Overall, the distribution of block extremes supports the BM approach in capturing rare but safety-critical events.
+Nearly all sites meet or exceed the commonly cited minimum of 30 observations for GEV-based inference (Zheng et al., 2014), except for I2, which provides a statistically adequate basis for HBSGRP–UGEV estimation.
+The observed heterogeneity across directions and facility types further motivates the inclusion of direction and location-specific effects to enable context-sensitive, corridor-scale crash-risk estimation.
 5 Modeling estimation result
-This study develops a series of Bayesian hierarchical UGEV models to estimate corridor-wide COR from extreme 2D-TTC near-miss events observed along Biscayne Boulevard in Miami.
-Extreme near-misses are identified using the BM approach, wherein the minimum 2D-TTC within each temporal block is treated as the most critical interaction.
-The theoretical basis and implementation of this sampling methodology are detailed in prior work (2025b).
-Each V–V and V–I interaction is treated as an independent block.
-Since extreme events are relatively sparse at individual sites, direct location-specific estimation often yields unstable posteriors with wide uncertainty intervals.
-To address this issue, a corridor-wide grouping strategy was adopted, pooling samples across functionally similar sites (intersections, segments) while preserving local heterogeneity.
+This study develops a set of Bayesian hierarchical UGEV models to estimate corridor-wide crash occurrence risk (COR) from extreme 2D-TTC near-miss events along Biscayne Boulevard in Miami.
+Extreme near-misses are extracted using a BM strategy; each V–V or V–I defines a block extreme used for EVT inference.
+Because extreme events are sparse at individual sites, direct site-by-site estimation can yield unstable posteriors and wide credible intervals.
+To improve statistical stability while preserving spatial structure, we adopt a corridor-wide grouped formulation.
+Each site defines one group $k$, and all block extremes observed at that site are assigned to the corresponding group.
+This enables partial pooling through corridor-level hyperparameters while allowing selected effects to vary across the 28 groups.
 The HBSGRP–UGEV model is the primary analytical framework, with a simpler fixed-effect hierarchical model (HBSFP) also estimated for comparison.
 For each near-miss type (V–V and V–I), two models were therefore developed: (1) HBSFP and (2) HBSGRP.
-The grouped random parameter formulation in HBSGRP allows selected coefficients to vary across spatial clusters, thereby capturing latent heterogeneity associated with interaction dynamics and roadway characteristics.
-Model performance was evaluated using the DIC (Spiegelhalter et al., 2002; El-Basyouny and Sayed, 2012b, a), with lower values indicating superior predictive fit (Zheng et al., 2019).
+The grouped random-parameters formulation in HBSGRP allows selected coefficients to vary across spatial clusters, thereby capturing latent heterogeneity in interaction dynamics and roadway characteristics.
+To compare alternative model specifications, we relied on several Bayesian information criteria commonly used in roadway safety research.
+The Deviance Information Criterion (DIC) (Spiegelhalter et al., 2002; El-Basyouny and Sayed, 2012b, a) is frequently reported as a summary measure of model fit, with lower values indicating superior predictive fit (Zheng et al., 2019).
+However, prior studies have demonstrated that its numerical value can be affected by the model’s parameterization (e.g., Geedipally et al., 2012, 2014).
+Align with earlier safety studies, a difference greater than five DIC units was taken to indicate a meaningfully better fit between competing models (El-Basyouny and Sayed, 2012b, a).
+To strengthen the assessment of predictive performance and reduce sensitivity to parameterization, additionally report the Widely Applicable Information Criterion (WAIC) and the Leave-One-Out Cross-validation Information Criterion (LOOIC), which have been recommended as more robust alternatives to DIC for Bayesian model comparison (Vehtari et al., 2017) and have seen growing use in recent roadway safety applications (Dzinyela et al., 2024; Islam et al., 2023; Khodadadi et al., 2023).
 Estimation was conducted using MultiBUGS, a parallelized extension of OpenBUGS tailored for high-dimensional Bayesian inference.
-Its support for block updating and parallel execution makes it well-suited to modeling spatial dependence, random effects, and extreme value behavior within complex data structures.
+Its support for block updates and parallel execution makes it well-suited for fitting hierarchical models with grouped random effects and extreme-value behavior within complex data structures.
 Parameters were estimated by Markov Chain Monte Carlo (MCMC) simulation; two parallel MCMC chains of 50000 iterations were run, with the first 20000 discarded as burn-in and the remaining 30000 retained for posterior inference.
@@
 For both V–V and V–I interactions, dynamic covariates included relative (speed, distance, acceleration, deceleration, headway, steering angle), jerk, and traffic volume, thereby capturing multiple dimensions of vehicle dynamics.
-Geometric and infrastructure related covariates were also included: lane number, lane width, driveway intensity, median type, and lane position of vehicle.
+Geometric and infrastructure-related covariates were also included: lane number, lane width, driveway intensity, median type, and vehicle lane position.
 All covariates were specified in both the $\mu$ and $\sigma$ components, enabling the models to capture variation in both the central tendency and dispersion of extreme 2D-TTC.
@@
 5.1 Posterior results and model performance
-Tables 5 through 8 report the posterior summaries for both HBSFP and HBSGRP models, estimated separately for V–V and V–I near-misses at intersections and segments.
-Across all cases, the HBSGRP structure consistently outperformed the HBSFP baseline model, confirming its capacity to capture unobserved heterogeneity and location-level variability.
-The DIC values for the HBSGRP models were 508.9 (V–V) and 3179.6 (V–I) at intersections, and 4189.0 (V–V) and 6550.0 (V–I) at segments.
+Tables 58 report the posterior summaries for both HBSFP and HBSGRP models, estimated separately for V–V and V–I near-miss events at intersections and segments.
+Across all cases, the HBSGRP structure consistently outperforms the HBSFP baseline, indicating a better ability to capture unobserved heterogeneity and location-level variability.
+For the HBSGRP models, the DIC values are 508.9 (V–V) and 3179.6 (V–I) at intersections and 4189.0 (V–V) and 6550.0 (V–I) at segments.
 These correspond to relative reductions of 3.41% and 0.54% at intersections, and 7.49% and 3.11% at segments, compared with their HBSFP counterparts.
 Such improvements exceed the widely accepted threshold of five DIC units, indicating substantively better model fit (El-Basyouny and Sayed, 2012b, a).
-To further evaluate robustness, WAIC and LOOIC were also computed, both of which produced rankings consistent with DIC, thereby reinforcing the superior performance of the HBSGRP.
+To further evaluate robustness, WAIC and LOOIC were also computed, both of which produced rankings consistent with DIC, thereby reinforcing the HBSGRP’s superior performance.
 By allowing selected coefficients to vary across spatial groups, the HBSGRP models capture latent heterogeneity in both the central tendency and dispersion, a capability absent in fixed-effect structures.
-Collectively, these results demonstrate that grouped random parameter models provide a more flexible and accurate framework for modeling COR dynamics in complex urban traffic environments.
+Collectively, these results demonstrate that grouped random parameters models provide a more flexible and accurate framework for modeling COR dynamics in complex urban traffic environments.
 5.1.1 Intersection-level
-The intersection-level HBSGRP–UGEV model was designed to capture the heterogeneity of extreme 2D-TTC distributions from V-V interactions at signalized intersections.
-In the modeling stage, all covariates were considered as random effects; however, roadway characteristics and few vehicle dynamics were exhibited negligible variation across sites and did not improve posterior fit.
-To maintain parsimony and interpretability, random effects were restricted to the intercept and two dynamic covariates (relative speed and distance), while other covariates were retained as fixed effects, following recommendations from (Anis et al., 2025a).
-This choice reflects theoretical expectations: vehicle dynamics are highly variable across intersections depending on signal phasing, approach alignment, and driver behavior, whereas roadway structure tends to exert a more stable influence.
-The resulting specification captures both site-specific dynamics and broader geometric constraints, group-specific random effects illustrated in Fig.
-5.1.1 with posterior summaries reported in Table 5.
+The intersection-level HBSGRP–UGEV model was designed to capture heterogeneity in extreme 2D-TTC from V–V interactions at signalized intersections.
+In an initial specification, all covariates were allowed to vary across site groups; however, the final random-parameter set was determined using a theory-first rationale and then empirically confirmed and parsimonious.
+Specifically, random effects were retained for the intercept and two varying-dynamics covariates, relative (speed and distance).
+These quantities directly govern the short-horizon evolution of 2D-TTC and are expected to vary across intersections with signal phasing, approach geometry, queue discharge conditions, and driver compliance.
+In contrast, other vehicle dynamics (e.g., relative acceleration/deceleration, jerk, heading/steering differences) are typically more episodic and noise-sensitive at high frequency; once closing dynamics are accounted for, their between-site variation was not consistently identifiable and contributed limited additional explanatory power.
+To avoid overfitting and preserve interpretability, restricted random effects to this minimal, theoretically motivated subset and relied on hierarchical partial pooling to shrink site-specific coefficients toward corridor-level means when information was limited.
+Candidate random-effects structures were compared using posterior stability (non-degenerate random-effect variances), predictive information criteria, and posterior predictive checks; covariates with random-effect variances concentrated near zero or unstable posteriors were retained as fixed effects (Anis et al., 2025a).
+The resulting specification captures intersection-to-intersection variability in closing behavior while maintaining a parsimonious, generalizable structure; group-specific random effects are illustrated in Fig.
+5.1.1, with posterior summaries reported in Table 5.
 (a) $\beta_{\mu,0[1,10]}$
@@
   • HBSFP
-    Hierarchical Bayesian Spatial Fixed Parameter
+    Hierarchical Bayesian Structure Fixed Parameter
   • HBSGRP
-    Hierarchical Bayesian Spatial Grouped Random Parameter
+    Hierarchical Bayesian Structure Grouped Random Parameters
   • a
@@
 The direction and magnitude of these coefficients provide insights into the mechanisms underlying COR.
-Relative speed and distance emerge as the dominant dynamics, influencing both the location ($\mu$) and scale ($\sigma$) parameters.
-For $\mu$, relative speed has a positive effect, indicating that larger speed differentials between vehicles sharply reduce 2D-TTC and intensify near-miss severity.
-In contrast, relative distance shows a negative effect, meaning shorter headways systematically elevate severity by limiting available buffer space for evasive action.
-For $\sigma$, both covariates carry negative signs, implying that high speeds and short distances compress variability, yielding consistently critical outcomes with little room for driver recovery through braking or swerving.
-Taken together, these results confirm that speed amplifies severity while short headways exacerbate risk and reduce variability, producing deterministic near-miss outcomes under constrained conditions.
+Relative speed and distance emerge as the dominant dynamics, influencing both $\mu$ and $\sigma$ parameters.
+Relative speed exhibits a positive association with the location component and a negative association with the scale component (Fig.
+8(b), Fig.
+8(e)), whereas relative distance exhibits a negative association with the location component and a negative association with the scale component (Fig.
+8(c), Fig.
+8(f)).
+Taken together, these changes imply that higher closing-rate conditions and tighter spacing shift the fitted extreme distribution toward more critical tail behavior while simultaneously concentrating the distribution.
 Fig.
-5.1.1 illustrates the spatial heterogeneity across intersections.
-While most locations show consistent trends, the I2 site exhibits unusually wide credible intervals for the $\mu$ effects and intervals exceeding zero for the $\sigma$ across the intercept, relative speed, and relative distance.
-This instability is attributable to I2’s small BM sample size, which weakens posterior precision and increases uncertainty in group-level effects.
-Despite localized uncertainty, the dominant intersection-level trend persists, with relative speed and distance exerting stable and significant effects on near-miss outcomes.
-Beyond random effects, several covariates were modeled as fixed parameters due to their stable contributions across intersections.
+5.1.1 illustrates the spatial heterogeneity across intersections, and the corresponding group-level posterior intervals are provided.
+While most locations show consistent trends, site I2 exhibits wider credible intervals for the location effects and less stable scale-related posteriors, attributable to its smaller BM sample size and the resulting weaker posterior precision.
+Beyond random effects, several covariates were modeled as fixed effects because of their stable contributions across intersections.
 These effects can be interpreted in terms of driver behavior.
 These factors highlight the roles of longitudinal dynamics, lateral control, maneuver type, and exposure in shaping near-miss outcomes.
-Table 5 shows that relative acceleration has a significant negative effect on both the location and scale parameters, suggesting smoother and less variable near misses when vehicles accelerate through green phases or clear intersections without abrupt interaction.
-By contrast, relative deceleration increases severity ($\mu$) but reduces variability ($\sigma$), capturing emergency braking events.
-These situations sharply intensify near misses but also produce uniform outcomes, since drivers in crisis have limited response options.
-Jerk, representing the rate of change in acceleration, shows the opposite pattern: a negative effect on $\mu$ but a positive effect on $\sigma$.
-This suggests that rapid throttle–brake oscillations, often signs of hesitation or overcorrection, slightly lower average severity but introduce greater unpredictability, underscoring the instability of abrupt corrective maneuvers.
-Lateral vehicle dynamics also play an important role in near-miss risk estimation.
-Steering difference is a significant factor in the HBSGRP model, showing a positive association with $\mu$ and a negative association with $\sigma$.
-This indicates that abrupt steering maneuvers intensify near-miss severity but produce more consistent outcomes within constrained intersection geometries.
-Heading difference shows a similar pattern, with a positive effect on $\mu$ and a negative effect on $\sigma$.
-This suggests that even small misalignments in vehicle trajectory elevate severity while reducing variability, reflecting the rigid geometric constraints of signalized approaches.
-The categorical indicator for left-turn movements is significant, with positive effects on both $\mu$ and $\sigma$.
-This finding is consistent with the well-documented complexity of left-turn phases, which involve interactions with opposing through traffic and diverse driver decision strategies.
-Left turns therefore not only elevate near-miss severity but also increase variability, reflecting differences in intersection design and driver behavior under turning conditions.
-By contrast, traffic volume is statistically insignificant for both parameters but was retained in the model to represent exposure.
-Its lack of effect suggests that short-duration extreme 2D-TTC events are driven more by instantaneous vehicle interactions than by aggregate demand.
+Table 5 shows that relative acceleration is negatively associated with both the location and scale components, implying lower predicted exceedance probability and a more concentrated extreme distribution under higher-acceleration conditions.
+By contrast, relative deceleration is positively associated with the location component and negatively associated with the scale component, implying higher predicted exceedance probability with a more concentrated tail under braking-dominant interactions.
+Jerk is negatively associated with the location component but positively associated with the scale component, implying lower central extreme levels but greater dispersion in tail behavior, i.e., more variability in predicted exceedance probability across comparable contexts.
+Lateral control variables also contribute meaningfully at intersections.
+Steering difference shows a positive association with the location component and a negative association with the scale component, while heading difference exhibits the same sign pattern.
+Taken together, these coupled changes imply that larger lateral corrections and trajectory misalignment shift the fitted extreme distribution toward more crash-prone tail behavior while concentrating it, resulting in higher exceedance probability under unstable lateral maneuvers in constrained intersection geometry.
+The left-turn indicator is significant, with positive associations in both the location and scale components, indicating that left-turn maneuvers shift the extreme 2D-TTC distribution toward more crash-prone behavior while increasing dispersion, consistent with the added conflict complexity of opposing flows and heterogeneous gap-acceptance.
+Traffic volume is not statistically significant in either component; it is retained as an exposure control, and its weak effect suggests that block-level extremes are dominated by instantaneous interaction states rather than by aggregate demand within the short scenario windows.
 Finally, the estimated shape parameter ($\xi$) is negative and significant across all intersections (Fig.
@@
 Importantly, group-level variability in $\xi$ suggests that intersection design and control strategies influence not only the frequency but also the extremal nature of near-miss outcomes.
-The HBSGRP model for V–I near misses specified the intercept as a random parameter in both the location and scale components, capturing baseline heterogeneity in severity and variability across intersections.
-Relative speed (in $\mu$) and relative distance (in $\sigma$) were also modeled as random effects (Fig.
-5.1.1), although their 95% credible intervals overlapped zero.
-They were nevertheless retained on theoretical grounds, as their inclusion improved overall model fit.
-This specification reflects behavioral intuition: heterogeneity in speed effects is more pronounced in severity (location), whereas heterogeneity in clearance effects is more evident in variability (scale).
-All other covariates were treated as fixed effects, reflecting their comparatively stable behavioral and geometric influences.
-Fit statistics (Table 6) show modest but consistent improvements in DIC, WAIC, and LOOIC relative to the HBSFP specification, suggesting that the grouped random parameter structure adds explanatory nuance but produces smaller performance gains than in the V–V case.
+The intersection-level HBSGRP model for V–I near-misses specified the intercept as a grouped random parameters in both the location and scale components (Fig.
+5.1.1) to capture baseline heterogeneity across intersections.
+Relative speed was modeled as a grouped random effect in the location component, and relative distance was modeled as a grouped random effect in the scale component.
+Although these effects are weakly identified at some sites (credible intervals include zero), they were retained because they are mechanistically relevant to boundary-encroachment dynamics and because they improve predictive fit relative to the fixed-parameter specification.
+All remaining covariates were treated as fixed effects.
+Fit statistics (Table 6) show modest but consistent improvements in DIC, WAIC, and LOOIC versus HBSFP, indicating smaller gains from grouped random parameters for V–I than for V–V at intersections.
 Table 6: Posterior estimates for V-I near misses at Intersections
@@
 $\beta_{\sigma,2}$ Fixed Relative acceleration 0.012 0.015 [-0.015, 0.042] 0.046 0.011 $0.009,0.084]^{\dagger}$
-$\beta_{\sigma,3}$ Fixed Relative deceleration -0.147 0.056 $0.285,-0.027]^{\dagger}$g -0.177 0.056 $0.289,-0.047]^{\dagger}$
+$\beta_{\sigma,3}$ Fixed Relative deceleration -0.147 0.056 $0.285,-0.027]^{\dagger} -$0.177 0.056 $0.289,-0.047]^{\dagger}$
 $\beta_{\sigma,4}$ Fixed Relative distance 0.209 0.038 $0.140,0.282]^{\dagger} - - -$
@@
   • HBSFP
-    Hierarchical Bayesian Spatial Fixed Parameter
+    Hierarchical Bayesian Structure Fixed Parameter
   • HBSGRP
-    Hierarchical Bayesian Spatial Grouped Random Parameter
+    Hierarchical Bayesian Structure Grouped Random Parameters
   • a
@@
 Fig. 10.: Posterior estimates for V-I near-misses at Intersections
-In the location parameter, several interpretable effects emerge.
-Relative distance is strongly negative and significant, confirming that reduced clearance to the roadway edge sharply lowers 2D-TTC and elevates severity.
-Relative deceleration also enters with a significant negative sign, contrasting with its positive effect in V–V near misses.
-This suggests that braking near infrastructure often reflects proactive yielding or evasive maneuvers, which mitigate severity rather than intensify it.
-Heading and steering differences are positive and significant, indicating that lateral deviations (e.g., turning toward or away from curbs) increase near-miss severity.
-Left-turn movement, by contrast, shows a significant negative association, suggesting that left-turn trajectories provide greater clearance to boundaries and reduce severity compared with right turns.
-This pattern may reflect geometric characteristics such as wider turning radii or reduced overlap with curbs and crosswalks during left-turn phasing.
-Finally, relative speed—although modeled as a random effect—does not reach significance in the pooled analysis, reinforcing that V–I severity is less sensitive to speed differentials than V–V encounters.
-The scale parameter results provide additional insight into variability.
-The negative random intercept indicates baseline differences across intersections, with some sites producing more predictable outcomes than others.
-Relative speed shows a positive and significant association with $\sigma$, implying that faster approaches to infrastructure increase variability in severity—in contrast to the V–V case, where speed compressed dispersion.
-This divergence reflects the diverse ways drivers respond to boundary proximity: some brake aggressively, while others maintain speed or adjust trajectory.
-Relative acceleration is also positive and significant, reinforcing that heterogeneous acceleration behavior near boundaries broadens outcome variability.
-By contrast, relative deceleration has a significant negative effect, indicating that decisive braking compresses the distribution into consistently critical outcomes.
-Jerk is similarly positive, confirming that abrupt throttle–brake oscillations amplify variability.
-Lateral dynamics play the opposite role: heading and steering differences both reduce variability, suggesting that once a lateral maneuver is committed, outcomes become more predictable within intersection geometry.
-Finally, left-turn movement increases variability despite lowering mean severity, consistent with heterogeneous layouts and driver strategies in left-turn contexts.
-The shape parameter remains negative and significant across intersections, supporting the Weibull domain for V–I near misses and confirming that these events have a bounded upper tail.
+Relative deceleration shows a negative association with the location parameter and a negative association with the scale parameter, indicating that stronger braking near roadway boundaries is associated with a systematic shift in extreme behavior and a concentration of the distribution.
+Relative distance shows a negative association with the location parameter, whereas its contribution to the scale component is not statistically significant at the pooled level, suggesting that reduced clearance robustly shifts the distribution toward more critical extremes across intersections, whereas its influence on dispersion is more site-dependent and uncertain.
+Lateral control indicators show a consistent coupled structure: both heading and steering differences exhibit positive associations with the location parameter and negative associations with the scale parameter, indicating that greater lateral misalignment is associated with shifted extreme behavior accompanied by reduced dispersion.
+The left-turn indicator shows a negative association with the location parameter and a positive association with the scale parameter, suggesting that left-turn contexts shift the location component while broadening dispersion, which supports threshold-dependent risk responses when exceedance probabilities are evaluated across severity levels.
+In contrast, speed and throttle-related dynamics act primarily through dispersion: relative speed and relative acceleration show positive associations with the scale parameter, while their pooled associations with the location component are weak, indicating that approach-speed and acceleration behavior near infrastructure is more strongly linked to variability in extremes than to a uniform location shift.
+The shape parameter remains negative and significant across intersections (9(c)), supporting the Weibull domain for V–I near misses and confirming that these events have a bounded upper tail.
 This indicates a physical and behavioral ceiling on severity, constrained by roadway geometry and driver response capacity.
 Spatial variation in $\xi$ further underscores the influence of intersection design and infrastructure placement on the extremal properties of near-miss distributions.
-Taken together, the intersection-level results suggest that V–I near misses are shaped more by clearance, braking, and lateral maneuvers than by speed differentials.
-In contrast, V–V near misses are dominated by high speeds, which produce deterministic severity, whereas V–I outcomes exhibit greater variability driven by the interplay of infrastructure geometry, driver avoidance strategies, and intersection control.
+Taken together, the intersection-level results suggest that V–I near misses are more strongly shaped by clearance, braking, and lateral maneuvers than by speed differentials.
+In contrast, V–V near misses are dominated by high speeds, which yield deterministic severity, whereas V–I outcomes exhibit greater variability driven by the interplay among infrastructure geometry, driver avoidance strategies, and intersection control.
 5.1.2 Segment-level
-The HBSGRP model for segment-level V–V near misses introduces random parameters within a hierarchical structure (Table 7).
-Specifically, the intercept and three time-varying vehicle dynamics—relative speed, relative deceleration, and relative distance—were allowed to vary across segments (Fig.
-5.1.2).
-This specification recognizes that longitudinal interactions in uninterrupted midblock contexts are shaped by both driver behavior and roadway characteristics (lane configuration, median type, driveway density).
-All remaining covariates—relative acceleration, jerk, heading and steering differences, traffic volume, lane attributes, and access density—were treated as fixed effects given their comparatively stable influence across segments.
-The HBSGRP model yields substantially better fit than the fixed-effect alternative, confirming the value of partial pooling for capturing heterogeneity across the nine directional segments.
-Refer to caption Refer to caption
-(a) $\beta_{\mu,0[1,18]}$ (b) $\gamma_{\mu,1[1,18]}$
-Refer to caption Refer to caption
-(c) $\gamma_{\mu,2[1,18]}$ (d) $\gamma_{\mu,3[1,18]}$
-Refer to caption Refer to caption
-(e) $\beta_{\sigma,0[1,18]}$ (f) $\gamma_{\sigma,1[1,18]}$
-Refer to caption Refer to caption
-(g) $\gamma_{\sigma,2[1,18]}$ (h) $\gamma_{\sigma,3[1,18]}$
-(i) $\beta_{\xi,0[1,18]}$
-Fig. 11.: Posterior estimates for V-V near-misses at segments
+The segment-level HBSGRP model for V–V near-misses specifies grouped random parameters to capture heterogeneity across directional segments (Table 7).
+Random effects were assigned to the location and scale intercepts and to three time variant covariates: relative speed, relative deceleration, and relative distance, allowing their contributions to vary by segment .
+All remaining covariates (relative acceleration, jerk, heading and steering differences, traffic volume, lane attributes, and driveway/access density) were treated as fixed effects to maintain a parsimonious specification when additional between-segment variation was not reliably identifiable.
+The HBSGRP model provides a better fit than the HBSFP alternative, supporting partial pooling for modeling segment-level variation in extreme near-miss risk.
 Table 7: Posterior estimates V-V near misses at Segments
@@
   • HBSGRP
-    Hierarchical Bayesian Spatial Grouped Random Parameter
+    Hierarchical Bayesian Spatial Grouped Random Parameters
   • a
@@
     Indicates statistical significance at the 95% level (interval excludes 0)
-For the location parameter ($\mu$), the random effects reveal meaningful differences across segments that align with expected driving dynamics.
-Relative speed shows a consistently positive effect, confirming that faster closing speeds sharply reduce 2D-TTC and intensify near-miss severity.
-The magnitude of this effect, however, varies across roadway contexts Fig.
-5.1.2.
-Segments with uninterrupted high-speed flows (e.g., S11) show the strongest impacts, while access-rich segments (e.g., S6, S7, and S13) display wider credible intervals, reflecting greater variability in interactions.
-Relative deceleration carries a negative sign overall, but its effect is less stable, suggesting that braking sometimes helps dissipate conflicts in smoother-flow environments, whereas in higher-speed segments the same braking often signals critical near-miss events.
-Relative distance exhibits a consistently negative association, indicating that shorter headways reliably elevate severity by restricting the available buffer for evasive maneuvers.
-The posterior variation across segments suggests that spacing behavior differs between divided and undivided facilities.
-Finally, in some segments with sparse block maxima samples, the intercept’s credible intervals approach zero, highlighting weaker estimation stability under limited data conditions.
-Among the fixed effects, relative acceleration shows a significant negative association with $\mu$, indicating that acceleration phases are systematically linked to more severe near-misses on segments.
-This pattern contrasts with intersections, where acceleration often reflects flow resumption under green signals.
-Along segments, however, aggressive acceleration during overtaking or merging reduces spacing and elevates risk.
-Jerk also exhibits a negative effect, underscoring how abrupt longitudinal control contributes to critical outcomes.
-Lateral control indicators—heading and steering differences—display positive coefficients, suggesting that lateral adjustments provide modest increases in 2D-TTC where lane-changing flexibility is available.
-Driveway density is positively associated with $\mu$, a result that may seem counterintuitive but likely reflects anticipatory caution by drivers in access-rich environments.
-In contrast, undivided medians show a strong negative effect, confirming that exposure to opposing traffic flows produces consistently more severe interactions.
-For the scale parameter ($\sigma$), the random effects for relative speed, deceleration, and distance are all negative, indicating that extreme longitudinal dynamics compress outcome variability and yield uniformly critical near-miss conditions.
-This pattern is evident in Fig.
-5.1.2, where high-speed and short-spacing conditions cluster tightly around lower 2D-TTC values, producing narrow posteriors in segments such as S12 and S15.
-By contrast, segments like S10 display wider credible intervals, reflecting weaker estimation stability under sparse BM samples.
-Among the fixed effects, jerk and heading difference increase $\sigma$, suggesting that volatile acceleration or lateral corrections introduce greater heterogeneity in near-miss outcomes.
-In contrast, lane number and lane width reduce $\sigma$, implying that additional maneuvering space fosters more predictable interactions.
-Driveway density increases dispersion, consistent with the variability created by frequent merging and crossing maneuvers.
-Median type again shows a strong negative effect, reinforcing that undivided facilities compress severity outcomes into consistently critical ranges.
-Finally, the shape parameter ($\xi$) remains negative and statistically significant across segment groups, confirming a bounded upper tail consistent with the Weibull domain.
-This finding indicates physical and behavioral constraints on extreme near-miss severity, governed by speed environment, roadway geometry, and driver reaction capacity.
-Overall, the segment-level results emphasize the central role of longitudinal dynamics—acceleration, braking, and headway—in determining near-miss severity, while roadway features such as medians, lane width, and access density primarily shape variability.
-The strong performance of the HBSGRP model underscores its value in capturing these segment-specific dynamics, which fixed-parameter structures fail to represent, particularly in accounting for the spatially contingent effects of speed and spacing in midblock contexts.
+For the segment-level V–V model, the grouped random effects indicate that closing dynamics shift the fitted extreme distribution through paired changes in the location and scale components, and that these joint shifts vary across segment contexts.
+Relative speed exhibits a positive association with the location component and a negative association with the scale component, implying that as closing speed increases, the fitted distribution moves toward more critical extremes while simultaneously becoming more concentrated around those extremes.
+The posterior variation across segments indicates that this coupled shift is strongest on segments characterized by more uninterrupted progression, whereas access-rich segments show wider posterior intervals, consistent with more heterogeneous interaction patterns.
+Relative distance shows a negative association with the location component and a negative association with the scale component, indicating that tighter spacing is linked to more critical tail behavior and less dispersion in extreme outcomes.
+Relative deceleration contributes through a coupled location/scale pattern whose magnitude is segment-dependent, consistent with the fact that braking can represent either conflict dissipation or crisis response depending on prevailing speeds and spacing; this heterogeneity is reflected in the segment-specific posterior intervals.
+In segments with fewer block-maxima observations, the intercept posteriors are less precise, but the corridor-level pattern remains: closing-rate and spacing jointly govern how the extreme distribution shifts and concentrates, which is the mechanism through which COR changes.
+Among the fixed effects, relative acceleration and jerk have negative associations with the location component, whereas jerk has a positive association with the scale component, suggesting that aggressive longitudinal control is associated with a shift in the external distribution and increased dispersion in outcomes.
+Heading difference has positive associations with both the location and scale components, whereas steering difference has a positive location association but a scale association whose interval overlaps zero, indicating that lateral misalignment tends to shift the external distribution, with less consistent evidence that it systematically changes dispersion.
+Exposure and geometry covariates also show coupled behavior: volume and lane width have positive location associations, while lane number and lane width have negative scale associations, suggesting that cross-sectional capacity primarily manifests through changes in dispersion rather than a consistent shift in the center of extremes.
+Driveway density shows positive associations at both the location and scale levels, indicating that access activity is linked to both a shift and a widening of the extreme distribution.
+The median type (undivided) shows negative associations in both location and scale, and left-lane position shows a positive location association and a negative scale association, indicating that cross-section and lane placement shape COR through simultaneous changes in the fitted center and the dispersion of extreme near-miss outcomes.
+Overall, segment-level COR is best interpreted as the result of these coupled location–scale shifts, with grouped vehicle dynamics providing the primary source of spatial heterogeneity (Table 7).
+Finally, the shape parameter ($\xi$) is negative and statistically significant across segment groups, indicating a bounded upper tail for extreme near-miss behavior.
+Overall, longitudinal dynamics (speed, braking, and spacing) drive segment-level extremes, while roadway features (median type, lane width, and driveway density) primarily modulate variability.
+The improved fit of the HBSGRP model shows that allowing key dynamics to vary by segment is important for capturing midblock heterogeneity that fixed-parameter models miss.
 The HBSGRP model for segment-level V–I near-misses offers insights into how longitudinal dynamics and roadway design jointly shape the severity and variability of extreme 2D-TTC outcomes (Table 8, Fig.
 5.1.2).
-In this model, the intercept and relative distance are specified as group-level random effects in both the location and scale components, capturing spatial heterogeneity in how spacing influences near-miss severity across segments.
-All other covariates, including relative speed, acceleration, deceleration, jerk, heading difference, steering angle difference, traffic volume, lane attributes, driveway density, median type, and left-lane presence—are modeled as fixed effects to represent their more consistent contributions across segment environments.
+In particular, the location intercept is random and significant, and the relative distance is random in both components, with a negative association in the location component and a positive association in the scale component.
+Implying that reduced clearance shifts extremes toward more critical outcomes while increasing dispersion, and that this sensitivity varies by segment.
+All other covariates, including relative speed, acceleration, deceleration, jerk, heading difference, steering angle difference, traffic volume, lane attributes, driveway density, median type, and left-lane presence, are modeled as fixed effects to represent their more consistent contributions across segment environments.
 (a) $\beta_{\mu,0[1,18]}$
@@
   • HBSGRP
-    Hierarchical Bayesian Spatial Grouped Random Parameter
+    Hierarchical Bayesian Spatial Grouped Random Parameters
   • a
@@
     Indicates statistical significance at the 95% level (interval excludes 0)
-For the location parameter, reduced clearance to roadway boundaries exerts a strong negative effect on 2D-TTC, confirming that closer spacing to infrastructure consistently produces more critical near-miss outcomes.
-The random effect for distance further indicates that this influence varies across segment groups, reflecting sensitivity to local geometric and operational conditions such as median presence or driveway density.
-Relative speed shows a notable divergence: while the fixed-parameter model suggests higher speeds intensify severity, the grouped specification produces a small positive estimate, implying that in some contexts faster travel may coincide with greater clearance from boundaries (e.g., drivers maintaining lane center on free-flow arterials).
-This reversal highlights how heterogeneous speed environments fundamentally alter the role of speed in shaping V–I severity.
-Relative acceleration is positive and significant, suggesting that acceleration phases often represent clearance from roadside constraints (e.g., moving away from curbs or barriers) rather than heightened risk.
-By contrast, relative deceleration retains a negative effect across both models, indicating that braking near boundaries is typically associated with abrupt evasive actions, sharply reducing 2D-TTC and producing more critical near-miss conditions.
-Several geometric and operational features also shape V–I near-miss severity at the segment level.
-Heading and steering differences exhibit significant positive coefficients, indicating that lateral adjustments toward or away from boundaries provide modest clearance gains and mitigate severity.
-Traffic volume, lane number, and lane width also emerge as positive contributors in the HBSGRP specification, suggesting that wider, higher-capacity segments allow greater separation from roadside features, producing slightly larger 2D-TTC values.
-The effect of driveway density, however, is particularly notable: while insignificant in the fixed-parameter model, the grouped specification yields a large negative estimate with a wide CI.
-This suggests that dense roadside access points intensify boundary conflicts by forcing frequent merging and turning maneuvers, though the data remain sensitive to localized characteristics.
-Median type is also significant, with undivided facilities consistently producing lower 2D-TTC values, reflecting heightened exposure to opposing flows and limited recovery space near boundaries.
-The left-lane variable shows a sign reversal, with the fixed model suggesting higher severity in interior lanes and the grouped model indicating greater severity in outer lanes.
-This pattern highlights heterogeneity across segment groups, where lane position relative to curbs, medians, or roadside access alters the likelihood of boundary-related near misses.
-The scale parameter provides complementary insights into the variability of boundary-related near misses.
-The random effect for relative distance is positive, indicating that reduced clearance to roadside features not only increases severity but also broadens variability, as drivers adopt divergent strategies when negotiating minimal spacing with curbs, medians, or barriers.
-Relative speed also shows a positive association with scale, suggesting that higher speeds amplify heterogeneity in boundary interactions—contrasting with V–V near misses, where speed reduces dispersion—because drivers respond differently to infrastructure proximity under fast-moving conditions.
-Acceleration, by contrast, has a negative coefficient in the grouped model, implying that acceleration phases near boundaries often reflect smoother clearance maneuvers, yielding more predictable outcomes.
-Relative deceleration also reduces variability, indicating that braking close to infrastructure compresses outcomes into uniformly critical events.
-Geometric attributes further shape variability.
-Lane width exerts a positive effect, showing that while wider lanes provide greater average clearance, they also increase variability by enabling a broader range of lateral positioning relative to boundaries.
-The undivided median type strongly increases scale, underscoring that the absence of a raised median not only heightens risk but also expands outcome variability due to heterogeneous driver avoidance maneuvers.
-The left-lane variable is also positive, indicating that boundary proximity in outer lanes produces more heterogeneous near-miss outcomes compared to interior lanes.
-Finally, driveway density, though statistically unstable, suggests that variability is strongly site-dependent, as frequent access points create unpredictable driver–infrastructure interactions.
-The shape parameter remains negative and significant across both specifications, confirming that the upper tail of the 2D-TTC distribution is bounded.
-This supports the Weibull domain for extremes and indicates that even in V–I segment interactions, physical and behavioral limits constrain how severe the most critical boundary-related near-misses can become.
-Variation in the shape parameter across segment groups underscores the role of local geometry, access density, and operational conditions in shaping the extremal behavior of these events.
-Segment-level analyses show that roadway geometry and vehicle dynamics jointly shape near-miss severity and variability.
-For V–V events, relative speed and distance dominate—speed amplifies severity while distance compresses variability, producing more deterministic outcomes under close-following conditions.
-For V–I events, the effect of speed is heterogeneous, sometimes reducing severity when clearance to boundaries is maintained, but elevating it when high speeds bring vehicles closer to curbs or medians.
-Access-related factors, particularly driveway density and median type, are critical for V–I outcomes, underscoring the role of access management and median design in roadside safety.
-Across both interaction types, the negative shape parameter indicates bounded extremes, reflecting physical and behavioral limits on crash risk escalation.
+Regarding the fixed parameters, for instance Relative speed is positive in both components, indicating that higher-speed boundary interactions are associated with a systematic shift and greater dispersion.
+Relative acceleration is positive in the location component and negative in the scale component, whereas relative deceleration is negative in both components, indicating that braking-dominated boundary encounters are associated with a shift toward more critical extremes and a more concentrated distribution.
+Heading and steering differences are positive in the location component and negative in the scale component, suggesting that lateral deviations coincide with shifted extremes but with reduced dispersion.
+Roadway and exposure variables show similar coupled behavior.
+Lane number is positive in the location component and negative in the scale component, while lane width is positive in both components, indicating that additional lanes align with reduced dispersion, whereas wider lanes align with greater dispersion.
+Driveway density is negative in the location component and does not show a stable scale association, indicating access-rich segments align with more critical extremes, with dispersion effects less consistent.
+Undivided medians are negative in the location component and positive in the scale component, linking undivided sections to more critical and more variable extremes.
+Vehicle position (left lane) is negative in the location component and positive in the scale component, indicating lane context shifts extremes while increasing dispersion.
+The shape parameter is negative and significant across sites (Fig.
+10(e)), indicating a bounded upper tail and supporting the Weibull domain for extreme 2D-TTC.
+This implies that, even at the segment level, physical and behavioral constraints limit how severe boundary-related extremes can become.
+Differences in this parameter across segments suggest that local geometry and operating conditions influence tail behavior.
+Overall, segment results show that vehicle dynamics and roadway features jointly govern both the shift and spread of extremes.
+For V–V interactions, relative speed, relative deceleration, and relative distance are the primary drivers, jointly shifting the distribution towards more critical outcomes and concentrating it under high closing rate, hard braking, and tight spacing conditions.
+For V–I interactions, relative distance is the dominant dynamic driver, with its segment-specific variability reflecting how clearance to curbs/medians governs boundary-related extremes.
 6 Model validation
-The study corridor dataset provided precise geolocation for each extreme near-miss event (Eqn.
-32) along with annual corridor-wide crash totals.
-Although annual crash data offer a useful aggregate benchmark, they lack precise geolocation of individual crashes and therefore cannot be directly matched to near-miss predictions at the event level.
-To address this limitation, validation was instead designed as an event-level case–control exercise, assessing whether the model assigns higher exceedance probabilities to more severe near-miss events.
-Posterior exceedance probabilities were computed (Eqn.
-30) across severity thresholds from –0.1 to –0.9 seconds, and classification performance was evaluated using ROC–AUC, a threshold-independent metric well suited to rare-event prediction.
+The trajectory-based near-miss dataset provides precise geolocation for each extracted extreme interaction (Eq.
+32) because all vehicle states are map-referenced and time-stamped at high frequency.
+These geolocated extremes can be aggregated to corridor elements (intersections and directional segments) and compared spatially.
+In contrast, the available historical crash records for the study corridor are reported with imprecise geolocation; therefore, they cannot be matched to specific near-miss trajectories.
+Consequently, direct external validation that links predicted near-miss risk to individual crash records is not feasible with the available crash data.
+Validation is therefore designed to assess: (i) out-of-sample discrimination on unseen extreme blocks and (ii) spatial generalizability through site-by-site comparisons of expected versus observed extremes with uncertainty.
+Both HBSFP–UGEV and HBSGRP–UGEV were evaluated using these validation procedures; only the superior-performing HBSGRP–UGEV results are reported.
+Model performance was evaluated using an out-of-sample event case–control design.
+The dataset was split at the block level into 70% training and 30% testing.
+For each severity threshold $\omega\in{-0.1,-0.2,\ldots,-0.9}$ sec, a test-set block was labeled as severe if its observed block maximum exceeded $\omega$.
+For each held-out block, the posterior exceedance probability was computed from the fitted HBSGRP–UGEV and HBSFP-UGEV model using Eq.
+(30).
+Because the HBSGRP–UGEV specification consistently outperformed the HBSFP–UGEV baseline in predictive criteria and posterior stability, only HBSGRP–UGEV validations are reported here.
+Receiver operating characteristic (ROC) curves were constructed by varying the classification probability cutoff, and predictive performance was summarized using the area under the ROC curve (ROC-AUC), which is threshold-independent and suitable for rare-event classification.
 (a) V–V Segment
@@
 Fig. 12.: ROC curves for four HBSGRP–UGEV models across severity thresholds from $\omega=-0.1$s to $\omega=-0.9$s
-ROC curves were generated for the four HBSGRP–UGEV variants, covering both V-V and V-I interactions at the segment and intersection levels.
-The results (Figure 6) reveal clear distinctions across near-miss types and roadway contexts.
-The V–V segment model achieved the highest predictive accuracy (AUC $\approx$ 0.89), reflecting the influence of dynamic variability in driver behavior.
-Relative speed, deceleration, and distance emerged as significant random parameters, capturing heterogeneity in conflict severity across segments.
-At the same time, fixed effects such as heading difference, steering difference, jerk, lane width, and driveway density contributed additional explanatory power, indicating that both dynamic interactions and roadway design shape V–V near-miss risks.
-The V–I model at the segment level achieved moderately strong accuracy (AUC $\approx$ 0.79) but displayed greater sensitivity to severity thresholds.
-Only relative distance emerged as a significant random parameter, while other vehicle dynamics (e.g., speed, deceleration, etc.) and roadway features (e.g., lane width, driveway density, median type, and etc.) contributed as fixed effects.
-This suggests that V–I risks are shaped primarily by localized conditions, with dynamic variability playing a comparatively smaller role than in V–V interactions.
-At intersections, the V–V model achieved moderately strong accuracy (AUC $\approx$ 0.82) with low threshold sensitivity.
-Random parameters for relative speed and relative distance captured cross-location heterogeneity in approach and gap-acceptance conditions, while fixed effects— including deceleration/acceleration, steering and heading differences, jerk, and a left-turn indicator—explained lateral maneuvering and control effects.
-Together, these terms allow the model to represent both longitudinal and lateral conflict dynamics under signalized conditions, yielding stable performance.
-In contrast, the intersection-level V-I model performed weakest (AUC $\approx$ 0.75) and showing large fluctuations across thresholds, which reflects the complexity of boundary-related near-misses in intersections where turning trajectories, curb proximity, traffic signals, and pedestrian activity all interact.
-Taken together, the results demonstrate that the HBSGRP–UGEV framework yields robust and consistent predictive accuracy for V–V interactions, particularly along corridor segments, where conflict dynamics are more regular and less influenced by intersectional variability.
-By contrast, V–I models—especially at intersections—showed lower stability and greater sensitivity to threshold specification, reflecting the complexity of boundary-related near-miss shaped by turning flows, curb geometry, and multimodal activity.
-These findings suggest that segment-level V–V estimates are the most operationally reliable for proactive monitoring, while intersection-level V–I models may require further refinement.
-Incorporating richer representations of curb design, pedestrian exposure, and high-resolution V-I proximity measures could strengthen stability and predictive performance in future applications.
+ROC curves for the four HBSGRP–UGEV variants (V–V and V–I; segments and intersections) are shown in Fig.
+6.
+The V–V segment model achieved the strongest identification (AUC $\approx 0.89$), followed by the V–V intersection model (AUC $\approx 0.82$).
+The V–I models exhibited lower identification, with AUC $\approx 0.79$ for segments and AUC $\approx 0.75$ for intersections.
+These results indicate that the framework more reliably distinguishes severe V–V near-misses, particularly along midblock segments, than infrastructure-related V–I near-misses in intersection environments, where turning trajectories, curb/median proximity, and signalized operations introduce greater variability.
+While ROC–AUC assesses overall discrimination, it does not indicate whether the model reproduces spatial variation across corridor elements.
+To evaluate spatial generalizability, expected versus observed extremes were compared site by site, accounting for uncertainty.
+For each site $k$ and threshold $\omega$, the observed number of severe extremes in the test set was computed as the count of held-out blocks at site $k$ whose observed block maximum exceeded $\omega$.
+The corresponding expected number of severe extremes was computed by summing the model-based exceedance probabilities across the same held-out blocks.
+Uncertainty in expected counts was quantified using posterior predictive simulation, propagating posterior uncertainty in exceedance probabilities to replicated exceedance counts and summarizing the resulting predictive intervals.
+The expected-observed results indicate that the fitted models generally reproduce the spatial ordering of extremes across corridor elements, supporting the ability of the HBSGRP–UGEV framework to capture site-to-site variation rather than only aggregate frequency.
+The comparisons also reveal a systematic calibration pattern across thresholds: the model tends to overestimate exceedances at more severe thresholds (i.e., deeper tail levels) and underestimates them at less severe thresholds.
+This behavior suggests that the fitted tail response may be overly sensitive for the most extreme conditions and comparatively conservative for milder near-miss conditions.
+Together, the hold-out ROC–AUC results (Fig.
+6) and the site-by-site expected–observed comparisons with uncertainty provide complementary evidence of predictive performance: the former evaluates out-of-sample discrimination on unseen blocks, while the latter tests whether predicted extremes generalize spatially across corridor elements.
+The primary limitation remains the absence of event-level crash geolocation and timestamps, which preclude direct external validation of crash matches.
+Future work can strengthen external validation by leveraging crash datasets with event-level spatial–temporal precision and by implementing explicit spatial cross-validation designs (e.g., leave-one-site-out) when site-level sample sizes become sufficiently large.
 7 Summary and conclusions
-This study addresses the persistent challenge of estimating short-term crash occurrence risk (COR) in urban corridor-wide traffic environments, where conventional surrogate safety models often overlook detailed vehicle geometry, roadway context, and site-specific behavioral variability.
-Existing approaches have typically focused on vehicle–vehicle (V–V) interactions and neglected vehicle–infrastructure (V–I) conflicts, limiting their applicability at the corridor scale.
-To overcome these gaps, we introduced a geometry-aware two-dimensional time-to-collision (2D-TTC) indicator that incorporates vehicle dimensions, dynamic states, and roadway boundaries, enabling the detection of extreme near-miss events from both V–V and V–I interactions.
-Extreme events were modeled through a nonstationary Univariate Generalized Extreme Value (UGEV) distribution embedded within a Hierarchical Bayesian Spatial Grouped Random Parameter (HBSGRP) framework.
-This structure allowed partial pooling across intersections and directional segments while preserving local heterogeneity, addressing data sparsity and contextual variation.
-High-resolution vehicle trajectories from the Argoverse-2 dataset were integrated with HD maps to extract near-miss events.
-Vehicle dynamics (relative speed, acceleration, deceleration, jerk, heading and steering differences, and traffic volume) and roadway features (lane width, driveway density, median type, turning movements, and lane position) were incorporated as covariates, with grouped random parameters specified for key dynamic variables.
-Grouped random parameters were specified for temporally varying dynamics, relative speed, relative deceleration, and relative distance, to capture their context-dependent effects.
-For benchmarking, a baseline hierarchical Bayesian fixed-parameter (HBSFP) model was also estimated.
-Results showed that the HBSGRP model consistently outperformed its fixed-parameter counterpart, yielding lower DIC, WAIC, and LOOIC values and demonstrating greater stability.
-Random-parameter vehicle dynamics revealed that relative speed, deceleration, and distance dominated V–V near misses at the segment level, whereas only relative distance was significant for V–I near misses.
+This study addresses the challenge of estimating short-term crash occurrence risk (COR) in urban corridors, where conventional surrogate safety approaches often rely on simplified interaction geometry and provide limited support for corridor-scale inference under heterogeneous operations.
+Existing trajectory-based EVT studies have primarily emphasized vehicle–vehicle (V–V) interactions and rarely incorporate vehicle–infrastructure (V–I) conflicts within a unified corridor-scale framework, thereby limiting their applicability for proactive safety monitoring and infrastructure-relevant risk assessment.
+To bridge these gaps, we proposed a geometry-aware two-dimensional time-to-collision (2D-TTC) indicator that propagates oriented vehicle footprints and integrates HD-map roadway boundaries, enabling consistent extraction of extreme near-miss events from both V–V and V–I interactions.
+Extreme near-misses were modeled using a nonstationary univariate generalized extreme value (UGEV) distribution embedded in a Hierarchical Bayesian Structure Grouped Random Parameters (HBSGRP) framework.
+The grouped structure supports partial pooling across intersections and directional segments while retaining location-specific heterogeneity, improving stability under sparse extreme samples.
+High-resolution trajectories from the Argoverse-2 dataset were combined with HD maps to construct interaction-window block maxima of 2D-TTC and to derive covariates representing (i) interaction dynamics (e.g., relative speed, acceleration/deceleration, jerk, and heading/steering differences, with exposure represented by traffic volume) and (ii) roadway context (e.g., lane width, driveway density, median type, turning movements, and lane position).
+Grouped random parameters were assigned to a small, theoretically motivated subset of high-impact dynamics to capture context-dependent effects without overfitting, and a baseline hierarchical Bayesian fixed-parameter model (HBSFP) was estimated for benchmarking.
+Results indicate that the HBSGRP specification consistently improves predictive performance and posterior stability relative to HBSFP, as reflected by lower DIC, WAIC, and LOOIC values.
+Importantly, covariate effects were interpreted in terms of their implied impact on COR (i.e., on the fitted UGEV distribution and crash-level exceedance probability), rather than through individual GEV parameters in isolation.
+Under this distribution-level interpretation, closing-dynamics variables (relative speed, deceleration, and distance) emerged as the most influential drivers of crash-prone extreme risk for V–V interactions, particularly on segments, while V–I risk was more strongly tied to spacing/proximity conditions and boundary encroachment.
 At intersections, V–V near misses were dominated by relative speed and distance, while no dynamic variables were significant for V–I near misses.
-Fixed covariates such as heading difference, steering difference, lane width, driveway density, and median type also showed significant effects, demonstrating the ability of the model to capture both dynamic behaviors and roadway design influences.
-Validation results further confirmed model credibility.
-ROC–AUC validation confirmed robust predictive accuracy, with V–V segment models performing best (AUC $\approx$ 0.89), followed by intersection V–V ($\approx$ 0.82)), while V–I models showed moderate performance ($\approx$ 0.79 for segments, $\approx$ 0.75 for intersections).
-These findings indicate that while V–V risks can be reliably modeled, V–I interactions, particularly at intersections, remain more variable and complex.
-While the proposed framework advances corridor-wide COR modeling, several aspects warrant further development.
-First, reliance on AV onboard sensors may introduce measurement bias in traffic state estimation and near-miss detection.
-Second, the 2D-TTC formulation simplifies nonlinear vehicle dynamics, which could understate risks during extreme evasive maneuvers.
-Third, the framework does not yet capture near-miss severity stratification.
-Future research offers opportunities to enrich this framework.
-Extending it to multivariate HBSGRP-UGEV models with complementary surrogate measures such as post-collision velocity change ($\Delta V$) would provide deeper insights into crash severity dynamics.
-Incorporating vulnerable road user (VRU) near-misses, validating across diverse urban networks, calibrating for transferability, and benchmarking against GPD methods will further enhance robustness and generalizability.
-As AV deployment expands, increasingly rich trajectory datasets will strengthen the statistical power of HBSGRP–UGEV and open new opportunities for nuanced, multimodal safety assessments.
-Within this evolving data landscape, the proposed framework provides a scalable and adaptable foundation for proactive corridor-wide safety monitoring.
-By enabling early identification of high-risk conditions and supporting real-time traffic management, the framework aligns with the broader Vision Zero agenda, advancing the long-term goal of safer, more efficient, and equitable urban mobility.
+Roadway and maneuver context (e.g., steering/heading deviations, lane width, driveway density, median type, and turning movements) also contributed meaningfully to COR variation, reinforcing that corridor risk is jointly shaped by short-horizon behavior and geometric/infrastructure constraints.
+Model credibility was supported through validation against held-out outcomes using ROC–AUC.
+Predictive discrimination was strongest for V–V segment models (AUC $\approx$ 0.89), followed by V–V intersection models (AUC $\approx$ 0.82), while V–I models achieved moderate performance (segments: $\approx$ 0.79; intersections: $\approx$ 0.75).
+These results suggest that V–V crash-prone extremes are more consistently explainable from interaction dynamics at the corridor scale, whereas V–I interactions—particularly at intersections—exhibit greater variability due to complex boundary approach behavior, turning paths, and map-defined infrastructure constraints.
+While the proposed framework advances corridor-wide COR estimation, several limitations motivate future research.
+First, Argoverse-2 motion-forecasting scenarios are intentionally curated toward safety-critical, high-interaction episodes; therefore, the estimated COR should be interpreted as conditional on these traffic states.
+Second, the 2D-TTC projection assumes simplified short-horizon dynamics, which may underrepresent risk under extreme evasive maneuvers or highly nonlinear control responses.
+Third, V–V and V–I extremes are modeled as separate univariate processes under the interaction-window block definition; extending to a bivariate or multivariate hierarchical EVT formulation would require a common space–time block that yields paired extremes and sufficient information to identify joint tail dependence.
+Finally, the current study focuses on crash occurrence risk from extreme near-misses, without explicit severity stratification within the EVT layer.
+Several extensions can further strengthen the proposed framework.
+First, incorporating longer-duration and more representative trajectory streams from continuously operating AV fleets would allow benchmarking the estimated COR against network-wide baseline risk and assessing temporal stability beyond short interaction windows.
+Second, the current univariate EVT formulation could be extended to multivariate hierarchical EVT models that jointly represent V–V and V–I extremes, provided that a common space–time blocking strategy yields paired extremes with sufficient information to identify tail dependence.
+Third, integrating complementary physics-informed severity surrogates, such as post-impact delta-V or energy-based indicators, would enable joint modeling of crash occurrence and severity rather than occurrence alone.
+Fourth, extending the framework to include vulnerable road user (VRU) near-misses would improve relevance for urban safety applications.
+Finally, systematic comparison with POT–GPD-based formulations under comparable blocking and dependence controls would help clarify trade-offs between BM and POT approaches for short-horizon, high-frequency trajectory data.
+As trajectory data scale in coverage and duration, hierarchical EVT models such as HBSGRP–UGEV will gain additional statistical power, enabling more nuanced, multimodal safety assessments.
+Within this evolving data landscape, the proposed framework provides a scalable foundation for proactive corridor-wide monitoring, supporting earlier identification of crash-prone conditions and enabling data-driven traffic management consistent with Vision Zero goals.
 Acknowledgement: